### v1

# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

Abstract
The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and the underutilization of spectral complementarity. Existing methods often fail to decouple shared structural features from modality-specific radiometric attributes, leading to feature conflicts and information loss. To address this issue, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-specific) components in the Fourier domain. Specifically, PAD consists of two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features through convolution-guided scaling to enhance geometric consistency, and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency details and low-frequency structures using frequency-adaptive multilayer perceptrons. This approach leverages SAR’s sensitivity to morphological features and RGB’s spectral richness. Extensive experiments on WHU-OPT-SAR and DDHR-SK datasets demonstrate state-of-the-art performance. Our work establishes a new paradigm for physics-aware multi-modal fusion in remote sensing. The code will be available at https://github.com/RanFeng2/PAD.
Index Terms:
Land Cover Classification, RGB-SAR Multi-modality, Multi-modal Segmentation, Synthetic Aperture Radar, Remote Sensing
I Introduction
Land cover classification (LCC) is a fundamental task in Earth observation that involves categorizing various surface materials, such as farmlands, forests, water bodies, and urban areas, using remote sensing imagery. This capability is essential for environmental monitoring, resource management, and urban development [1]. In recent years, fusing Synthetic Aperture Radar (SAR) and RGB imagery has emerged as a promising approach to improve LCC performance. SAR provides robust structural information that is less susceptible to atmospheric and illumination variations, while RGB imagery offers rich spectral details, compensating for SAR-specific artifacts such as speckle noise [2, 3].
Despite their complementary relation, a key challenge remains: the inherent imaging heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB), which leads to significant spectral misalignment and hinders the effective fusion of cross-modal features. Compared to conventional multi-modal fusion in natural images, integrating SAR and RGB data presents more pronounced radiometric and geometric disparities, limiting the performance of existing frameworks [4, 5, 6].
Figure 1: SAR-RGB Spatial-Frequency Complementarity Analysis. Frequency-domain representations capture modality complementarity more effectively than spatial features. (a) Spatial discrepancies exhibit anisotropic distributions; (b) relative amplitude differences are pronounced in high-frequency regions; (c) phase discrepancies are minimal and clustered in low-frequency bands; (e)–(f) variations in sampling rates significantly affect the spatial discrepancy distribution. Analysis conducted on WHU-OPT-SAR (see Section III-A).
Deep learning methods, particularly convolutional neural networks (CNNs), have achieved notable success in LCC tasks due to their powerful feature extraction capabilities [7]. To mitigate modality heterogeneity in SAR-RGB fusion, most approaches employ dual-branch architectures to independently extract multi-scale features from each modality [8], followed by fusion via simple operations such as element-wise summation, mean aggregation, or channel-wise concatenation, before processing with a shared decoder or segmentation head. Although this fusion scheme relies on accurately modeling complementary features, the intrinsic heterogeneity between SAR and RGB often impedes the extraction of truly complementary representations. Recent studies, inspired by contrastive learning [9, 10, 11], have explored aligning highly correlated, modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts. For example, Li et al. [12] considered phase as a modality-shared feature and used it as input in the intermediate network stage. However, they not only disregarded high-frequency amplitude patterns, which are essential for terrain delineation, thus limiting the effectiveness of the fusion, but also failed to systematically prove the complementarity of phase between the two modalities.
Through comprehensive spectral analysis (Fig. 1), we identify three critical frequency-domain priors that have been previously neglected: 1) Spatial-domain alignment (Fig. 1(a)) is significantly hampered by anisotropic discrepancies between modalities, rendering direct pixel- or feature-level matching unreliable; 2) Existing phase-based strategies overlook the inevitable slight differences between different frequency bands (Fig. 1(c)), leading to information loss; and 3) Amplitude mismatches between SAR and RGB modalities present different patterns in the high/low-frequency regions, most pronounced in the high-frequency band (Fig. 1(b)), highlighting the need for frequency-aware difference modeling.
Motivated by these insights, we propose a novel fusion method, Phase-Amplitude Decoupling (PAD). The central idea behind PAD is to disentangle SAR and RGB modalities into phase components, which capture modality-invariant structural information, and amplitude components, which encode modality-specific details. By explicitly aligning the shared phase representations, PAD enhances the extraction of complementary amplitude features, thereby reducing feature conflicts caused by cross-modal heterogeneity. Specifically, PAD comprises two key modules: 1) a Phase Spectrum Correction (PSC) module that ensures robust phase alignment across modalities, and 2) an Amplitude Spectrum Fusion (ASF) module that selectively integrates complementary amplitude cues for improved fusion performance.
Our contributions are summarized fourfold:
  • Through the systematic analysis of conventional registration datasets, we establish the frequency distribution law of SAR-RGB common patterns (phase) and unique patterns (amplitude) for the first time, providing an interpretable physical prior for multi-source remote sensing fusion.
  • We propose a novel spectral decomposition method that orthogonally separates phase (modality-shared) and amplitude (modality-specific) features, effectively resolving feature conflicts in SAR-RGB fusion while maintaining geometric integrity.
  • To address phase inconsistencies, the PSC module employs convolution-guided multiplicative scaling to align and enhance critical structural details, ensuring robust fusion.
  • The ASF module leverages frequency-domain multilayer perceptrons to dynamically integrate high-frequency (local details) and low-frequency (global structure) components, capitalizing on the complementary strengths of SAR and RGB data for improved segmentation.
Figure 2: Flowchart of the PAD Framework. The framework comprises the (a) SCF module, (c) ASF module, and (d) PSC module. Registered RGB and SAR images, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are first input, and their features, $x_{RGB}$ and $x_{SAR}$, are extracted at each backbone stage before being fed into $n$ PAD fusion modules. Meanwhile, (b) FD and (e) FR denote the transformations from the spatial to the frequency domain and vice versa, respectively. Each stage’s final output is fed into a shared decoder. Here, $\mathcal{A}$ and $\mathcal{P}$ represent amplitude and phase, and $C$ (outer circle) indicates channel-wise concatenation.
II Related Work
II-A Multi-modal Fusion for Land Cover Classification
While single-modality approaches [13, 14, 15, 16, 17, 18, 19, 20] have achieved reasonable performance in homogeneous data scenarios, their inability to leverage cross-modal synergies limits their discriminative power in complex environments. SAR-RGB fusion methods generally follow two paradigms: knowledge transfer via domain adaptation and hierarchical feature integration. In this section, we analyze their technical evolution and inherent limitations.
II-A1 Cross-Modal Transfer Learning
Transfer learning [21] seeks to adapt knowledge from one modality to another. However, the significant differences between SAR and RGB data often degrade learned representations, limiting their applicability in multi-modal tasks.
II-A2 Hierarchical Feature Fusion
Feature fusion strategies have progressed through three developmental phases: a) Basic Fusion Operators: Early methods utilized basic operations such as concatenation or weighted summation [22]. However, these approaches are prone to modality dominance and information redundancy in complex scenes. b) Attention-Guided Fusion: Spatial-channel attention mechanisms were introduced to address these issues [23, 24]. For instance, MCANet introduces cross-modal attention (MCAM) to model inter-modality dependencies, MRFS [25] develops progressive cycle attention (PC-Att) for infrared-visible fusion, and CMX [26] introduces bidirectional channel-spatial attention to refine cross-modal features. However, most existing attention paradigms have high computational costs. c) Complementarity-Aware Fusion: Recent state-of-the-art approaches explicitly model modality-specific and modality-shared components. ASANet [27] uses semantic focusing modules (SFM) to emphasize complementary features, CDDFuse [10] applies contrastive learning to disentangle private and shared representations, and CEN [28] proposes a method that achieves complementary perceptual fusion through dynamic channel exchange and parameter-sharing-independent batch normalization. Furthermore, recent work [12] pioneers phase-structure alignment by leveraging phase information as a cross-modal invariant feature. Despite these advancements, spatial-domain methods still struggle with geometric misalignment and computational inefficiency.
II-B Frequency Representation Learning in Vision Tasks
Frequency analysis offers powerful tools to address challenges in spatial fusion [29, 30, 31]. Three primary application directions are discussed below:
II-B1 Phase as a Modality-Invariant Representation
Decoupling amplitude and phase spectra has proven effective in domain generalization (DG), where models trained on multiple source domains must generalize to unseen target domains. In DG, phase spectra encode structural semantics (e.g., object shapes and edges), while amplitude spectra capture modality-specific styles (e.g., illumination and texture patterns) [32, 33]. Early DG methods [34, 35] attempted to disentangle style and content by normalizing feature statistics (mean/variance), assuming that style corresponds to statistical variations. However, such heuristic separations can distort structural content. Recent approaches employ Fourier transforms as explicit style-content separators. For example, amplitude-phase recombination [36] swaps the amplitude spectrum of a source image with that of a target domain while preserving the source phase, thereby synthesizing cross-domain training samples. This strategy compels networks to focus on phase-centric structural features, improving robustness to style shifts. Building on this, decompose-adjust-compose (DAC) [37] integrates Phase-Amplitude Decoupling with feature normalization to dynamically balance style and content in the frequency domain. These methods highlight a critical advantage of spectral decomposition: phase alignment preserves structural consistency across modalities, while amplitude manipulation enables adaptive style fusion.
II-B2 Frequency-Driven Global Context Modeling
Beyond style-content separation, the Fourier transform’s global receptive field has been exploited to overcome two key limitations of spatial-domain networks: restricted long-range dependency modeling and the quadratic complexity of self-attention. For instance, fast Fourier convolution (FFC) [38] introduced hybrid spatial-frequency processing with parallel branches, capturing global patterns via spectral operations while preserving local details through convolution. These principles have also been applied to optimize attention mechanisms. FFTformer [39] replaces dot-product attention with spectral element-wise multiplication, reducing complexity from $O(N^{2})$ to $O(N\log N)$ while maintaining global interactions. Similarly, GFNet [40] and AFNO [41] employ learnable frequency filters as token mixers, achieving significant speedups compared to efficient self-attention variants. These methods demonstrate that global context aggregation in the frequency domain is not only computationally efficient but also capable of fusing multi-scale information.
II-B3 Frequency Priors as Inductive Biases
Recent studies have shown that frequency-domain representations can be explicitly designed as task-specific inductive biases [42, 43, 44]. For DG, DFF [45] exploits the assumption that low frequencies control cross-domain structural consistency, introducing a frequency-domain dynamic filtering mechanism to amplify domain-invariant components while suppressing high-frequency variations. In dense prediction tasks, FreqFusion [46] applies adaptive frequency filtering: an Adaptive Low-Pass Filter (ALPF) generator smooths high-frequency perturbations to homogenize intra-class features, while an Adaptive High-Pass Filter (AHPF) generator restores critical boundary information lost during downsampling. This strategy encodes spectral priors that low frequencies ensure semantic consistency, while high frequencies enhance boundary discriminability. In the context of RGB-NIR denoising, FCENet [47] utilizes a Frequency Dynamic Selection Module (FDSM) to select useful frequency components across modalities and a Frequency Exhaustive Fusion Module (FEFM) to refine shared and modality-specific spectral features. These approaches demonstrate that decoupling and resetting frequency components effectively encode task-related physical principles, such as structural invariance and boundary sensitivity, injecting strong inductive biases into deep networks.
III Proposed Method
This paper introduces a multi-modal fusion framework based on Phase-Amplitude Decoupling (PAD) for land cover classification using SAR and RGB images. PAD effectively integrates complementary information from both modalities while mitigating heterogeneous feature conflicts that often arise with traditional spatial fusion techniques. The overall architecture, depicted in Fig. 2, consists of three main modules: the Amplitude Spectrum Fusion (ASF) module, the Phase Spectrum Correction (PSC) module, and the Spatial Complementary Fusion (SCF) module.
Input : SAR/RGB image pairs $\{I_{\mathrm{SAR}}^{(i)},I_{\mathrm{RGB}}^{(i)}\}_{i=1}^{N}$
Output : RSD, RAD, APPD metrics for each image pair
1
1: Preprocessing
//Convert RGB image to grayscale
2 $I_{\mathrm{RGB}}\leftarrow 0.299\,R+0.587\,G+0.114\,B$
//Retain SAR image as is
3 $I_{\mathrm{SAR}}\leftarrow I_{\mathrm{SAR}}$
4
2: Spectral Decomposition
5 for each pair $(i)$ do
       //Frequency domain conversion
6       Compute $\mathcal{A}_{\mathrm{SAR}}^{(i)}$ and $\mathcal{P}_{\mathrm{SAR}}^{(i)}$ via Eqs. 11, 12 and 13
7       Compute $\mathcal{A}_{\mathrm{RGB}}^{(i)}$ and $\mathcal{P}_{\mathrm{RGB}}^{(i)}$ similarly
8
      //Spatial difference computation
9       Compute $\Delta I^{(i)}\leftarrow\mathrm{RSD}^{(i)}$ via Eq. 1
       //Amplitude difference computation
10       Compute $\Delta\mathcal{A}^{(i)}\leftarrow\mathrm{RAD}^{(i)}$ via Eq. 2
       //Phase difference computation
11       Compute $\Delta\mathcal{P}^{(i)}\leftarrow\mathrm{APPD}^{(i)}$ via Eq. 3
12
13 end for
14
3: Dataset-level Aggregation
//Mean spatial, amplitude, phase difference computation
15 $\mathrm{RSD}\leftarrow\mathtt{Mean}(\{\Delta I^{(i)}\}_{i=1}^{N})$
16 $\mathrm{RAD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{A}^{(i)}\}_{i=1}^{N})$
17 $\mathrm{APPD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{P}^{(i)}\}_{i=1}^{N})$
Algorithm 1 Spectral Difference Analysis.
III-A Frequency-domain Analysis for Modality Decoupling
To validate the physical significance of frequency-domain decoupling in multi-modal data, we conducted a systematic spectral analysis on WHU-OPT-SAR. This analysis includes visualizations of the averaged difference spectra (Fig. 1(a)–(c)) computed using Algorithm 1 to compare the differences between two modalities in different domains. The difference spectra are defined as:
$\mathrm{RSD}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|I_{\mathrm{SAR}}^{(i)}-I_{\mathrm{RGB}}^{(i)}\right|}{I_{\mathrm{RGB}}^{(i)}+\epsilon},$ (1)
$\mathrm{RAD}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|\mathcal{A}_{\mathrm{SAR}}^{(i)}-\mathcal{A}_{\mathrm{RGB}}^{(i)}\right|}{\mathcal{A}_{\mathrm{RGB}}^{(i)}+\epsilon},$ (2)
$\mathrm{APPD}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\arg\exp j(\mathcal{P}_{\mathrm{SAR}}^{(i)}-\mathcal{P}_{\mathrm{RGB}}^{(i)})\right|,$ (3)
where $I_{\mathrm{SAR}}$, $\mathcal{A}_{\mathrm{SAR}}$, and $\mathcal{P}_{\mathrm{SAR}}$ denote the pixel values, amplitude, and phase spectra of SAR images, respectively; $I_{\mathrm{RGB}}$, $\mathcal{A}_{\mathrm{RGB}}$, and $\mathcal{P}_{\mathrm{RGB}}$ correspond to those of the RGB images; $\arg(\cdot)$ computes the angle of a complex number; ||$\cdot$|| ⋅ | denotes the absolute value; and $\epsilon$ is a small constant to prevent division by zero. By leveraging the periodicity of complex exponentials, APPD maps phase differences into the principal value range $[-\pi,\pi)$, eliminating abnormally large differences caused by phase wrapping (e.g., a true phase difference of $2\pi$ is corrected to 00).
Figure 1 reveals three key observations: (1) Spatial differences (RSD) exhibit anisotropic distributions; (2) Amplitude differences (RAD) predominantly emerge in high-frequency regions; (3) Phase differences (APPD) remain minimal with slight low-frequency clustering. The frequency-domain metrics (RAD, APPD) demonstrate superior specificity compared to spatial-domain RSD, confirming that spectral decomposition better captures cross-modal complementarity.
These analyses lead to two fundamental insights that drive our architecture design: a) Phase Sharing: The consistently low APPD values indicate strong phase consistency between modalities. This suggests that phase spectra contain shared information that can guide cross-modal learning. Our PSC Module thus performs pixel-level phase alignment while preserving this inherent consistency. b) Amplitude Complementarity: The RAD distribution reveals frequency-dependent amplitude divergence, with high-frequency components showing greater discrepancy than low-frequency ones. This localized spectral pattern motivates our ASF Module, which adjusts the frequency radius dynamically to handle frequency-specific amplitude variations.
These phenomena can be physically interpreted based on the inherent characteristics of each imaging modality. Specifically, phase discrepancies manifest differently across frequency bands. Low-frequency disparities primarily stem from geometric distortions in imaging, leading to overall deviation between the two modalities that cannot be registered, whereas high-frequency variations arise from SAR speckle noise affecting fine details. On the other hand, amplitude disparities emerge from varying sensitivities to different ground objects, with low/high frequency contrasts reflecting modality-specific responses to large/small scale structures (e.g., farmlands and forests vs. roads).
We further validate our spectral analysis through sensitivity experiments on sampling rates (Fig. 1(e)–(f)). By analyzing the mean and variance of the difference spectra, we found that both RAD and APPD metrics are largely invariant to sampling rate, confirming that the observed phase-amplitude complementarity originates from intrinsic modality properties, rather than from sampling artifacts.
TABLE I: Statistical Analysis Results for the Average Periodic Phase Difference Spectrum.
Metric Low Freq High Freq Whole Image
Shapiro-Wilk (Stat) 1.0044 1.0050 1.0131
Shapiro-Wilk (p) 1.0 1.0 1.0
Skewness -0.0521 0.0020 -0.0265
Kurtosis 0.0750 -0.0117 0.0322
Figure 3: Statistical Analysis of APPD. (a) APPD spectrum; (b)–(d) Frequency histograms with Gaussian fits for low-frequency (LF, radius = 0.5 $\times$ half-diagonal), high-frequency (HF), and full spectra (ALL); (e)–(f) Mean-variance trajectories across sampling rates. The HF, LF, and ALL spectra show near-Gaussian invariance, while LF/ALL deviates slightly due to the dominant structural noise in the LF range. These results suggest that phase, especially HF, can serve as a robust shared feature.
To assess the statistical consistency of cross-modal phase spectra, we conducted normality tests (Shapiro-Wilk), along with skewness and kurtosis analyses (Table I). Frequency histograms and mean-variance trajectories (Fig. 3) indicate that the frequency difference spectra follow a near-Gaussian distribution, particularly for high-frequency components, whereas low-frequency components exhibit minor systematic deviations, likely due to structural noise.
Figure 4: Visualization of Spatial and Frequency Spectrum Differences in a Scenario with Missing RGB Regions. (a)–(b) show RGB images with missing regions; (d)–(f) present the difference spectrum when the missing regions are included; and (g)–(i) display the difference spectrum after excluding the missing regions. (c) indicates that variation in RSD is significantly higher than in RAD and APPD, demonstrating that frequency-domain representations are more robust to missing data.
Figure 5: Visualization of Spatial and Frequency Spectrum Differences in a Cloud-Obscured Observations Scenario. (a)–(c) present an example of cloud-obscured observations, with cloud masks generated using RGB, SAR, and NIR images from WHU-OPT-SAR [48]. (d) compares the stability of RSD, RAD, and APPD between cloud-covered and cloud-free regions. Numbers next to the arrow indicates the value of decline.
We also evaluate phase spectrum invariance under challenging conditions through two controlled experiments: (1) partial RGB data missing (Fig. 4) and (2) cloud-obscured observations (Fig. 5). In the missing RGB scenario, RSD is highly sensitive to data incompleteness (Fig. 4(d) vs. (g)), while phase differences remain consistent (Fig. 4(f) vs. (i)). Similarly, in cloud-affected cases, the APPD variance ($\Delta\approx 10^{-0.037}\%$) is negligible compared to that of RSD ($\Delta\approx 10^{5.81}\%$) and RAD ($\Delta\approx 10^{0.58}\%$) as shown in Fig. 5(d). Collectively, these results demonstrate that phase information (particularly in the high-frequency region) exhibits robust consistency between SAR and RGB modalities, supporting its role as a shared feature to mitigate inter-modal conflicts and leverage complementary information in multi-modal frameworks.
III-B Overall Framework
As illustrated in Fig. 2, the proposed PAD framework integrates SAR and RGB modalities via a dual-branch encoder and a shared decoder, with its primary innovation in cross-modal feature fusion through phase-amplitude decoupling. The encoder utilizes asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$ with independent parameters to preserve modality-specific attributes. Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave scattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images. Both backbones produce hierarchical feature maps at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$, such that $i\in\{0,1,2,3\}$:
$x_{RGB}^{i}=\mathcal{B}_{\mathrm{RGB}}^{(i)}\left(I_{\mathrm{RGB}}^{i}\right),\quad x_{SAR}^{i}=\mathcal{B}_{\mathrm{SAR}}^{(i)}\left(I_{\mathrm{SAR}}^{i}\right).$ (4)
Initially, the SCF module (Eq. 5) processes the bimodal features by computing pixel-wise fusion weights through spatial attention gating. This operation preliminarily integrates complementary features in the spatial domain and reduces feature dimensionality, thereby constraining the solution space for subsequent frequency-domain operations.
$X^{i}=\mathrm{SCF}\left(x_{RGB}^{i},x_{SAR}^{i}\right),$ (5a)
$\mathcal{S}=\mathrm{Sigmoid}\left(\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1},x_{2}\right)\right)\right),$ (5b)
$x_{1}^{\prime}=\mathcal{S}\odot x_{1},\quad x_{2}^{\prime}=\left(1-\mathcal{S}\right)\odot x_{2},$ (5c)
$x=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1}^{\prime},x_{2}^{\prime}\right)\right).$ (5d)
The fused features are then transformed via FD into separate amplitude and phase spectra:
$\mathcal{A}^{i},\mathcal{P}^{i}=\mathrm{FD}\left(X^{i}\right).$ (6)
The amplitude spectrum $\mathcal{A}^{i}$ is enhanced by the ASF module, which distinguishes high-frequency details from low-frequency structures using a trainable frequency radius $r$:
$\mathcal{A^{\prime}}^{i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right).$ (7)
Simultaneously, the phase spectrum is corrected via multiplicative scaling in the PSC module:
$\mathcal{P^{\prime}}^{i}=\mathrm{PSC}\left(\mathcal{P}^{i}\right).$ (8)
Finally, FR reconstructs the features, yielding the final fusion output:
$\bm{F^{\prime}}^{i}=\mathrm{FR}\left(\mathcal{A^{\prime}}^{i},\mathcal{P^{\prime}}^{i}\right).$ (9)
The shared decoder, $\mathcal{D}$, aggregates the multi-stage fused features to generate high-precision land-cover segmentation maps:
$\bm{\hat{Y}}=\mathcal{D}\left(\bm{F^{\prime}}^{0},\bm{F^{\prime}}^{1},\bm{F^{\prime}}^{2},\bm{F^{\prime}}^{3}\right).$ (10)
III-C Frequency Decoupling & Frequency Recoupling
To achieve physically disentangled representations in the spectral domain, we propose a bidirectional transformation framework comprising FD and its inverse, FR. First, the input image $X\in\mathbb{R}^{H\times W}$ is transformed from the spatial domain to the frequency domain using a 2D real-valued fast Fourier transform (rFFT) [49], yielding the spectrum $\mathcal{F}\in\mathbb{C}^{H\times W}$:
$\mathcal{F}\left(u,v\right)=\sum_{x=0}^{W-1}\sum_{y=0}^{H-1}X\left(x,y\right)\exp\left(-i2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
where $(u,v)$ represents frequency-domain coordinates. Exploiting the conjugate symmetry of real-valued signals, only the non-redundant region (with $20\leq u\leq W/$ 2 and $0\leq v<H$) is retained. Next, an FFT Shift operation relocates low-frequency components to the center of the spectrum:
$\mathcal{F}_{\mathrm{shift}}\left(u,v\right)=\mathcal{F}\left(u\oplus\frac{W}{2},v\oplus\frac{H}{2}\right),$ (12)
where $\oplus$ denotes modulo addition. Finally, the amplitude and phase spectra are derived as:
$\mathcal{A}=|\mathcal{F}|,\quad\mathcal{P}=\angle\mathcal{F},$ (13)
with $\mathcal{A}\in\mathbb{R}^{H\times W}$ capturing radiometric properties and $\mathcal{P}\in\mathbb{R}^{H\times W}$ preserving structural details. This decoupling enables separate learning of SAR scattering features and RGB structural priors.
III-D Phase Spectrum Correction Module
As illustrated in Fig. 1(c), minor misalignments in the low-frequency phase spectra between SAR and RGB images, attributable to inherent geometric distortions (e.g., side-looking SAR vs. RGB orthographic projection), can compromise spatial consistency and boundary precision. Traditional phase alignment methods relying on rigid transformations are inadequate for non-rigid distortions. Thus, we propose a lightweight PSC module that performs adaptive geometric correction using learnable multiplicative masks. Given an input phase spectrum tensor $P\in\mathbb{R}^{C\times H\times W}$, the module first compresses the channel dimension via a $1\times 1$ convolution (reducing $C$ to $C/r$, where $r\in\{4,8\}$) and applies a Leaky ReLU activation [50]. A subsequent $1\times 1$ convolution produces a single-channel Phase Correction Mask (PCM) constrained to $(0,1)$ by a Sigmoid function. Pixel-wise multiplicative scaling is then applied as follows:
$P^{\prime}=P\odot\left(1+\text{PCM}\right),$ (14)
where $\odot$ denotes the Hadamard product. This design limits the maximum scaling factor to 2, ensuring the preservation of phase periodicity while requiring significantly fewer parameters than conventional $3\times 3$ convolutions.
III-E Amplitude Spectrum Fusion Module
Our analysis in Fig. 1(b) indicates that the spectral discrepancy between SAR and RGB modalities is primarily concentrated in the medium-to-high frequency components of the amplitude spectra. To exploit this prior while maintaining gradient stability, we design an adaptive spectrum fusion module with three sequential operations. First, we compute a normalized frequency distance metric:
$\mathrm{Dist}(i)=\frac{\sqrt{\left(X_{i}-X_{\mathrm{center}}\right)^{2}+\left(Y_{i}-Y_{\mathrm{center}}\right)^{2}}}{\sqrt{X_{\mathrm{max}}^{2}+Y_{\mathrm{max}}^{2}}},$ (15)
which quantifies component positions relative to the spectrum center $(X_{\mathrm{center}},Y_{\mathrm{center}})$. A trainable boundary radius $r$, constrained via a Sigmoid activation to yield $R=\sigma(r)$, dynamically partitions the spectrum into low- and high-frequency regions. To maintain differentiability, a soft partition is employed:
$\mathrm{Mask}_{\mathrm{high-freq}}=\sigma\left(\left(\mathrm{Dist}-R\right)\cdot\tau\right),$ (16)
with temperature parameter $\tau=10$. The high-frequency components are then enhanced using a position-wise multilayer perceptron (MLP):
$A_{\mathrm{high-freq}}=\mathrm{MLP}\left(A\odot\mathrm{Mask}_{\mathrm{high-freq}}\right),$ (17)
and concatenated with the original amplitude spectra. A subsequent $1\times 1$ convolution and residual connection yield the fused amplitude:
$A^{\prime}=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(A,A_{\mathrm{high-freq}}\right)\right),\quad A\leftarrow A+A^{\prime}.$ (18)
This module preserves low-frequency structures while adaptively reinforcing high-frequency details that capture complementary semantic information. The learnable radius $R$ self-adjusts to the input characteristics, eliminating the need for manual threshold tuning.
III-F Loss Function
Our optimization framework integrates three objectives to jointly promote semantic segmentation accuracy, multi-scale feature learning, and amplitude-phase coherence.
III-F1 Primary Semantic Supervision
The primary segmentation loss is defined as the standard cross-entropy loss:
$\mathcal{L}_{\mathrm{seg}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}_{i,c},$ (19)
where $\Omega$ denotes the set of foreground pixels (excluding background label 255), $y_{i,c}\in\{0,1\}$ indicates the ground-truth class, and $\hat{p}_{i,c}$ is the predicted probability.
III-F2 Auxiliary Semantic Supervision
An auxiliary segmentation head is added at the second-to-last stage ($t=2$) using an identical cross-entropy loss:
$\mathcal{L}_{\mathrm{aux}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}^{(a)}_{i,c}.$ (20)
III-F3 Phase-Amplitude Decoupling Constraint
To resolve the coupling between amplitude and phase spectra, we introduce an Amplitude Consistency Loss that regularizes the fused amplitude $\mathcal{A}^{\prime}_{t}$ to remain faithful to its original decoupled state $\mathcal{A}_{t}$ across feature scales:
$\mathcal{L}_{\mathrm{amp}}=\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}\alpha_{t}\left\|\mathcal{A}^{\prime}_{t}-\mathcal{A}_{t}\right\|_{F}^{2},$ (21)
where $\mathcal{T}=\{0,1,2,3\}$ indexes the feature scales, $\alpha_{t}=0.5+(1.0-0.5)\times\frac{t}{3}$ imposes depth-progressive weighting, and $\|\cdot\|_{F}$ is the Frobenius norm.
III-F4 Composite Objective
The overall loss is formulated as:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{seg}}+\lambda_{1}\mathcal{L}_{\mathrm{aux}}+\lambda_{2}\mathcal{L}_{\mathrm{amp}},$ (22)
with $\lambda_{1}=0.4$ balancing auxiliary supervision and $\lambda_{2}=0.1$ controlling the amplitude regularization strength. This composite loss promotes both geometric invariance, by stabilizing phase-sensitive features through $\mathcal{L}_{\mathrm{amp}}$, and semantic invariance via hierarchical amplitude refinement.
IV Experimental Results
IV-A Dataset Description
WHU-OPT-SAR [23] is a multi-modal dataset developed for land use classification in Hubei Province, China. The dataset comprises 100 pairs of RGB, near-infrared (NIR), and SAR images covering identical geographic regions, with corresponding land type labels. The RGB images are acquired from GF1 satellite data, while the SAR images are sourced from GF3 satellite data. Both image types have been resampled to a uniform spatial resolution of 5 meters, with dimensions of 5556 $\times$ 3704 pixels. The dataset spans seven primary land use categories: farmland, urban areas, villages, water bodies, forests, roads, and others. In our study, images are cropped into non-overlapping patches of 512 $\times$ 512 pixels, resulting in 7000 total patches; 80% (5600 patches) are used for training and 20% (1400 patches) for testing.
DDHR-SK is a subset of the DDHR multi-modal dataset developed for land cover classification (LCC) in Pohang City, South Korea [24]. RGB images are collected from GF2 satellite data and SAR images from GF3 satellite data. To simulate adverse conditions such as clouds and fog, the RGB images were processed with a fog filter using the GNU Image Manipulation Program (GIMP). The dataset comprises 6173 paired cloud-affected RGB and SAR images, resampled to a spatial resolution of 1 meter and resized to 256 $\times$ 256 pixels. It covers five land cover types: buildings, farmland, greenery, water, and roads, with a public split of 3087 training images and 3086 validation images.
IV-B Training Setting
All models were trained and evaluated on a single NVIDIA H800 GPU, while the comparative experiments were carried out on a single NVIDIA A100 GPU. Training was conducted using the MMSegmentation framework [51] with a batch size of 4 and 80k iterations, including validation every 8k iterations. We employed the AdamW [52] optimizer with an initial learning rate of $1\times 10^{-4}$ and a Cosine Annealing schedule (with T_max set to the maximum iteration count). A fixed random seed of 42 was used to ensure reproducibility.
Evaluation metrics include mean intersection over union (mIoU), Kappa coefficient, overall accuracy (OA), and F1 score, with performance reported at the iteration yielding the best mIoU. To assess robustness across diverse scenarios and mitigate class imbalance bias, the Kappa coefficient is computed for each sample, and the mean value is reported:
$\kappa_{i}=\frac{P_{o}^{(i)}-P_{e}^{(i)}}{1-P_{e}^{(i)}},\quad\mathrm{mKappa}=\frac{1}{M}\sum_{i=1}^{M}\kappa_{i},$ (23)
where $P_{o}^{(i)}$ and $P_{e}^{(i)}$ denote the observed and expected agreements for the $i$-th sample, respectively.
IV-C Experimental Results on Datasets
To assess the effectiveness of the proposed PAD method, we compared its performance with 17 different models, including 9 state-of-the-art multi-modal segmentation approaches and 8 cutting-edge unimodal segmentation methods. All experimental results were obtained through our replication of the original experimental configurations. Experiments were performed on two datasets: WHU-OPT-SAR, the largest multi-modal LCC dataset, and DDHR-SK, which contains RGB images captured under cloudy conditions.
TABLE II: Experiment Results (%) on WHU-OPT-SAR.
Method Venue IoU per Category OA mKappa mF1 mIoU
Farmland City Village Water Forest Road Others
RGB DANet [13] CVPR’19 67.51 55.22 46.71 60.85 83.13 38.89 20.43 83.19 49.74 67.40 53.25
OCRNet [14] ECCV’20 67.68 56.48 47.47 61.22 82.98 37.20 17.97 83.23 49.11 66.95 53.00
CGNet [15] TIP’20 61.33 45.96 37.98 49.33 80.98 1.92 0.19 79.26 37.42 50.54 39.67
SegFormer [16] NeurIPS’21 68.18 56.91 46.82 60.85 83.19 35.24 18.73 83.37 49.46 66.79 52.84
SETR [17] CVPR’21 65.25 55.07 42.97 57.80 82.44 31.91 18.36 81.94 46.19 64.73 50.54
Mask2former [18] CVPR’22 67.09 56.64 43.97 60.09 81.47 36.24 21.03 82.69 47.69 66.65 52.36
PIDNet [19] CVPR’23 66.20 55.20 44.95 58.90 82.20 36.36 16.40 82.25 47.20 65.53 51.46
TransUNet [20] MIA’24 67.62 56.28 46.08 60.66 83.25 38.76 19.32 83.18 49.89 67.20 53.14
SAR DANet [13] CVPR’19 60.42 54.12 33.05 54.45 80.57 24.89 15.11 79.44 35.28 60.16 46.09
OCRNet [14] ECCV’20 59.81 52.48 31.21 54.03 79.57 22.49 13.92 78.94 32.62 58.74 44.79
CGNet [15] TIP’20 56.72 46.81 26.32 46.30 77.53 0.27 0.00 76.39 24.64 47.00 36.28
SegFormer [16] NeurIPS’21 60.17 54.71 32.30 54.07 79.79 21.02 15.13 79.31 33.60 59.24 45.31
SETR [17] CVPR’21 60.35 52.29 31.41 53.56 80.15 18.66 15.18 79.12 34.25 58.33 44.51
Mask2former [18] CVPR’22 60.39 52.86 35.69 53.85 78.36 29.16 17.20 79.26 32.76 61.35 46.79
PIDNet [19] CVPR’23 58.59 52.35 29.65 53.60 79.35 23.06 13.74 78.42 32.15 58.33 44.34
TransUNet [20] MIA’24 61.93 54.85 33.42 56.26 80.63 27.56 14.96 80.14 37.45 61.13 47.09
SAR & RGB MCANet [23] JAG’21 64.21 45.49 42.93 56.20 81.81 25.93 8.47 80.80 45.47 59.94 46.44
CEN [28] TPAMI’22 66.46 56.28 45.46 58.09 82.34 32.51 17.82 82.32 47.47 65.36 51.28
TokenFusion [53] CVPR’22 67.87 57.26 46.96 61.66 83.31 36.34 18.76 83.38 49.87 67.09 53.16
DDHRNet [24] JAG’22 59.12 45.67 30.77 47.96 79.59 0.47 0.03 78.00 27.18 48.36 37.66
CDDFuse [10] CVPR’23 59.88 35.08 25.22 52.31 78.77 17.64 15.00 77.02 38.69 54.29 40.56
CMX [26] TITS’23 69.24 58.20 48.45 64.06 83.83 39.99 21.92 84.18 52.76 69.01 55.10
MRFS [25] CVPR’24 69.39 58.13 49.43 64.90 83.87 41.26 21.72 84.21 54.33 69.38 55.53
FTransUNet [5] TGRS’24 64.87 53.84 43.71 55.63 81.53 0.00 0.00 81.25 43.53 52.97 42.80
ASANet [27] ISPRS’24 69.18 56.95 48.58 64.06 83.79 41.48 21.74 84.15 52.62 69.05 55.11
PAD (Ours) 69.85 58.10 49.70 65.39 84.05 43.24 23.47 84.56 53.93 70.14 56.26
IV-C1 Evaluations on WHU-OPT-SAR
As summarized in Table II, our PAD framework achieves significant performance improvements on WHU-OPT-SAR. Compared with the best unimodal method (DANet-RGB [13], achieving 53.25% mIoU and 83.19% OA) and the multi-modal baseline MRFS [25] (55.53% mIoU, 84.21% OA), PAD establishes a new state-of-the-art with 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1. In fine-grained categories, PAD attains 43.24% IoU on Road and 49.70% on Village segmentation, outperforming MRFS by 1.98% and 0.27%, respectively, thereby underlining the precision of our cross-modal fusion strategy.
Figure 6: Visualization Results on WHU-OPT-SAR. Four representative cases are shown: (a) building-dense area, (b) specular water bodies, (c) vegetated region, and (d) urban–rural transition zone. For each case, the following images are presented from left to right: RGB image, SAR image, pseudo-color ground truth, and prediction maps from six competing methods. PAD is compared against leading single-modal baselines (DANet-RGB, TransUNet-SAR) as well as multi-modal approaches (CDDFuse, CMX, MRFS). Regions of interest are highlighted with gray boxes (for building/water areas) and circles (for vegetation zones); further details are provided in Section IV-C.
Visual analyses support these quantitative results. For example, in Fig. 6(a), PAD successfully recovers occluded road segments (highlighted in cyan) in building-dense regions through SAR phase analysis, while RGB-only models ignore these areas. In Fig. 6(b), PAD effectively distinguishes specular reflection artifacts from actual water boundaries by leveraging SAR’s dielectric properties, overcoming limitations inherent in optical sensors under cloud interference. Additionally, Fig. 6(c) demonstrates PAD’s ability to resolve spectral ambiguities in vegetation classification, and Fig. 6(d) shows that PAD maintains cross-category consistency in complex urban–rural transitions.
TABLE III: Experiment Results (%) on DDHR-SK.
Method Venue IoU per Category OA mKappa mF1 mIoU
Buildings Roads Greenery Water Farmland
RGB DANet [13] CVPR’19 94.46 75.11 91.10 99.03 90.59 96.23 73.35 94.57 90.06
OCRNet [14] ECCV’20 90.18 60.49 84.42 97.87 82.84 93.17 58.96 90.26 83.16
CGNet [15] TIP’20 80.19 27.57 72.77 90.47 69.97 86.16 33.90 78.76 68.19
SegFormer [16] NeurIPS’21 92.44 67.11 88.76 98.82 88.18 95.00 66.73 92.71 87.06
SETR [17] CVPR’21 94.17 74.10 91.43 99.18 90.88 96.22 73.79 94.49 89.95
Mask2former [18] CVPR’22 94.16 80.46 91.08 98.68 86.88 96.07 74.05 94.76 90.25
PIDNet [19] CVPR’23 87.77 59.01 80.57 96.89 77.64 91.44 53.43 88.56 80.38
TransUNet [20] MIA’24 90.67 62.68 85.64 98.41 83.71 93.64 61.82 90.95 84.22
SAR DANet [13] CVPR’19 91.20 58.97 85.93 98.13 82.31 93.57 59.35 90.28 83.31
OCRNet [14] ECCV’20 86.38 38.70 78.36 96.18 71.34 89.52 41.07 83.54 74.19
CGNet [15] TIP’20 81.16 19.58 69.45 90.27 60.78 84.63 25.42 74.96 64.25
SegFormer [16] NeurIPS’21 88.81 46.92 81.99 97.57 76.75 91.51 49.55 86.73 78.41
SETR [17] CVPR’21 90.97 59.57 84.71 98.16 80.51 93.27 59.39 89.99 82.79
Mask2former [18] CVPR’22 91.34 58.52 86.40 93.29 80.08 93.20 62.69 89.49 81.92
PIDNet [19] CVPR’23 84.37 36.77 73.28 94.91 66.91 87.55 38.63 81.49 71.25
TransUNet [20] MIA’24 86.56 38.45 77.56 96.53 71.30 89.49 42.21 83.44 74.08
SAR & RGB MCANet [23] JAG’21 85.12 47.88 78.69 96.96 73.18 89.59 46.13 85.55 76.37
CEN [28] TPAMI’22 91.78 58.73 87.70 98.33 85.73 94.26 60.19 90.93 84.46
TokenFusion [53] CVPR’22 93.90 70.67 90.62 98.94 89.21 95.79 69.64 93.70 88.67
DDHRNet [24] JAG’22 79.78 13.85 69.85 93.73 62.17 84.73 24.59 73.75 63.88
CDDFuse [10] CVPR’23 96.59 83.02 94.59 99.11 93.30 97.61 80.65 96.46 93.32
CMX [26] TITS’23 96.43 82.59 94.37 99.38 93.72 97.58 81.08 96.44 93.30
MRFS [25] CVPR’24 96.48 83.68 94.48 99.39 93.84 97.65 82.24 96.60 93.57
FTransUNet [5] TGRS’24 84.69 35.21 76.51 96.40 71.11 88.56 36.55 82.35 72.78
ASANet [27] ISPRS’24 95.04 77.09 92.64 99.24 91.82 96.73 75.77 95.21 91.17
PAD (Ours) 97.09 86.12 95.38 99.47 94.89 98.04 84.38 97.16 94.59
TABLE IV: Ablation Experimental Results (%) on Main Modules on WHU-OPT-SAR.
ID Components IoU per Category aAcc mKappa mF1 mIoU
MSF ASF PSC Farmland City Village Water Forest Road Others
[…]
IV-C2 Evaluations on DDHR-SK
As shown in Table III, PAD exhibits superior boundary segmentation performance on the relatively simplified scenes of DDHR-SK. Compared with leading unimodal models (e.g., Mask2former-RGB [18], achieving 90.25% mIoU and 94.76% mF1, DANet-SAR [13], yielding 83.31% mIoU and 90.28% mF1) and the multi-modal baseline MRFS [25] (93.57% mIoU, 82.24% mKappa), PAD attains new state-of-the-art performance with 94.59% mIoU, 97.65% OA, 82.24% mKappa, and 96.60% mF1. For boundary-sensitive categories, PAD achieves 99.47% IoU on Water and 86.12% IoU on Road edges, surpassing MRFS by 2.44% and 0.08%, respectively.
Figure 7: Visualization Results on DDHR-SK. Four boundary-focused cases are shown: (a) coastal demarcation, (b) water–cropland interface, (c) vegetation–building complex, and (d) road network topology. PAD is compared with leading unimodal methods (Mask2former-RGB, DANet-SAR) and state-of-the-art multi-modal approaches (CDDFuse, CMX, MRFS). For each case, the panels display the RGB input, SAR coherence map, pseudo-color ground truth, and prediction maps from six methods. Critical boundary regions are highlighted with orange boxes; further analysis is provided in Section IV-C.
The visual results further reveal PAD’s capability in cross-modal boundary refinement. In Fig. 7(a), PAD accurately reconstructs periodically submerged coastal lines via SAR coherence analysis, whereas RGB models tend to confuse tidal sediments with terrestrial areas. In Fig. 7(b), PAD leverages SAR features to delineate water infiltration boundaries in irrigated fields, overcoming the over-segmentation issues common to optical sensors. Moreover, in Fig. 7(c) and Fig. 7(d), PAD achieves continuous road networks under vegetation occlusion and precise material boundaries in vegetation–building complexes, respectively.
IV-C3 Ablation Study
We performed comprehensive ablation studies on WHU-OPT-SAR to validate the contributions of the core components of our approach: Amplitude Spectrum Fusion module (ASF), Phase Spectrum Correction module (PSC), and Multi-Stage Fusion (MSF).
Main Modules
Eight model variants (see Table IV) were tested. The baseline variant (I), which replaces MSF with early fusion and substitutes the PAD (ASF+PSC) with pixel-wise additive fusion, achieves the lowest mIoU (54.98%). Enabling MSF alone (II) improves mIoU by 0.39%, while using ASF alone (III) attains 55.86% mIoU, and solely PSC (IV) adds a smaller gain. Combining MSF with either ASF (V) or PSC (VI) yields similar improvements (mIoU around 56.11–56.12%). Integrating all components achieves the best performance (56.26% mIoU), demonstrating the effective synergy of all modules.
TABLE V: Ablation Experiment Results (%) on Loss Function Weights on WHU-OPT-SAR.
Configuration IoU per Category aAcc Kappa mF1 mIoU
Farmland City Village Water Forest Road Others
Fixed $\lambda_{2}=0.1$
$\lambda_{1}=0.01$ 69.87 57.71 49.43 65.31 84.12 43.35 23.22 84.54 53.60 70.02 56.14
$\lambda_{1}=0.4$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
$\lambda_{1}=1.0$ 69.56 57.30 49.10 65.39 83.84 41.59 20.97 84.37 52.20 69.21 55.39
Fixed $\lambda_{1}=0.4$
$\lambda_{2}=0.01$ 69.83 56.80 48.67 65.63 84.19 43.17 22.82 84.52 53.54 69.76 55.87
$\lambda_{2}=0.1$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
$\lambda_{2}=1.0$ 69.83 57.74 49.37 64.95 83.95 43.45 22.50 84.46 53.87 69.84 55.97
Loss Function Weights
We conducted weight ablation studies (see Table V) to examine the impact of the trade-off between the amplitude consistency loss ($\mathcal{L}_{\mathrm{amp}}$) and the segmentation loss ($\mathcal{L}_{\mathrm{seg}}$). The best performance (mIoU = 56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$. Increasing $\lambda_{1}$ above 0.4 degrades performance, notably in under-represented classes, while an overly low $\lambda_{2}$ significantly diminishes performance, confirming that segmentation supervision must remain the dominant objective.
TABLE VI: Ablation Results (%) on the Initial Frequency Radius on WHU-OPT-SAR.
Configuration IoU per Category aAcc Kappa mF1 mIoU
Farmland City Village Water Forest Road Others
$radius=0.01$ 70.08 57.44 49.31 66.29 84.18 43.08 22.25 84.66 53.92 69.88 56.09
$radius=0.1$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
$radius=1$ 69.79 58.19 49.71 64.81 84.15 43.44 23.03 84.52 53.88 70.04 56.16
$radius=10$ 69.71 57.71 49.69 65.63 84.18 43.21 23.56 84.57 53.93 70.12 56.24
$radius=100$ 69.90 57.23 49.57 65.23 84.13 43.34 22.94 84.53 53.88 69.93 56.05
Initial Frequency Radius
The initial frequency radius parameter $r$, which controls the proportion of preserved low-frequency components in our phase-aware filtering, was varied (see Table VI). Our experiments show that $r=0.1$ achieves an optimal balance (56.26% mIoU). Values that overly suppress low-frequency information or over-smooth the features lead to inferior performance, though the method remains robust for $r\in[0.1,10]$.
IV-D Visualization at Distribution Level
To further analyze the feature distribution, we extracted multi-scale features from the RGB and SAR encoders, as well as from the fusion modules. After upsampling to the original resolution, the features were concatenated channel-wise. To balance computational efficiency with preserving discriminative information, a two-stage dimensionality reduction was applied: 1) random spatial sampling, followed by 2) principal component analysis (PCA), retaining 50 principal components (accounting for over 89% of the variance). The final visualization was generated using t-SNE [54] (with a perplexity of 30 and 1,000 iterations).
Refer to caption Refer to caption Refer to caption
(a) Before PAD Fusion (RGB branch) (b) Before PAD Fusion (SAR branch) (c) After PAD Fusion
Figure 8: t-SNE Visualization of Cross-modal Feature Distribution Evolution. (a) RGB encoder features exhibit spectral continuity but have blurred boundaries; (b) SAR encoder features, while strongly sensitive to geometry, are affected by speckle noise; (c) PAD-fused features display enhanced intra-class compactness and improved semantic separation.
As illustrated in Fig. 8, the PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability. Quantitatively, the mean intra-class standard variance is reduced by 20.39% for Road, 13.16% for Water, 6.46% for Forest, and 4.43% for Farmland, compared to the best unimodal baselines. In addition, significant increases in Euclidean distances between critical class pairs, such as a 75.75% increase in the City/Village pair (66.56 to 116.95) and a 13.29% increase for the Farmland/Forest pair (107.64 to 121.89), indicate that the fused features capture cross-modal discriminative patterns effectively.
Overall, these experimental results confirm that the PAD framework significantly improves classification accuracy and prediction consistency by effectively integrating multi-modal information and suppressing noise, while maintaining the discriminative spectral and geometric characteristics inherent in the data.
V Conclusion
In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land cover classification. Our frequency-domain analysis revealed that phase components are consistent across modalities, while amplitude spectra capture modality-specific details, providing the first quantitative physical prior for multi-source remote sensing fusion. PAD separates and aligns phase features while adaptively fusing amplitude components. The Phase Spectrum Correction (PSC) module resolves structural discrepancies, and the Amplitude Spectrum Fusion (ASF) module dynamically integrates SAR’s high-frequency details with RGB’s low-frequency context. Extensive experiments demonstrate that PAD outperforms existing methods in both accuracy and robustness. Future work will focus on improving phase alignment, particularly under significant geometric distortions due to imprecise cross-modal registration.

### v2

@@
 Abstract
-The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and the underutilization of spectral complementarity.
-Existing methods often fail to decouple shared structural features from modality-specific radiometric attributes, leading to feature conflicts and information loss.
-To address this issue, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-specific) components in the Fourier domain.
-Specifically, PAD consists of two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features through convolution-guided scaling to enhance geometric consistency, and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency details and low-frequency structures using frequency-adaptive multilayer perceptrons.
-This approach leverages SAR’s sensitivity to morphological features and RGB’s spectral richness.
+The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and underutilized spectral complementarity.
+Existing methods often fail to decouple shared structural features from modality-complementary radiometric attributes, causing feature conflicts and information loss.
+To address this, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-complementary) components in the Fourier domain, thus reinforcing shared structures while preserving complementary characteristics to improve fusion quality.
+Unlike prior approaches that overlook the distinct physical properties encoded in frequency spectra, PAD is the first to introduce explicit amplitude-phase decoupling for multi-modal fusion.
+Specifically, PAD comprises two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features via convolution-guided scaling to enhance geometric consistency; and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive multilayer perceptrons, leveraging SAR’s morphological sensitivity and RGB’s spectral richness.
 Extensive experiments on WHU-OPT-SAR and DDHR-SK datasets demonstrate state-of-the-art performance.
@@
 I Introduction
-Land cover classification (LCC) is a fundamental task in Earth observation that involves categorizing various surface materials, such as farmlands, forests, water bodies, and urban areas, using remote sensing imagery.
+Land cover classification (LCC) is a fundamental Earth observation task that categorizes surface materials, such as farmlands, forests, water bodies, and urban areas, using remote sensing imagery.
 This capability is essential for environmental monitoring, resource management, and urban development [1].
-In recent years, fusing Synthetic Aperture Radar (SAR) and RGB imagery has emerged as a promising approach to improve LCC performance.
-SAR provides robust structural information that is less susceptible to atmospheric and illumination variations, while RGB imagery offers rich spectral details, compensating for SAR-specific artifacts such as speckle noise [2, 3].
-Despite their complementary relation, a key challenge remains: the inherent imaging heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB), which leads to significant spectral misalignment and hinders the effective fusion of cross-modal features.
-Compared to conventional multi-modal fusion in natural images, integrating SAR and RGB data presents more pronounced radiometric and geometric disparities, limiting the performance of existing frameworks [4, 5, 6].
-Figure 1: SAR-RGB Spatial-Frequency Complementarity Analysis.
-Frequency-domain representations capture modality complementarity more effectively than spatial features.
-(a) Spatial discrepancies exhibit anisotropic distributions; (b) relative amplitude differences are pronounced in high-frequency regions; (c) phase discrepancies are minimal and clustered in low-frequency bands; (e)–(f) variations in sampling rates significantly affect the spatial discrepancy distribution.
-Analysis conducted on WHU-OPT-SAR (see Section III-A).
-Deep learning methods, particularly convolutional neural networks (CNNs), have achieved notable success in LCC tasks due to their powerful feature extraction capabilities [7].
-To mitigate modality heterogeneity in SAR-RGB fusion, most approaches employ dual-branch architectures to independently extract multi-scale features from each modality [8], followed by fusion via simple operations such as element-wise summation, mean aggregation, or channel-wise concatenation, before processing with a shared decoder or segmentation head.
-Although this fusion scheme relies on accurately modeling complementary features, the intrinsic heterogeneity between SAR and RGB often impedes the extraction of truly complementary representations.
-Recent studies, inspired by contrastive learning [9, 10, 11], have explored aligning highly correlated, modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts.
+Recently, fusing Synthetic Aperture Radar (SAR) and RGB imagery has emerged as a promising approach to enhance LCC performance: SAR provides robust structural information resilient to atmospheric and illumination variations, while RGB imagery offers rich spectral details that mitigate SAR artifacts such as speckle noise [2, 3].
+Figure 1: Motivation and SAR-RGB Spatial-Frequency Difference Analysis.
+(a) SAR and RGB imagery exhibit modality gaps due to distinct sensing mechanisms (backscattering vs.
+reflectance).
+(b) Fusion feature space comprises heterogeneous, shared, and complementary components; effective fusion suppresses heterogeneous features while enhancing shared and complementary patterns.
+(c)-(g) Frequency-domain analysis on WHU-OPT-SAR (see Section III-A): (c) spatial discrepancies exhibit anisotropic distributions; (d) relative amplitude differences concentrate in high-frequency regions; (e) phase discrepancies are minimal and cluster in low-frequency bands; (f)-(g) sampling-rate variations significantly affect spatial discrepancy distributions.
+Despite this complementarity, RGB-SAR fusion faces a key challenge: inherent imaging heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB) leads to significant spectral misalignment and hinders effective cross-modal feature integration (see Fig.
+1(a)).
+In SAR images, objects often appear as discrete backscatter responses, lacking fundamental visual cues, such as geometric contours, textures, and consistent shapes, that are readily available in RGB imagery [4].
+Compared to natural-image fusion, integrating SAR and RGB data presents pronounced radiometric and geometric disparities, limiting existing frameworks [5, 6, 7].
+Deep learning, particularly convolutional neural networks (CNNs), has achieved notable success in LCC by extracting powerful features [8, 9].
+To mitigate modality heterogeneity, most approaches use dual-branch architectures to extract multi-scale features independently, then fuse them via element-wise summation, mean aggregation, or channel-wise concatenation before a shared decoder or segmentation head [10].
+However, accurate modeling of complementary features is still impeded by SAR-RGB heterogeneity.
+Inspired by contrastive learning [11, 12, 13], recent works align modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts (see Fig.
+1(b)), yet spatial-domain modeling remains inefficient.
+To better capture RGB-SAR complementarity, some methods explore frequency-domain fusion [14, 15, 16].
 For example, Li et al.
-[12] considered phase as a modality-shared feature and used it as input in the intermediate network stage.
-However, they not only disregarded high-frequency amplitude patterns, which are essential for terrain delineation, thus limiting the effectiveness of the fusion, but also failed to systematically prove the complementarity of phase between the two modalities.
-Through comprehensive spectral analysis (Fig.
-1), we identify three critical frequency-domain priors that have been previously neglected: 1) Spatial-domain alignment (Fig.
-1(a)) is significantly hampered by anisotropic discrepancies between modalities, rendering direct pixel- or feature-level matching unreliable; 2) Existing phase-based strategies overlook the inevitable slight differences between different frequency bands (Fig.
-1(c)), leading to information loss; and 3) Amplitude mismatches between SAR and RGB modalities present different patterns in the high/low-frequency regions, most pronounced in the high-frequency band (Fig.
-1(b)), highlighting the need for frequency-aware difference modeling.
-Motivated by these insights, we propose a novel fusion method, Phase-Amplitude Decoupling (PAD).
-The central idea behind PAD is to disentangle SAR and RGB modalities into phase components, which capture modality-invariant structural information, and amplitude components, which encode modality-specific details.
-By explicitly aligning the shared phase representations, PAD enhances the extraction of complementary amplitude features, thereby reducing feature conflicts caused by cross-modal heterogeneity.
-Specifically, PAD comprises two key modules: 1) a Phase Spectrum Correction (PSC) module that ensures robust phase alignment across modalities, and 2) an Amplitude Spectrum Fusion (ASF) module that selectively integrates complementary amplitude cues for improved fusion performance.
-Our contributions are summarized fourfold:
-  • Through the systematic analysis of conventional registration datasets, we establish the frequency distribution law of SAR-RGB common patterns (phase) and unique patterns (amplitude) for the first time, providing an interpretable physical prior for multi-source remote sensing fusion.
-  • We propose a novel spectral decomposition method that orthogonally separates phase (modality-shared) and amplitude (modality-specific) features, effectively resolving feature conflicts in SAR-RGB fusion while maintaining geometric integrity.
-  • To address phase inconsistencies, the PSC module employs convolution-guided multiplicative scaling to align and enhance critical structural details, ensuring robust fusion.
-  • The ASF module leverages frequency-domain multilayer perceptrons to dynamically integrate high-frequency (local details) and low-frequency (global structure) components, capitalizing on the complementary strengths of SAR and RGB data for improved segmentation.
+[17] treat phase as modality-shared features but discard amplitude information essential for terrain delineation and do not systematically validate phase sharing.
+We conduct comprehensive spectral analysis (see Fig.
+1(c)-(g)) to reveal three critical frequency-domain priors: 1) SAR-RGB spatial-domain differences are anisotropic and susceptible to pixel-level mismatch (see Fig.
+1(c)); 2) average phase spectra differ minimally and mainly in low frequencies, indicating phase as shared information (see Fig.
+1(e)); and 3) amplitude mismatches exhibit distinct high/low-frequency patterns, most pronounced in the high-frequency band, highlighting the need for frequency-aware modeling (see Fig.
+1(d)).
+Motivated by these insights, we propose Phase-Amplitude Decoupling (PAD), which disentangles phase components capturing modality-invariant structures from amplitude components encoding modality-specific details.
+By explicitly aligning shared phase representations, PAD enhances complementary amplitude feature extraction and reduces cross-modal conflicts.
+PAD comprises two modules: Phase Spectrum Correction (PSC), which aligns and reinforces structural details via convolution-guided scaling; and Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive MLPs.
+Our contributions are fourfold:
+  • We systematically analyze registration datasets to establish frequency distributions of SAR-RGB common (phase) and unique (amplitude) patterns, offering interpretable physical priors for multi-source fusion.
+  • We introduce frequency-domain amplitude-phase decoupling for multi-modal fusion, orthogonally separating phase (modality-shared) and amplitude (modality-specific) features to resolve heterogeneity-induced conflicts.
+  • PSC employs convolution-guided multiplicative scaling to align and reinforce critical structural details for robust fusion.
+  • ASF uses frequency-domain MLPs to integrate complementary high-frequency (local details) and low-frequency (global structure) components, leveraging SAR’s morphological sensitivity and RGB’s spectral richness for improved segmentation.
 Figure 2: Flowchart of the PAD Framework.
-The framework comprises the (a) SCF module, (c) ASF module, and (d) PSC module.
-Registered RGB and SAR images, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are first input, and their features, $x_{RGB}$ and $x_{SAR}$, are extracted at each backbone stage before being fed into $n$ PAD fusion modules.
-Meanwhile, (b) FD and (e) FR denote the transformations from the spatial to the frequency domain and vice versa, respectively.
-Each stage’s final output is fed into a shared decoder.
-Here, $\mathcal{A}$ and $\mathcal{P}$ represent amplitude and phase, and $C$ (outer circle) indicates channel-wise concatenation.
+Registered RGB and SAR images, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are processed by asymmetric backbones to extract features $x_{\mathrm{RGB}}$ and $x_{\mathrm{SAR}}$ at each stage.
+These features are passed through $n$ PAD fusion modules, each comprising: (a) the SCF module, (c) the ASF module, and (d) the PSC module.
+(b) FD and (e) FR denote the spatial-to-frequency and frequency-to-spatial transformations, respectively.
+Each stage’s fused features are concatenated channel-wise (outer circle $C$) and fed into a shared decoder.
+Here, $\mathcal{A}$ and $\mathcal{P}$ represent amplitude and phase spectra.
 II Related Work
 II-A Multi-modal Fusion for Land Cover Classification
-While single-modality approaches [13, 14, 15, 16, 17, 18, 19, 20] have achieved reasonable performance in homogeneous data scenarios, their inability to leverage cross-modal synergies limits their discriminative power in complex environments.
-SAR-RGB fusion methods generally follow two paradigms: knowledge transfer via domain adaptation and hierarchical feature integration.
-In this section, we analyze their technical evolution and inherent limitations.
+Single-modality approaches [18, 19, 20, 21, 22, 23, 24, 25] perform well on homogeneous data but cannot exploit cross-modal synergies in complex environments.
+SAR-RGB fusion methods generally follow two paradigms: cross-modal transfer learning and hierarchical feature integration.
 II-A1 Cross-Modal Transfer Learning
-Transfer learning [21] seeks to adapt knowledge from one modality to another.
-However, the significant differences between SAR and RGB data often degrade learned representations, limiting their applicability in multi-modal tasks.
+Transfer learning [26] adapts knowledge from one modality to another, but the pronounced differences between SAR and RGB data often degrade learned representations, limiting multi-modal applicability.
 II-A2 Hierarchical Feature Fusion
-Feature fusion strategies have progressed through three developmental phases: a) Basic Fusion Operators: Early methods utilized basic operations such as concatenation or weighted summation [22].
-However, these approaches are prone to modality dominance and information redundancy in complex scenes.
-b) Attention-Guided Fusion: Spatial-channel attention mechanisms were introduced to address these issues [23, 24].
-For instance, MCANet introduces cross-modal attention (MCAM) to model inter-modality dependencies, MRFS [25] develops progressive cycle attention (PC-Att) for infrared-visible fusion, and CMX [26] introduces bidirectional channel-spatial attention to refine cross-modal features.
-However, most existing attention paradigms have high computational costs.
-c) Complementarity-Aware Fusion: Recent state-of-the-art approaches explicitly model modality-specific and modality-shared components.
-ASANet [27] uses semantic focusing modules (SFM) to emphasize complementary features, CDDFuse [10] applies contrastive learning to disentangle private and shared representations, and CEN [28] proposes a method that achieves complementary perceptual fusion through dynamic channel exchange and parameter-sharing-independent batch normalization.
-Furthermore, recent work [12] pioneers phase-structure alignment by leveraging phase information as a cross-modal invariant feature.
-Despite these advancements, spatial-domain methods still struggle with geometric misalignment and computational inefficiency.
+Feature fusion has evolved through three stages: a) Basic Fusion Operators: Early methods used concatenation or weighted summation [27], but these are prone to modality dominance and redundancy.
+b) Attention-Guided Fusion: Spatial-channel attention mechanisms address these issues [28, 29].
+For example, MCANet’s cross-modal attention (MCAM) models inter-modality dependencies; MRFS [30] introduces progressive cycle attention for infrared-visible fusion; and CMX [31] employs bidirectional channel-spatial attention.
+However, these paradigms incur high computational costs.
+c) Complementarity-Aware Fusion: Recent approaches explicitly disentangle shared and modality-specific features.
+ASANet [32] uses semantic focusing modules to emphasize complementary cues; CDDFuse [12] applies contrastive learning to separate private and shared representations; and CEN [33] leverages dynamic channel exchange with independent batch normalization.
+Despite these advances, spatial-domain methods still struggle with geometric misalignment and inefficiency.
 II-B Frequency Representation Learning in Vision Tasks
-Frequency analysis offers powerful tools to address challenges in spatial fusion [29, 30, 31]. Three primary application directions are discussed below:
-II-B1 Phase as a Modality-Invariant Representation
-Decoupling amplitude and phase spectra has proven effective in domain generalization (DG), where models trained on multiple source domains must generalize to unseen target domains.
-In DG, phase spectra encode structural semantics (e.g., object shapes and edges), while amplitude spectra capture modality-specific styles (e.g., illumination and texture patterns) [32, 33].
-Early DG methods [34, 35] attempted to disentangle style and content by normalizing feature statistics (mean/variance), assuming that style corresponds to statistical variations.
-However, such heuristic separations can distort structural content.
-Recent approaches employ Fourier transforms as explicit style-content separators.
-For example, amplitude-phase recombination [36] swaps the amplitude spectrum of a source image with that of a target domain while preserving the source phase, thereby synthesizing cross-domain training samples.
-This strategy compels networks to focus on phase-centric structural features, improving robustness to style shifts.
-Building on this, decompose-adjust-compose (DAC) [37] integrates Phase-Amplitude Decoupling with feature normalization to dynamically balance style and content in the frequency domain.
-These methods highlight a critical advantage of spectral decomposition: phase alignment preserves structural consistency across modalities, while amplitude manipulation enables adaptive style fusion.
-II-B2 Frequency-Driven Global Context Modeling
-Beyond style-content separation, the Fourier transform’s global receptive field has been exploited to overcome two key limitations of spatial-domain networks: restricted long-range dependency modeling and the quadratic complexity of self-attention.
-For instance, fast Fourier convolution (FFC) [38] introduced hybrid spatial-frequency processing with parallel branches, capturing global patterns via spectral operations while preserving local details through convolution.
-These principles have also been applied to optimize attention mechanisms.
-FFTformer [39] replaces dot-product attention with spectral element-wise multiplication, reducing complexity from $O(N^{2})$ to $O(N\log N)$ while maintaining global interactions.
-Similarly, GFNet [40] and AFNO [41] employ learnable frequency filters as token mixers, achieving significant speedups compared to efficient self-attention variants.
-These methods demonstrate that global context aggregation in the frequency domain is not only computationally efficient but also capable of fusing multi-scale information.
-II-B3 Frequency Priors as Inductive Biases
-Recent studies have shown that frequency-domain representations can be explicitly designed as task-specific inductive biases [42, 43, 44].
-For DG, DFF [45] exploits the assumption that low frequencies control cross-domain structural consistency, introducing a frequency-domain dynamic filtering mechanism to amplify domain-invariant components while suppressing high-frequency variations.
-In dense prediction tasks, FreqFusion [46] applies adaptive frequency filtering: an Adaptive Low-Pass Filter (ALPF) generator smooths high-frequency perturbations to homogenize intra-class features, while an Adaptive High-Pass Filter (AHPF) generator restores critical boundary information lost during downsampling.
-This strategy encodes spectral priors that low frequencies ensure semantic consistency, while high frequencies enhance boundary discriminability.
-In the context of RGB-NIR denoising, FCENet [47] utilizes a Frequency Dynamic Selection Module (FDSM) to select useful frequency components across modalities and a Frequency Exhaustive Fusion Module (FEFM) to refine shared and modality-specific spectral features.
-These approaches demonstrate that decoupling and resetting frequency components effectively encode task-related physical principles, such as structural invariance and boundary sensitivity, injecting strong inductive biases into deep networks.
+Frequency representation learning has emerged as a powerful tool across various vision tasks due to its capacity for global modeling and structural decomposition.
+In this subsection, we review three relevant areas:
+II-B1 Spectral Decomposition for Structure-Style Separation
+Fourier-based decomposition isolates phase (structural semantics) and amplitude (appearance) for domain generalization and style transfer [34, 35].
+Early methods used feature statistics for style normalization [36, 37], while recent works employ Fourier transforms for clearer separation, enabling content-preserving augmentation by swapping amplitudes across domains [38, 39].
+II-B2 Frequency-Driven Global Modeling
+Fourier transforms provide a global receptive field useful for efficient attention and context modeling.
+Fast Fourier Convolution (FFC) [40] and FFTformer [41] leverage spectral processing to replace or accelerate spatial attention.
+Spectral token mixing has also been effective in lightweight transformers [42, 43].
+II-B3 Frequency-Based Inductive Biases in Dense Prediction
+Spectral priors enhance dense prediction tasks.
+DFF [44] uses frequency-domain dynamic filtering to amplify invariant low-frequency features, improving cross-domain generalization.
+FreqFusion [45] designs learnable high- or low-pass filters for semantic consistency and boundary refinement.
+FCENet [14] employs frequency-selective modules for cross-modal spectral correlations in RGB-NIR denoising.
+In summary, frequency-based representations offer a flexible, physically interpretable design space for vision tasks; yet, few frameworks target task-specific fusion in remote sensing.
+We address this gap in the next section.
+II-C Frequency-Domain Fusion Methods
+Applying frequency representations to multi-modal fusion is recent.
+Shi et al.
+[15] introduce Frefusion, using spectral attention to merge modalities in the frequency domain.
+Frequency spectra naturally decompose signals into phase (modality-invariant structures) and amplitude (modality-specific details), offering a principled fusion framework.
+However, existing methods often transplant spatial operators, MLPs [16], convolutions [14], or attention [15], into the spectral domain, modulating magnitude while neglecting the distinct roles of phase and amplitude.
+Consequently, shared information may be preserved, but complementary cues are suppressed.
+In contrast, our PAD framework explicitly decouples phase and amplitude spectra.
+Through systematic spectral analysis (see Section III-A), we show that SAR and RGB share strong phase consistency while amplitude captures complementary patterns.
+PAD introduces (1) a Phase Spectrum Correction module for robust structural alignment and (2) an Amplitude Spectrum Fusion module with frequency-adaptive refinement.
+This decomposition-correction-fusion pipeline distinguishes PAD from previous FFT-based approaches by providing a physically interpretable strategy for multi-modal fusion.
+The two most related works, FeINFN [16] and SFINet++ [46], also leverage amplitude and phase in the frequency domain for high-frequency detail enhancement and global context modeling.
+However, they lack a physically grounded interpretation of amplitude-phase roles and do not explicitly decouple shared versus modality-specific information.
+In contrast, PAD is the first framework to formulate and exploit amplitude-phase decomposition for principled alignment and complementary integration across modalities.
 III Proposed Method
-This paper introduces a multi-modal fusion framework based on Phase-Amplitude Decoupling (PAD) for land cover classification using SAR and RGB images.
-PAD effectively integrates complementary information from both modalities while mitigating heterogeneous feature conflicts that often arise with traditional spatial fusion techniques.
-The overall architecture, depicted in Fig.
-2, consists of three main modules: the Amplitude Spectrum Fusion (ASF) module, the Phase Spectrum Correction (PSC) module, and the Spatial Complementary Fusion (SCF) module.
+To address the fundamental conflict between modality-invariant geometry and modality-specific appearance in SAR-RGB fusion, we introduce Phase-Amplitude Decoupling (PAD), a frequency-aware multi-modal fusion framework.
+The overall architecture, shown in Fig.
+2, comprises three modules: the Spatial Complementary Fusion (SCF) module, which preserves fine-grained spatial priors; the Phase Spectrum Correction (PSC) module, which aligns and reinforces shared structural details; and the Amplitude Spectrum Fusion (ASF) module, which adaptively integrates complementary high- and low-frequency amplitude cues.
+Unlike previous methods that apply spectral operations uniformly, PAD explicitly decouples and recombines phase and amplitude with dedicated modules to enable physics-consistent fusion.
 Input : SAR/RGB image pairs $\{I_{\mathrm{SAR}}^{(i)},I_{\mathrm{RGB}}^{(i)}\}_{i=1}^{N}$
@@
 Algorithm 1 Spectral Difference Analysis.
-III-A Frequency-domain Analysis for Modality Decoupling
-To validate the physical significance of frequency-domain decoupling in multi-modal data, we conducted a systematic spectral analysis on WHU-OPT-SAR.
-This analysis includes visualizations of the averaged difference spectra (Fig.
-1(a)–(c)) computed using Algorithm 1 to compare the differences between two modalities in different domains.
+III-A Frequency-domain Priors for Amplitude-Phase Decoupling
+To validate the physical significance of phase-amplitude decoupling and reveal the shared and complementary relations between RGB and SAR, we conducted a systematic spectral analysis on WHU-OPT-SAR dataset, the largest RGB-SAR segmentation benchmark.
+Firstly, we visualized the averaged difference spectra (see Fig.
+1(c)-(e)) using Algorithm 1 to compare modality differences in spatial, amplitude, and phase domains.
+This analysis shows that frequency-domain representations capture cross-modal sharing and complementarity more effectively than spatial representations.
 The difference spectra are defined as:
@@
 $\mathrm{APPD}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\arg\exp j(\mathcal{P}_{\mathrm{SAR}}^{(i)}-\mathcal{P}_{\mathrm{RGB}}^{(i)})\right|,$ (3)
-where $I_{\mathrm{SAR}}$, $\mathcal{A}_{\mathrm{SAR}}$, and $\mathcal{P}_{\mathrm{SAR}}$ denote the pixel values, amplitude, and phase spectra of SAR images, respectively; $I_{\mathrm{RGB}}$, $\mathcal{A}_{\mathrm{RGB}}$, and $\mathcal{P}_{\mathrm{RGB}}$ correspond to those of the RGB images; $\arg(\cdot)$ computes the angle of a complex number; ||$\cdot$|| ⋅ | denotes the absolute value; and $\epsilon$ is a small constant to prevent division by zero.
-By leveraging the periodicity of complex exponentials, APPD maps phase differences into the principal value range $[-\pi,\pi)$, eliminating abnormally large differences caused by phase wrapping (e.g., a true phase difference of $2\pi$ is corrected to 00).
-Figure 1 reveals three key observations: (1) Spatial differences (RSD) exhibit anisotropic distributions; (2) Amplitude differences (RAD) predominantly emerge in high-frequency regions; (3) Phase differences (APPD) remain minimal with slight low-frequency clustering.
-The frequency-domain metrics (RAD, APPD) demonstrate superior specificity compared to spatial-domain RSD, confirming that spectral decomposition better captures cross-modal complementarity.
-These analyses lead to two fundamental insights that drive our architecture design: a) Phase Sharing: The consistently low APPD values indicate strong phase consistency between modalities.
-This suggests that phase spectra contain shared information that can guide cross-modal learning.
-Our PSC Module thus performs pixel-level phase alignment while preserving this inherent consistency.
-b) Amplitude Complementarity: The RAD distribution reveals frequency-dependent amplitude divergence, with high-frequency components showing greater discrepancy than low-frequency ones.
-This localized spectral pattern motivates our ASF Module, which adjusts the frequency radius dynamically to handle frequency-specific amplitude variations.
-These phenomena can be physically interpreted based on the inherent characteristics of each imaging modality.
-Specifically, phase discrepancies manifest differently across frequency bands.
-Low-frequency disparities primarily stem from geometric distortions in imaging, leading to overall deviation between the two modalities that cannot be registered, whereas high-frequency variations arise from SAR speckle noise affecting fine details.
-On the other hand, amplitude disparities emerge from varying sensitivities to different ground objects, with low/high frequency contrasts reflecting modality-specific responses to large/small scale structures (e.g., farmlands and forests vs.
-roads).
-We further validate our spectral analysis through sensitivity experiments on sampling rates (Fig.
-1(e)–(f)).
-By analyzing the mean and variance of the difference spectra, we found that both RAD and APPD metrics are largely invariant to sampling rate, confirming that the observed phase-amplitude complementarity originates from intrinsic modality properties, rather than from sampling artifacts.
-TABLE I: Statistical Analysis Results for the Average Periodic Phase Difference Spectrum.
-Metric Low Freq High Freq Whole Image
-Shapiro-Wilk (Stat) 1.0044 1.0050 1.0131
-Shapiro-Wilk (p) 1.0 1.0 1.0
+where $I$, $\mathcal{A}$, and $\mathcal{P}$ denote pixel values, amplitude spectra, and phase spectra for SAR and RGB images; $\arg(\cdot)$ computes the complex angle; and $\epsilon$ prevents division by zero.
+By leveraging the periodicity of complex exponentials, APPD maps phase differences into $[-\pi,\pi)$, correcting wrap-around artifacts (e.g., $2\pi\to 0$).
+Figure 1(c)-(e) reveals three priors: 1) Spatial differences (RSD) exhibit anisotropic distributions.
+2) Amplitude differences (RAD) concentrate in high-frequency regions.
+3) Phase differences (APPD) remain minimal, with slight low-frequency clustering.
+These frequency-domain metrics demonstrate superior explainability over spatial RSD, confirming that spectral decomposition better captures cross-modal specificity.
+Physically, low-frequency phase disparities stem from geometric distortions (e.g., side-looking SAR vs.
+orthographic RGB), while high-frequency variations arise from SAR speckle noise.
+Amplitude disparities reflect modality-specific sensitivity to object scales: low frequencies encode large structures (farmlands, forests) and high frequencies capture fine details (roads).
+These analyses yield two key insights that drive our architecture design: 1) Phase Sharing: Consistently low APPD values indicate strong phase consistency.
+The PSC module thus performs pixel-level alignment while preserving this shared information.
+2) Amplitude Complementarity: RAD’s frequency-dependent divergence motivates ASF’s dynamic frequency-radius adjustment to handle amplitude variations.
+Secondly, we conducted sensitivity experiments on sampling rates (see Fig.
+1(f)-(g)) to evaluate the stability of our spectral metrics.
+By downsampling input images at various ratios and computing the mean and variance of RSD, RAD, and APPD across the dataset, we observed that RAD and APPD remain nearly constant, whereas RSD exhibits larger fluctuations.
+This invariance of RAD and APPD to sampling artifacts confirms that frequency-domain complementarity arises from intrinsic modality characteristics rather than from individual changes.
+TABLE I: Statistical Analysis of the Average Periodic Phase Difference Spectrum.
+Metric Low Frequency High Frequency Whole Image
+Shapiro-Wilk Statistic 1.0044 1.0050 1.0131
+Shapiro-Wilk p-value 1.0000 1.0000 1.0000
 Skewness -0.0521 0.0020 -0.0265
@@
 Figure 3: Statistical Analysis of APPD.
-(a) APPD spectrum; (b)–(d) Frequency histograms with Gaussian fits for low-frequency (LF, radius = 0.5 $\times$ half-diagonal), high-frequency (HF), and full spectra (ALL); (e)–(f) Mean-variance trajectories across sampling rates.
-The HF, LF, and ALL spectra show near-Gaussian invariance, while LF/ALL deviates slightly due to the dominant structural noise in the LF range.
-These results suggest that phase, especially HF, can serve as a robust shared feature.
-To assess the statistical consistency of cross-modal phase spectra, we conducted normality tests (Shapiro-Wilk), along with skewness and kurtosis analyses (Table I).
-Frequency histograms and mean-variance trajectories (Fig.
-3) indicate that the frequency difference spectra follow a near-Gaussian distribution, particularly for high-frequency components, whereas low-frequency components exhibit minor systematic deviations, likely due to structural noise.
-Figure 4: Visualization of Spatial and Frequency Spectrum Differences in a Scenario with Missing RGB Regions.
-(a)–(b) show RGB images with missing regions; (d)–(f) present the difference spectrum when the missing regions are included; and (g)–(i) display the difference spectrum after excluding the missing regions.
-(c) indicates that variation in RSD is significantly higher than in RAD and APPD, demonstrating that frequency-domain representations are more robust to missing data.
-Figure 5: Visualization of Spatial and Frequency Spectrum Differences in a Cloud-Obscured Observations Scenario.
-(a)–(c) present an example of cloud-obscured observations, with cloud masks generated using RGB, SAR, and NIR images from WHU-OPT-SAR [48].
-(d) compares the stability of RSD, RAD, and APPD between cloud-covered and cloud-free regions.
-Numbers next to the arrow indicates the value of decline.
-We also evaluate phase spectrum invariance under challenging conditions through two controlled experiments: (1) partial RGB data missing (Fig.
-4) and (2) cloud-obscured observations (Fig.
+(a) APPD spectrum.
+(b)-(d) Frequency histograms with Gaussian fits for low-frequency (LF; radius = 0.5 $\times$ half-diagonal), high-frequency (HF), and all spectra (ALL).
+(e)-(f) Mean-variance trajectories across sampling rates.
+All spectra exhibit near-Gaussian invariance, with LF and ALL showing slight deviations due to structural noise in the low-frequency range, indicating that phase, particularly HF, can serve as a robust shared feature.
+Thirdly, to assess the statistical consistency of cross-modal phase spectra, we performed normality tests (Shapiro-Wilk) and calculated skewness and kurtosis for APPD values at each frequency band (see Table I).
+Frequency histograms and mean-variance trajectories (see Fig.
+3) demonstrate that APPD distributions closely follow a Gaussian profile, especially in high-frequency regions.
+Low-frequency components show minor deviations, likely due to structural noise, but overall the near-Gaussian behavior validates phase spectra as a reliable shared feature.
+In a word, phase is so similar between RGB and SAR that it is almost unaffected by individual differences.
+Figure 4: Visualization of Spatial and Frequency Spectrum Differences under Missing RGB Regions.
+(a)-(b) RGB images with missing regions; (c) variation in RSD, RAD, and APPD, showing RSD’s greater sensitivity to missing data; (d)-(f) difference spectra including missing regions; (g)-(i) difference spectra excluding missing regions.
+These results demonstrate that frequency-domain metrics (RAD, APPD) are more robust to missing data than spatial-domain RSD.
+Figure 5: Visualization of Spatial and Frequency Spectrum Differences under Cloud Obscuration.
+(a)-(c) Example observations with cloud masks derived from RGB, SAR, and NIR channels in WHU-OPT-SAR [47].
+(d) Comparison of RSD, RAD, and APPD stability between cloud-covered and cloud-free areas; arrows denote the value of decline.
+Finally, we evaluated phase invariance under two challenging scenarios: 1) partial RGB data removal (see Fig.
+4) and 2) cloud-obscured observations (see Fig.
 5).
-In the missing RGB scenario, RSD is highly sensitive to data incompleteness (Fig.
-4(d) vs.
-(g)), while phase differences remain consistent (Fig.
-4(f) vs.
-(i)).
-Similarly, in cloud-affected cases, the APPD variance ($\Delta\approx 10^{-0.037}\%$) is negligible compared to that of RSD ($\Delta\approx 10^{5.81}\%$) and RAD ($\Delta\approx 10^{0.58}\%$) as shown in Fig.
+In the former, RSD increased dramatically when RGB pixels were missing (compare Fig.
+4(d) and (g)), while APPD remained stable (compare Fig.
+4(f) and (i)).
+In the latter, APPD variance ($\Delta\approx 10^{-0.037}\%$) was negligible compared to that of RSD ($\Delta\approx 10^{5.81}\%$) and RAD ($\Delta\approx 10^{0.58}\%$) as shown in Fig.
 5(d).
-Collectively, these results demonstrate that phase information (particularly in the high-frequency region) exhibits robust consistency between SAR and RGB modalities, supporting its role as a shared feature to mitigate inter-modal conflicts and leverage complementary information in multi-modal frameworks.
+These results confirm that phase information, particularly at high frequencies, retains robust consistency across modalities, underscoring its role as a shared feature to mitigate inter-modal conflicts and enhance complementary fusion.
 III-B Overall Framework
@@
 2, the proposed PAD framework integrates SAR and RGB modalities via a dual-branch encoder and a shared decoder, with its primary innovation in cross-modal feature fusion through phase-amplitude decoupling.
-The encoder utilizes asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$ with independent parameters to preserve modality-specific attributes.
+The encoder employs asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$, each with independent parameters to preserve modality-specific attributes.
 Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave scattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images.
-Both backbones produce hierarchical feature maps at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$, such that $i\in\{0,1,2,3\}$:
+Both backbones produce hierarchical feature maps at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$ for $i\in\{0,1,2,3\}$:
 $x_{RGB}^{i}=\mathcal{B}_{\mathrm{RGB}}^{(i)}\left(I_{\mathrm{RGB}}^{i}\right),\quad x_{SAR}^{i}=\mathcal{B}_{\mathrm{SAR}}^{(i)}\left(I_{\mathrm{SAR}}^{i}\right).$ (4)
-Initially, the SCF module (Eq.
-5) processes the bimodal features by computing pixel-wise fusion weights through spatial attention gating.
-This operation preliminarily integrates complementary features in the spatial domain and reduces feature dimensionality, thereby constraining the solution space for subsequent frequency-domain operations.
+Although frequency-domain fusion is more robust to modality differences, it can discard rich local interactions in the spatial domain.
+Therefore, we first introduce a lightweight SCF module to retain fine-grained spatial priors.
+The SCF module (see Eq.
+5) processes the bimodal features by computing pixel-wise fusion weights via spatial attention gating.
+This operation preliminarily integrates complementary features in the spatial domain and reduces feature dimensionality, constraining the solution space for subsequent frequency-domain operations:
 $X^{i}=\mathrm{SCF}\left(x_{RGB}^{i},x_{SAR}^{i}\right),$ (5a)
@@
 $x=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1}^{\prime},x_{2}^{\prime}\right)\right).$ (5d)
-The fused features are then transformed via FD into separate amplitude and phase spectra:
+The fused feature $X^{i}$ is then transformed via FD into separate amplitude $\mathcal{A}^{i}$ and phase $\mathcal{P}^{i}$ spectra:
 $\mathcal{A}^{i},\mathcal{P}^{i}=\mathrm{FD}\left(X^{i}\right).$ (6)
-The amplitude spectrum $\mathcal{A}^{i}$ is enhanced by the ASF module, which distinguishes high-frequency details from low-frequency structures using a trainable frequency radius $r$:
-$\mathcal{A^{\prime}}^{i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right).$ (7)
-Simultaneously, the phase spectrum is corrected via multiplicative scaling in the PSC module:
+The amplitude spectrum $\mathcal{A}^{i}$ is enhanced by the ASF module, which uses a trainable frequency radius $r$ to distinguish and reinforce high-frequency details while preserving low-frequency structures:
+$\mathcal{A^{\prime}}^{i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right),$ (7)
+where the PSC module applies multiplicative scaling to correct the phase:
 $\mathcal{P^{\prime}}^{i}=\mathrm{PSC}\left(\mathcal{P}^{i}\right).$ (8)
-Finally, FR reconstructs the features, yielding the final fusion output:
+Finally, FR reconstructs the fused features:
 $\bm{F^{\prime}}^{i}=\mathrm{FR}\left(\mathcal{A^{\prime}}^{i},\mathcal{P^{\prime}}^{i}\right).$ (9)
-The shared decoder, $\mathcal{D}$, aggregates the multi-stage fused features to generate high-precision land-cover segmentation maps:
+The shared decoder $\mathcal{D}$ then aggregates the multi-stage fused features to generate high-precision land-cover segmentation maps:
 $\bm{\hat{Y}}=\mathcal{D}\left(\bm{F^{\prime}}^{0},\bm{F^{\prime}}^{1},\bm{F^{\prime}}^{2},\bm{F^{\prime}}^{3}\right).$ (10)
@@
 To achieve physically disentangled representations in the spectral domain, we propose a bidirectional transformation framework comprising FD and its inverse, FR.
-First, the input image $X\in\mathbb{R}^{H\times W}$ is transformed from the spatial domain to the frequency domain using a 2D real-valued fast Fourier transform (rFFT) [49], yielding the spectrum $\mathcal{F}\in\mathbb{C}^{H\times W}$:
-$\mathcal{F}\left(u,v\right)=\sum_{x=0}^{W-1}\sum_{y=0}^{H-1}X\left(x,y\right)\exp\left(-i2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
-where $(u,v)$ represents frequency-domain coordinates.
-Exploiting the conjugate symmetry of real-valued signals, only the non-redundant region (with $20\leq u\leq W/$ 2 and $0\leq v<H$) is retained.
-Next, an FFT Shift operation relocates low-frequency components to the center of the spectrum:
+First, the input image $X\in\mathbb{R}^{H\times W}$ is transformed from the spatial domain to the frequency domain using a 2D real-input fast Fourier transform (rFFT) [48], yielding the spectrum $\mathcal{F}\in\mathbb{C}^{H\times(\lfloor W/2\rfloor+1)}$:
+$\mathcal{F}\left(u,v\right)=\sum_{y=0}^{H-1}\sum_{x=0}^{W-1}X\left(y,x\right)\exp\left(-i2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
+where $(u,v)$ are frequency-domain coordinates.
+Exploiting the conjugate symmetry of real-valued signals, we retain only the non-redundant region ($0\leq u\leq\lfloor W/2\rfloor$, $0\leq v<H$).
+Next, an FFT shift relocates low-frequency components to the spectrum center:
 $\mathcal{F}_{\mathrm{shift}}\left(u,v\right)=\mathcal{F}\left(u\oplus\frac{W}{2},v\oplus\frac{H}{2}\right),$ (12)
-where $\oplus$ denotes modulo addition. Finally, the amplitude and phase spectra are derived as:
+where $\oplus$ denotes modulo addition. Finally, we derive the amplitude and phase spectra as
 $\mathcal{A}=|\mathcal{F}|,\quad\mathcal{P}=\angle\mathcal{F},$ (13)
-with $\mathcal{A}\in\mathbb{R}^{H\times W}$ capturing radiometric properties and $\mathcal{P}\in\mathbb{R}^{H\times W}$ preserving structural details.
-This decoupling enables separate learning of SAR scattering features and RGB structural priors.
+where $\mathcal{A},\mathcal{P}\in\mathbb{R}^{H\times(\lfloor W/2\rfloor+1)}$ capturing radiometric and structural information, respectively.
+This decoupling allows the model to focus on modality-specific representations: SAR phase carries scattering patterns sensitive to physical surface characteristics, whereas RGB amplitude captures fine-grained textures and semantic content.
+By separating them, the network avoids early-stage interference between heterogeneous modalities, enhancing feature discriminability before integration.
 III-D Phase Spectrum Correction Module
+While previous works have assumed perfect phase alignment [17], our spectral analysis (see Fig.
+1(e)) reveals that low-frequency geometric inconsistencies persist, necessitating a learnable, spatially aware correction mechanism.
+Inspired by channel attention, we propose a lightweight Phase Spectrum Correction (PSC) module tailored for frequency-domain phase alignment.
 As illustrated in Fig.
-1(c), minor misalignments in the low-frequency phase spectra between SAR and RGB images, attributable to inherent geometric distortions (e.g., side-looking SAR vs.
+1(e), minor misalignments in the low-frequency phase spectra between SAR and RGB, attributable to geometric distortions (e.g., side-looking SAR vs.
 RGB orthographic projection), can compromise spatial consistency and boundary precision.
-Traditional phase alignment methods relying on rigid transformations are inadequate for non-rigid distortions.
-Thus, we propose a lightweight PSC module that performs adaptive geometric correction using learnable multiplicative masks.
-Given an input phase spectrum tensor $P\in\mathbb{R}^{C\times H\times W}$, the module first compresses the channel dimension via a $1\times 1$ convolution (reducing $C$ to $C/r$, where $r\in\{4,8\}$) and applies a Leaky ReLU activation [50].
-A subsequent $1\times 1$ convolution produces a single-channel Phase Correction Mask (PCM) constrained to $(0,1)$ by a Sigmoid function.
-Pixel-wise multiplicative scaling is then applied as follows:
-$P^{\prime}=P\odot\left(1+\text{PCM}\right),$ (14)
-where $\odot$ denotes the Hadamard product.
-This design limits the maximum scaling factor to 2, ensuring the preservation of phase periodicity while requiring significantly fewer parameters than conventional $3\times 3$ convolutions.
+Traditional rigid registration methods are inadequate for such non-rigid distortions.
+Thus, the PSC module performs adaptive geometric correction using learnable multiplicative masks.
+Unlike the global channel scaling in traditional attention mechanisms, the proposed PSC estimates a spatially varying modulation mask via a stack of $d$ consecutive $1\times 1$ convolutions, enabling localized, fine-grained phase adjustment.
+Given an input phase-spectrum tensor $P\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$, we first compress the channel dimension via $d$ $1\times 1$ convolution (reducing $C$ to $C/r$, $r\in\{4,8\}$) followed by leaky linear rectification function (Leaky ReLU) [49].
+A second $1\times 1$ convolution produces a single-channel Phase Correction Mask (PCM), constrained to $(0,1)$ by a Sigmoid:
+$\mathrm{PCM}=\sigma\left(\mathrm{Conv}_{d}^{1\times 1}\left(\delta\left(\cdots\delta\left(\mathrm{Conv}_{1}^{1\times 1}\left(P\right)\right)\cdots\right)\right)\right),$ (14)
+where $d$ denotes the number of stacked convolutions and $\delta(\cdot)$ is Leaky ReLU. Pixel-wise multiplicative scaling is then applied:
+$P^{\prime}=P\odot\left(1+\mathrm{PCM}\right),$ (15)
+where $\odot$ is the Hadamard product. This formulation enables localized, fine-grained phase adjustment while preserving phase periodicity and incurs minimal computational overhead.
 III-E Amplitude Spectrum Fusion Module
-Our analysis in Fig.
-1(b) indicates that the spectral discrepancy between SAR and RGB modalities is primarily concentrated in the medium-to-high frequency components of the amplitude spectra.
-To exploit this prior while maintaining gradient stability, we design an adaptive spectrum fusion module with three sequential operations.
-First, we compute a normalized frequency distance metric:
+Previous frequency-domain methods often apply fixed high- or low-pass filters or spatially uniform spectral attention, failing to adapt to content-specific frequency distributions.
+In contrast, our analysis in Fig.
+1(d) shows that the spectral discrepancy between SAR and RGB modalities is primarily concentrated in the medium-to-high frequency components of the amplitude spectra.
+To exploit this prior while maintaining gradient stability, we design the ASF module to employ a trainable frequency radius for dynamic separation and enhancement of complementary frequency bands via three sequential operations.
+First, we compute a normalized frequency distance metric for each spectral component $i$:
 $\mathrm{Dist}(i)=\frac{\sqrt{\left(X_{i}-X_{\mathrm{center}}\right)^{2}+\left(Y_{i}-Y_{\mathrm{center}}\right)^{2}}}{\sqrt{X_{\mathrm{max}}^{2}+Y_{\mathrm{max}}^{2}}},$ (16)
-which quantifies component positions relative to the spectrum center $(X_{\mathrm{center}},Y_{\mathrm{center}})$.
-A trainable boundary radius $r$, constrained via a Sigmoid activation to yield $R=\sigma(r)$, dynamically partitions the spectrum into low- and high-frequency regions.
-To maintain differentiability, a soft partition is employed:
+where quantifies component positions relative to the spectrum center $(X_{\mathrm{center}},Y_{\mathrm{center}})$.
+A trainable boundary radius $r$, constrained via a Sigmoid activation to yield $R=\sigma(r)\in[0,1]$, dynamically partitions the spectrum into low- and high-frequency regions.
+This constraint matches the normalized spectral dimensions and prevents $R$ from diverging to unreasonable values during training, which would otherwise lead to unstable partitioning and degraded performance.
+To maintain differentiability, we employ a soft partition:
 $\mathrm{Mask}_{\mathrm{high-freq}}=\sigma\left(\left(\mathrm{Dist}-R\right)\cdot\tau\right),$ (17)
-with temperature parameter $\tau=10$. The high-frequency components are then enhanced using a position-wise multilayer perceptron (MLP):
+where temperature $\tau=10$. High-frequency components are enhanced via a position-wise MLP:
 $A_{\mathrm{high-freq}}=\mathrm{MLP}\left(A\odot\mathrm{Mask}_{\mathrm{high-freq}}\right),$ (18)
-and concatenated with the original amplitude spectra. A subsequent $1\times 1$ convolution and residual connection yield the fused amplitude:
+then concatenated with the original amplitude $A$ and fused by a $1\times 1$ convolution and residual connection:
 $A^{\prime}=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(A,A_{\mathrm{high-freq}}\right)\right),\quad A\leftarrow A+A^{\prime}.$ (19)
 This module preserves low-frequency structures while adaptively reinforcing high-frequency details that capture complementary semantic information.
-The learnable radius $R$ self-adjusts to the input characteristics, eliminating the need for manual threshold tuning.
+The learnable radius $R$ self-adjusts to input characteristics, eliminating manual threshold tuning.
 III-F Loss Function
@@
 III-F2 Auxiliary Semantic Supervision
-An auxiliary segmentation head is added at the second-to-last stage ($t=2$) using an identical cross-entropy loss:
+An auxiliary segmentation head is added at the second-to-last stage ($t=2$), also using cross-entropy loss:
 $\mathcal{L}_{\mathrm{aux}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}^{(a)}_{i,c}.$ (21)
 III-F3 Phase-Amplitude Decoupling Constraint
-To resolve the coupling between amplitude and phase spectra, we introduce an Amplitude Consistency Loss that regularizes the fused amplitude $\mathcal{A}^{\prime}_{t}$ to remain faithful to its original decoupled state $\mathcal{A}_{t}$ across feature scales:
+To resolve the coupling between amplitude and phase spectra and enforce consistency between fused and original amplitudes, we introduce an amplitude consistency loss that regularizes $\mathcal{A}^{\prime}_{t}$ to remain faithful to its original decoupled state $\mathcal{A}_{t}$ across feature scales:
 $\mathcal{L}_{\mathrm{amp}}=\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}\alpha_{t}\left\|\mathcal{A}^{\prime}_{t}-\mathcal{A}_{t}\right\|_{F}^{2},$ (22)
-where $\mathcal{T}=\{0,1,2,3\}$ indexes the feature scales, $\alpha_{t}=0.5+(1.0-0.5)\times\frac{t}{3}$ imposes depth-progressive weighting, and $\|\cdot\|_{F}$ is the Frobenius norm.
+where $\mathcal{T}=\{0,1,2,3\}$ indexes the feature scales, $\alpha_{t}=0.5+0.5\frac{t}{3}$ imposes depth-progressive weighting, and $\|\cdot\|_{F}$ is the Frobenius norm.
 III-F4 Composite Objective
-The overall loss is formulated as:
+The overall loss is
 $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{seg}}+\lambda_{1}\mathcal{L}_{\mathrm{aux}}+\lambda_{2}\mathcal{L}_{\mathrm{amp}},$ (23)
@@
 IV-A Dataset Description
-WHU-OPT-SAR [23] is a multi-modal dataset developed for land use classification in Hubei Province, China.
-The dataset comprises 100 pairs of RGB, near-infrared (NIR), and SAR images covering identical geographic regions, with corresponding land type labels.
-The RGB images are acquired from GF1 satellite data, while the SAR images are sourced from GF3 satellite data.
-Both image types have been resampled to a uniform spatial resolution of 5 meters, with dimensions of 5556 $\times$ 3704 pixels.
-The dataset spans seven primary land use categories: farmland, urban areas, villages, water bodies, forests, roads, and others.
-In our study, images are cropped into non-overlapping patches of 512 $\times$ 512 pixels, resulting in 7000 total patches; 80% (5600 patches) are used for training and 20% (1400 patches) for testing.
-DDHR-SK is a subset of the DDHR multi-modal dataset developed for land cover classification (LCC) in Pohang City, South Korea [24].
-RGB images are collected from GF2 satellite data and SAR images from GF3 satellite data.
-To simulate adverse conditions such as clouds and fog, the RGB images were processed with a fog filter using the GNU Image Manipulation Program (GIMP).
-The dataset comprises 6173 paired cloud-affected RGB and SAR images, resampled to a spatial resolution of 1 meter and resized to 256 $\times$ 256 pixels.
-It covers five land cover types: buildings, farmland, greenery, water, and roads, with a public split of 3087 training images and 3086 validation images.
+WHU-OPT-SAR [28] is a multi-modal dataset for land use classification in Hubei Province, China.
+It contains 100 co-registered RGB, NIR, and SAR image triples with corresponding land type labels.
+RGB images are acquired by GF-1, and SAR images by GF-3; both are resampled to 5 m resolution and sized 5556 $\times$ 3704 pixels.
+Seven primary land use categories are covered: farmland, urban areas, villages, water bodies, forests, roads, and others.
+We crop each image into non-overlapping 512 $\times$ 512 patches, yielding 7,000 patches, of which 80% (5,600) are used for training and 20% (1,400) for testing.
+DDHR-SK is a subset of the DDHR multi-modal dataset for land cover classification in Pohang City, South Korea [29].
+RGB images from GF-2 and SAR images from GF-3 are paired, and RGB images are processed with a fog filter in GIMP to simulate adverse conditions.
+The dataset comprises 6,173 cloud-affected RGB-SAR pairs, resampled to 1 m resolution and resized to 256 $\times$ 256 pixels.
+It includes five land cover types: buildings, farmland, greenery, water, and roads, with 3,087 training and 3,086 validation images in the public split.
 IV-B Training Setting
-All models were trained and evaluated on a single NVIDIA H800 GPU, while the comparative experiments were carried out on a single NVIDIA A100 GPU.
-Training was conducted using the MMSegmentation framework [51] with a batch size of 4 and 80k iterations, including validation every 8k iterations.
-We employed the AdamW [52] optimizer with an initial learning rate of $1\times 10^{-4}$ and a Cosine Annealing schedule (with T_max set to the maximum iteration count).
-A fixed random seed of 42 was used to ensure reproducibility.
-Evaluation metrics include mean intersection over union (mIoU), Kappa coefficient, overall accuracy (OA), and F1 score, with performance reported at the iteration yielding the best mIoU.
-To assess robustness across diverse scenarios and mitigate class imbalance bias, the Kappa coefficient is computed for each sample, and the mean value is reported:
+All models were trained using the MMSegmentation framework [50] on a single NVIDIA H800 GPU (comparisons used an NVIDIA A100).
+We used a batch size of 4 for 80 k iterations, validating every 8 k iterations.
+Optimization was performed with AdamW [51], an initial learning rate of $1\times 10^{-4}$, and a cosine annealing schedule (T_max = total iterations).
+A fixed seed (42) ensured reproducibility.
+We report mean IoU (mIoU), Kappa coefficient, overall accuracy (OA), and F1 score, selecting the iteration with the best mIoU.
+To mitigate class-imbalance bias, we compute the Kappa coefficient per sample and report its mean:
 $\kappa_{i}=\frac{P_{o}^{(i)}-P_{e}^{(i)}}{1-P_{e}^{(i)}},\quad\mathrm{mKappa}=\frac{1}{M}\sum_{i=1}^{M}\kappa_{i},$ (24)
-where $P_{o}^{(i)}$ and $P_{e}^{(i)}$ denote the observed and expected agreements for the $i$-th sample, respectively.
+where $P_{o}^{(i)}$ and $P_{e}^{(i)}$ are the observed and expected agreements for sample $i$.
 IV-C Experimental Results on Datasets
-To assess the effectiveness of the proposed PAD method, we compared its performance with 17 different models, including 9 state-of-the-art multi-modal segmentation approaches and 8 cutting-edge unimodal segmentation methods.
-All experimental results were obtained through our replication of the original experimental configurations.
-Experiments were performed on two datasets: WHU-OPT-SAR, the largest multi-modal LCC dataset, and DDHR-SK, which contains RGB images captured under cloudy conditions.
-TABLE II: Experiment Results (%) on WHU-OPT-SAR.
+To assess the effectiveness of PAD, we replicated the original configurations of 17 comparison models, nine state-of-the-art multi-modal segmentation approaches and eight leading unimodal methods, on both WHU-OPT-SAR and DDHR-SK.
+TABLE II: Quantitative Results (%) on WHU-OPT-SAR.
+We report per-category IoU for seven land cover classes (Farmland, City, Village, Water, Forest, Road, Others), along with OA, mKappa, mF1, and mIoU.
+Bold and underlined values indicate the best and second-best performance, respectively.
 Method Venue IoU per Category OA mKappa mF1 mIoU
@@
 IV-C1 Evaluations on WHU-OPT-SAR
-As summarized in Table II, our PAD framework achieves significant performance improvements on WHU-OPT-SAR.
-Compared with the best unimodal method (DANet-RGB [13], achieving 53.25% mIoU and 83.19% OA) and the multi-modal baseline MRFS [25] (55.53% mIoU, 84.21% OA), PAD establishes a new state-of-the-art with 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1.
-In fine-grained categories, PAD attains 43.24% IoU on Road and 49.70% on Village segmentation, outperforming MRFS by 1.98% and 0.27%, respectively, thereby underlining the precision of our cross-modal fusion strategy.
+As shown in Table II, PAD achieves new state-of-the-art results: 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1.
+This surpasses the best unimodal method (DANet-RGB [18]: 53.25% mIoU, 83.19% OA) and the multi-modal baseline MRFS [30] (55.53% mIoU, 84.21% OA).
+In fine-grained classes, PAD attains 43.24% IoU on roads and 49.70% on villages, improving over MRFS by 1.98% and 0.27%, respectively.
 Figure 6: Visualization Results on WHU-OPT-SAR.
-Four representative cases are shown: (a) building-dense area, (b) specular water bodies, (c) vegetated region, and (d) urban–rural transition zone.
-For each case, the following images are presented from left to right: RGB image, SAR image, pseudo-color ground truth, and prediction maps from six competing methods.
-PAD is compared against leading single-modal baselines (DANet-RGB, TransUNet-SAR) as well as multi-modal approaches (CDDFuse, CMX, MRFS).
-Regions of interest are highlighted with gray boxes (for building/water areas) and circles (for vegetation zones); further details are provided in Section IV-C.
-Visual analyses support these quantitative results.
-For example, in Fig.
-6(a), PAD successfully recovers occluded road segments (highlighted in cyan) in building-dense regions through SAR phase analysis, while RGB-only models ignore these areas.
-In Fig.
-6(b), PAD effectively distinguishes specular reflection artifacts from actual water boundaries by leveraging SAR’s dielectric properties, overcoming limitations inherent in optical sensors under cloud interference.
-Additionally, Fig.
-6(c) demonstrates PAD’s ability to resolve spectral ambiguities in vegetation classification, and Fig.
-6(d) shows that PAD maintains cross-category consistency in complex urban–rural transitions.
-TABLE III: Experiment Results (%) on DDHR-SK.
+Four representative scenarios: (a) building-dense area, (b) specular water bodies, (c) vegetated region, and (d) urban-rural transition.
+For each scenario, images are shown (left to right): RGB, SAR, pseudo-color ground truth, and prediction maps from six competing methods.
+We compare PAD against single-modal baselines (DANet-RGB, TransUNet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS).
+Regions of interest are highlighted with gray boxes (buildings/water) and circles (vegetation); see Section IV-C for details.
+Visual results in Fig.
+6 further illustrate PAD’s benefits.
+In (a), PAD recovers occluded road segments (cyan) in building-dense areas via SAR phase cues, whereas RGB-only models miss them.
+In (b), PAD distinguishes specular reflections from true water boundaries under cloud interference.
+In (c)-(d), PAD resolves vegetation spectral ambiguities and maintains consistency across urban-rural transitions.
+TABLE III: Quantitative Results (%) on DDHR-SK. We report per-category IoU for seven land cover classes (Buildings, Roads, Greenery, Water, Farmland), along with OA, mKappa, mF1, and mIoU.
 Method Venue IoU per Category OA mKappa mF1 mIoU
@@
 PAD (Ours) 97.09 86.12 95.38 99.47 94.89 98.04 84.38 97.16 94.59
-TABLE IV: Ablation Experimental Results (%) on Main Modules on WHU-OPT-SAR.
-ID Components IoU per Category aAcc mKappa mF1 mIoU
+IV-C2 Evaluations on DDHR-SK
+As shown in Table III, PAD exhibits superior performance under adverse (cloudy) conditions on DDHR-SK.
+It achieves 94.59% mIoU and 97.65% OA, outperforming leading unimodal methods (e.g., Mask2former-RGB: 90.25% mIoU, 94.76% mF1; DANet-SAR: 83.31% mIoU, 90.28% mF1) and the multi-modal baseline MRFS (93.57% mIoU, 82.24% mKappa).
+PAD also attains 82.24% mKappa and 96.60% mF1, with boundary-sensitive categories achieving 99.47% IoU on water and 86.12% IoU on roads, surpassing MRFS by 2.44% and 0.08%, respectively.
+Qualitative results in Fig.
+7 further demonstrate PAD’s ability to delineate water infiltration in irrigated fields and maintain continuous road networks beneath fog.
+Figure 7: Visualization Results on DDHR-SK.
+Four boundary-focused scenarios: (a) coastal demarcation, (b) water-cropland interface, (c) vegetation-building complex, and (d) road network topology.
+For each scenario, panels show the RGB input, SAR coherence map, pseudo-color ground truth, and prediction maps from six methods.
+We compare PAD against unimodal baselines (Mask2former-RGB, DANet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS).
+Critical boundary regions are highlighted with orange boxes; see Section IV-C for details.
+The visual results further demonstrate PAD’s capability for cross-modal boundary refinement.
+In Fig.
+7(a), PAD accurately reconstructs periodically submerged coastlines via SAR coherence analysis, whereas RGB-only models confuse tidal sediments with terrestrial areas.
+In Fig.
+7(b), PAD leverages SAR features to delineate water-infiltration boundaries in irrigated fields, overcoming the over-segmentation common in optical sensors.
+Moreover, in Fig.
+7(c) and Fig.
+7(d), PAD achieves continuous road networks under vegetation occlusion and precise material boundaries in vegetation-building complexes, respectively.
+TABLE IV: Computational Efficiency Comparison.
+Model size (#Params), inference time per image, and throughput (FPS) for competitive methods.
+Lower #Params and inference time, and higher FPS, indicate better efficiency.
+Method #Params (M) Inference Time (ms) FPS (frames/s)
+CMX [31] 15.00 206.70 1,957.8
+MRFS [30] 10.30 695.85 1,342.2
+ASANet [32] 7.82 72.91 2,245.2
+PAD (Ours) 3.88 5.16 3,228.7
+IV-C3 Computational Efficiency Analysis
+We evaluate the computational efficiency of PAD in terms of model size (Fusion Modules Only), inference time per image, and frames per second (FPS), as summarized in Table IV.
+We compare against four high-mIoU methods, CMX [31], MRFS [30], ASANet [32], and CDDFuse [12] (excluded from detailed analysis due to its two-stage design and lower efficiency).
+All metrics are aggregated across fusion stages.
+As shown in Table IV, PAD uses only 3.882 M parameters, achieves the fastest inference time (5.16 ms per image), and delivers the highest FPS (3,228.7).
+These results demonstrate that PAD not only matches or exceeds accuracy of existing methods but also offers real-time efficiency suitable for operational deployment.
+IV-C4 Ablation Study
+We performed comprehensive ablation studies on WHU-OPT-SAR to validate the contributions of our core components: the Amplitude Spectrum Fusion (ASF) module, the Phase Spectrum Correction (PSC) module, and Multi-Stage Fusion (MSF).
+[…]
+ID Components IoU per Category OA mKappa mF1 mIoU
 MSF ASF PSC Farmland City Village Water Forest Road Others
@@
 […]
-IV-C2 Evaluations on DDHR-SK
-As shown in Table III, PAD exhibits superior boundary segmentation performance on the relatively simplified scenes of DDHR-SK.
-Compared with leading unimodal models (e.g., Mask2former-RGB [18], achieving 90.25% mIoU and 94.76% mF1, DANet-SAR [13], yielding 83.31% mIoU and 90.28% mF1) and the multi-modal baseline MRFS [25] (93.57% mIoU, 82.24% mKappa), PAD attains new state-of-the-art performance with 94.59% mIoU, 97.65% OA, 82.24% mKappa, and 96.60% mF1.
-For boundary-sensitive categories, PAD achieves 99.47% IoU on Water and 86.12% IoU on Road edges, surpassing MRFS by 2.44% and 0.08%, respectively.
-Figure 7: Visualization Results on DDHR-SK.
-Four boundary-focused cases are shown: (a) coastal demarcation, (b) water–cropland interface, (c) vegetation–building complex, and (d) road network topology.
-PAD is compared with leading unimodal methods (Mask2former-RGB, DANet-SAR) and state-of-the-art multi-modal approaches (CDDFuse, CMX, MRFS).
-For each case, the panels display the RGB input, SAR coherence map, pseudo-color ground truth, and prediction maps from six methods.
-Critical boundary regions are highlighted with orange boxes; further analysis is provided in Section IV-C.
-The visual results further reveal PAD’s capability in cross-modal boundary refinement.
-In Fig.
-7(a), PAD accurately reconstructs periodically submerged coastal lines via SAR coherence analysis, whereas RGB models tend to confuse tidal sediments with terrestrial areas.
-In Fig.
-7(b), PAD leverages SAR features to delineate water infiltration boundaries in irrigated fields, overcoming the over-segmentation issues common to optical sensors.
-Moreover, in Fig.
-7(c) and Fig.
-7(d), PAD achieves continuous road networks under vegetation occlusion and precise material boundaries in vegetation–building complexes, respectively.
-IV-C3 Ablation Study
-We performed comprehensive ablation studies on WHU-OPT-SAR to validate the contributions of the core components of our approach: Amplitude Spectrum Fusion module (ASF), Phase Spectrum Correction module (PSC), and Multi-Stage Fusion (MSF).
 Main Modules
-Eight model variants (see Table IV) were tested.
-The baseline variant (I), which replaces MSF with early fusion and substitutes the PAD (ASF+PSC) with pixel-wise additive fusion, achieves the lowest mIoU (54.98%).
-Enabling MSF alone (II) improves mIoU by 0.39%, while using ASF alone (III) attains 55.86% mIoU, and solely PSC (IV) adds a smaller gain.
-Combining MSF with either ASF (V) or PSC (VI) yields similar improvements (mIoU around 56.11–56.12%).
-Integrating all components achieves the best performance (56.26% mIoU), demonstrating the effective synergy of all modules.
-TABLE V: Ablation Experiment Results (%) on Loss Function Weights on WHU-OPT-SAR.
-Configuration IoU per Category aAcc Kappa mF1 mIoU
+Eight variants (see Table V) were evaluated.
+The baseline (I), which replaces MSF with early fusion and substitutes PAD (ASF+PSC) with pixel-wise addition, achieves the lowest mIoU (54.98%).
+Enabling MSF alone (II) yields a +0.39% gain, ASF alone (III) reaches 55.86% mIoU, and PSC alone (IV) provides a smaller improvement.
+Combining MSF with ASF (V) or PSC (VI) yields similar gains (56.11-56.12% mIoU).
+Integrating all three modules achieves the best performance (56.26% mIoU), demonstrating their effective synergy.
+TABLE VI: Ablation Study on Fusion Modules (%) for WHU-OPT-SAR. We compare alternative amplitude-only, phase-only, and full fusion designs by integrating each into our framework.
+Fusion Module IoU per Category OA mKappa mF1 mIoU
+Farmland City Village Water Forest Road Others
+Amplitude FelNFN [16] 66.93 55.80 45.15 61.27 82.77 35.54 19.17 82.95 47.31 66.46 52.38
+FCENet [14] 68.27 55.18 46.79 61.17 83.43 37.09 16.71 83.43 49.82 66.52 52.66
+FGSwin [53] 69.87 57.43 49.03 65.04 84.12 43.16 22.26 84.51 53.53 69.70 55.85
+Ours 69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
+Phase FelNFN [16] 60.66 53.25 31.99 54.14 80.64 19.84 15.02 79.42 33.32 58.89 45.08
+Channel Attention 69.58 58.66 49.78 64.28 83.96 43.43 22.65 84.39 53.54 69.93 56.05
+Ours 69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
+Whole FelNFN [16] 66.43 51.97 44.36 58.48 82.80 32.33 12.64 82.33 47.35 63.63 49.86
+FCENet++ [14] 66.87 53.01 46.41 60.30 82.64 30.21 14.51 82.78 43.83 64.33 50.56
+FGSwin [53] 69.88 57.86 49.48 65.26 84.15 42.49 22.36 84.56 53.41 69.76 55.92
+Frefusion [15] 69.91 57.99 49.35 64.22 84.00 42.72 22.46 84.41 53.98 69.69 55.81
+SFINet++ [46] 68.23 55.65 47.31 62.35 83.10 38.48 19.04 83.50 49.56 67.43 53.45
+Ours 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
+Fusion Modules
+To further validate ASF and PSC, we replaced each with five existing frequency-domain designs, FelNFN [16], FCENet [14], FGSwin [53], Frefusion [15], and SFINet++ [46], while keeping the rest of our framework unchanged (see Table VI).
+These methods represent common strategies such as fixed-kernel spectral convolution (FelNFN), frequency-spatial interaction via Hadamard products/matrix multiplication (FCENet), learned spectral filtering (FGSwin), attention-based cross-domain fusion (Frefusion), and amplitude-phase parallel processing (SFINet++).
+Our ASF outperforms other amplitude-specific methods (56.11% vs.
+55.85% mIoU for FGSwin), and PSC consistently exceeds phase baselines (56.12% vs.
+56.05% for channel attention and 45.08% for FelNFN).
+At the full-fusion level, our integrated design still leads (56.26% mIoU), indicating both modules are effective in isolation and synergistic when combined.
+These results confirm that adaptive frequency modulation (ASF) and spatially guided phase correction (PSC) offer structurally unique and more effective alternatives to existing paradigms.
+TABLE VII: Ablation Study on Loss Function Weights (%) for WHU-OPT-SAR. We vary $\lambda_{1}$ with fixed $\lambda_{2}=0.1$ and $\lambda_{2}$ with fixed $\lambda_{1}=0.4$.
+Configuration IoU per Category OA mKappa mF1 mIoU
 Farmland City Village Water Forest Road Others
@@
 Loss Function Weights
-We conducted weight ablation studies (see Table V) to examine the impact of the trade-off between the amplitude consistency loss ($\mathcal{L}_{\mathrm{amp}}$) and the segmentation loss ($\mathcal{L}_{\mathrm{seg}}$).
-The best performance (mIoU = 56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$.
-Increasing $\lambda_{1}$ above 0.4 degrades performance, notably in under-represented classes, while an overly low $\lambda_{2}$ significantly diminishes performance, confirming that segmentation supervision must remain the dominant objective.
-TABLE VI: Ablation Results (%) on the Initial Frequency Radius on WHU-OPT-SAR.
-Configuration IoU per Category aAcc Kappa mF1 mIoU
+We conducted weight ablations (see Table VII) to assess the trade-off between the amplitude consistency loss $\mathcal{L}_{\mathrm{amp}}$ and segmentation loss $\mathcal{L}_{\mathrm{seg}}$.
+The best mIoU (56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$.
+Increasing $\lambda_{1}$ above 0.4 degrades performance, especially on under-represented classes, while too small $\lambda_{2}$ significantly reduces accuracy, confirming segmentation supervision must remain dominant.
+TABLE VIII: Ablation Study on Initial Frequency Radius (%) for WHU-OPT-SAR. We vary the trainable radius parameter in the ASF module.
+Radius IoU per Category OA mKappa mF1 mIoU
 Farmland City Village Water Forest Road Others
-$radius=0.01$ 70.08 57.44 49.31 66.29 84.18 43.08 22.25 84.66 53.92 69.88 56.09
-$radius=0.1$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
-$radius=1$ 69.79 58.19 49.71 64.81 84.15 43.44 23.03 84.52 53.88 70.04 56.16
-$radius=10$ 69.71 57.71 49.69 65.63 84.18 43.21 23.56 84.57 53.93 70.12 56.24
-$radius=100$ 69.90 57.23 49.57 65.23 84.13 43.34 22.94 84.53 53.88 69.93 56.05
+$0.01$ 70.08 57.44 49.31 66.29 84.18 43.08 22.25 84.66 53.92 69.88 56.09
+$0.1$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
+$1$ 69.79 58.19 49.71 64.81 84.15 43.44 23.03 84.52 53.88 70.04 56.16
+$10$ 69.71 57.71 49.69 65.63 84.18 43.21 23.56 84.57 53.93 70.12 56.24
+$100$ 69.90 57.23 49.57 65.23 84.13 43.34 22.94 84.53 53.88 69.93 56.05
 Initial Frequency Radius
-The initial frequency radius parameter $r$, which controls the proportion of preserved low-frequency components in our phase-aware filtering, was varied (see Table VI).
-Our experiments show that $r=0.1$ achieves an optimal balance (56.26% mIoU).
-Values that overly suppress low-frequency information or over-smooth the features lead to inferior performance, though the method remains robust for $r\in[0.1,10]$.
+We varied the initial frequency radius $r$, which controls the preserved low-frequency proportion in ASF (see Table VIII).
+The optimal balance is at $r=0.1$ (56.26% mIoU).
+Values that overly suppress or over-smooth low-frequency content degrade performance, although the method remains robust for $r\in[0.1,10]$.
 IV-D Visualization at Distribution Level
-To further analyze the feature distribution, we extracted multi-scale features from the RGB and SAR encoders, as well as from the fusion modules.
-After upsampling to the original resolution, the features were concatenated channel-wise.
-To balance computational efficiency with preserving discriminative information, a two-stage dimensionality reduction was applied: 1) random spatial sampling, followed by 2) principal component analysis (PCA), retaining 50 principal components (accounting for over 89% of the variance).
-The final visualization was generated using t-SNE [54] (with a perplexity of 30 and 1,000 iterations).
+To further analyze feature distributions, we extracted multi-scale features from the RGB and SAR encoders and from the fusion modules.
+After upsampling to the original resolution, we concatenated features channel-wise.
+To balance computational efficiency with discriminative power, we applied a two-stage dimensionality reduction: a) random spatial sampling, and b) principal component analysis (PCA), retaining 50 components (over 89% variance).
+The final visualization was generated using t-SNE [54] (perplexity = 30, 1,000 iterations).
 Refer to caption Refer to caption Refer to caption
@@
 Figure 8: t-SNE Visualization of Cross-modal Feature Distribution Evolution.
-(a) RGB encoder features exhibit spectral continuity but have blurred boundaries; (b) SAR encoder features, while strongly sensitive to geometry, are affected by speckle noise; (c) PAD-fused features display enhanced intra-class compactness and improved semantic separation.
-As illustrated in Fig.
-8, the PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability.
-Quantitatively, the mean intra-class standard variance is reduced by 20.39% for Road, 13.16% for Water, 6.46% for Forest, and 4.43% for Farmland, compared to the best unimodal baselines.
-In addition, significant increases in Euclidean distances between critical class pairs, such as a 75.75% increase in the City/Village pair (66.56 to 116.95) and a 13.29% increase for the Farmland/Forest pair (107.64 to 121.89), indicate that the fused features capture cross-modal discriminative patterns effectively.
-Overall, these experimental results confirm that the PAD framework significantly improves classification accuracy and prediction consistency by effectively integrating multi-modal information and suppressing noise, while maintaining the discriminative spectral and geometric characteristics inherent in the data.
+(a) RGB encoder features show spectral continuity but blurred class boundaries.
+(b) SAR encoder features capture geometric details but are affected by speckle noise.
+(c) PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separation.
+As shown in Fig.
+8, PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability.
+Quantitatively, mean intra-class standard variance is reduced by 20.39% for roads, 13.16% for water, 6.46% for forest, and 4.43% for farmland, compared to the best unimodal baselines.
+Moreover, Euclidean distances between critical class pairs increase significantly: City/Village by 75.75% (66.56 $to$ 116.95) and Farmland/Forest by 13.29% (107.64 $to$ 121.89), indicating that PAD captures cross-modal discriminative patterns effectively.
+TABLE IX: Statistical Comparison Between Models w/ and w/o ASF.
+Metric w/ ASF w/o ASF
+Mean Pearson Correlation 0.0364 0.0369
+Mean Spearman Correlation 0.0482 0.0488
+Paired t-test p-value 0.0016
+Wilcoxon test p-value $6.4\times 10^{-7}$
+Cohen’s $d$ (effect size) -0.038
+IV-E Noise Sensitivity of ASF
+To assess whether ASF amplifies high-frequency speckle noise in SAR images, we conducted a correlation analysis using the Equivalent Number of Looks (ENL), a standard measure of speckle.
+For each image, we computed pixel-wise ENL maps and then measured the Pearson and Spearman correlations between ENL and classification accuracy.
+As shown in Table IX, the model with ASF exhibits nearly identical correlation coefficients to the baseline, differences in mean correlation are below 0.001, and Cohen’s $d=-0.038$ indicates a negligible effect size.
+These results confirm that ASF does not increase noise sensitivity and remains robust in low-SNR regions.
+Overall, these results confirm that PAD significantly improves classification accuracy and prediction consistency by integrating multi-modal information, suppressing noise, and preserving discriminative spectral and geometric characteristics inherent in the data.
 V Conclusion
 In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land cover classification.
-Our frequency-domain analysis revealed that phase components are consistent across modalities, while amplitude spectra capture modality-specific details, providing the first quantitative physical prior for multi-source remote sensing fusion.
-PAD separates and aligns phase features while adaptively fusing amplitude components.
-The Phase Spectrum Correction (PSC) module resolves structural discrepancies, and the Amplitude Spectrum Fusion (ASF) module dynamically integrates SAR’s high-frequency details with RGB’s low-frequency context.
-Extensive experiments demonstrate that PAD outperforms existing methods in both accuracy and robustness.
-Future work will focus on improving phase alignment, particularly under significant geometric distortions due to imprecise cross-modal registration.
+Our frequency-domain analysis quantitatively demonstrates that phase components remain consistent across modalities, while amplitude spectra capture modality-specific details, providing the first physical prior for multi-source remote sensing fusion.
+PAD disentangles and aligns phase features via the Phase Spectrum Correction (PSC) module, and adaptively fuses amplitude components using the Amplitude Spectrum Fusion (ASF) module, which integrates SAR’s high-frequency details with RGB’s low-frequency context.
+Extensive experiments on two challenging datasets show that PAD outperforms existing methods in both accuracy and robustness.
+Despite its demonstrated effectiveness, PAD assumes well-aligned input modalities and relies on hard-label supervision.
+In future work, we will integrate self-supervised cross-modal registration refinement and soft-label modeling, and investigate transformer-based frequency fusion to further enhance global context awareness.

### v3

@@
 Abstract
-The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and underutilized spectral complementarity.
-Existing methods often fail to decouple shared structural features from modality-complementary radiometric attributes, causing feature conflicts and information loss.
-To address this, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-complementary) components in the Fourier domain, thus reinforcing shared structures while preserving complementary characteristics to improve fusion quality.
-Unlike prior approaches that overlook the distinct physical properties encoded in frequency spectra, PAD is the first to introduce explicit amplitude-phase decoupling for multi-modal fusion.
-Specifically, PAD comprises two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features via convolution-guided scaling to enhance geometric consistency; and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive multilayer perceptrons, leveraging SAR’s morphological sensitivity and RGB’s spectral richness.
-Extensive experiments on WHU-OPT-SAR and DDHR-SK datasets demonstrate state-of-the-art performance.
-Our work establishes a new paradigm for physics-aware multi-modal fusion in remote sensing.
+The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and underexploited spectral complementarity.
+Existing approaches often fail to decouple shared structural features from modality-complementary radiometric attributes, resulting in feature conflicts and information loss.
+To address this, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-complementary) components in the Fourier domain.
+This design reinforces shared structures while preserving complementary characteristics, thereby enhancing fusion quality.
+Unlike previous methods that overlook the distinct physical properties encoded in frequency spectra, PAD explicitly introduces amplitude-phase decoupling for multi-modal fusion.
+Specifically, PAD comprises two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features via convolution-guided scaling to improve geometric consistency; and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high- and low-frequency patterns using frequency-adaptive multilayer perceptrons, effectively exploiting SAR’s morphological sensitivity and RGB’s spectral richness.
+Extensive experiments on WHU-OPT-SAR and DDHR-SK demonstrate state-of-the-art performance.
+This work establishes a new paradigm for physics-aware multi-modal fusion in remote sensing.
 The code will be available at https://github.com/RanFeng2/PAD.
-Index Terms:
-Land Cover Classification, RGB-SAR multi-modality, multi-modal segmentation, Synthetic Aperture Radar, Remote Sensing
 I Introduction
@@
 reflectance).
-(b) Fusion feature space comprises heterogeneous, shared, and complementary components; effective fusion suppresses heterogeneous features while enhancing shared and complementary patterns.
-(c)-(g) Frequency-domain analysis on WHU-OPT-SAR (see Section III-A): (c) spatial discrepancies exhibit anisotropic distributions; (d) relative amplitude differences concentrate in high-frequency regions; (e) phase discrepancies are minimal and cluster in low-frequency bands; (f)-(g) sampling-rate variations significantly affect spatial discrepancy distributions.
-Despite this complementarity, RGB-SAR fusion faces a key challenge: inherent imaging heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB) leads to significant spectral misalignment and hinders effective cross-modal feature integration (see Fig.
+(b) The fusion feature space contains heterogeneous, shared, and complementary components; effective fusion suppresses heterogeneous cues while enhancing shared and complementary patterns.
+(c-g) Frequency-domain analysis on WHU-OPT-SAR (see §III-A): (c) spatial discrepancies show anisotropic distributions; (d) relative amplitude differences concentrate in high-frequency regions; (e) phase discrepancies are minimal and primarily cluster in low-frequency bands; (f-g) varying the sampling rate significantly affects spatial discrepancy distributions.
+Despite this complementarity, RGB-SAR fusion faces a key challenge: the inherent heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB) causes spectral misalignment and hinders effective cross-modal feature integration (see Fig.
 1(a)).
 In SAR images, objects often appear as discrete backscatter responses, lacking fundamental visual cues, such as geometric contours, textures, and consistent shapes, that are readily available in RGB imagery [4].
-Compared to natural-image fusion, integrating SAR and RGB data presents pronounced radiometric and geometric disparities, limiting existing frameworks [5, 6, 7].
-Deep learning, particularly convolutional neural networks (CNNs), has achieved notable success in LCC by extracting powerful features [8, 9].
-To mitigate modality heterogeneity, most approaches use dual-branch architectures to extract multi-scale features independently, then fuse them via element-wise summation, mean aggregation, or channel-wise concatenation before a shared decoder or segmentation head [10].
-However, accurate modeling of complementary features is still impeded by SAR-RGB heterogeneity.
-Inspired by contrastive learning [11, 12, 13], recent works align modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts (see Fig.
+Compared with natural image fusion, integrating SAR and RGB data presents pronounced radiometric and geometric disparities, which limits existing frameworks [5, 6, 7].
+Deep learning, particularly convolutional neural networks (CNNs), has achieved remarkable success in LCC by extracting powerful features [8, 9, 10, 11].
+To mitigate modality heterogeneity, most approaches employ dual-branch architectures to extract multi-scale features independently and then fuse them via element-wise summation, mean aggregation, or channel-wise concatenation before a shared decoder or segmentation head [12].
+However, accurately modeling complementary features remains impeded by SAR-RGB heterogeneity.
+Inspired by contrastive learning [13, 14], recent works align modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts (see Fig.
 1(b)), yet spatial-domain modeling remains inefficient.
@@
 [18] treat phase as modality-shared features but discard amplitude information essential for terrain delineation and do not systematically validate phase sharing.
-We conduct comprehensive spectral analysis (see Fig.
+We conduct a comprehensive spectral analysis (see Fig.
 1(c)-(g)) to reveal three critical frequency-domain priors: 1) SAR-RGB spatial-domain differences are anisotropic and susceptible to pixel-level mismatch (see Fig.
-1(c)); 2) average phase spectra differ minimally and mainly in low frequencies, indicating phase as shared information (see Fig.
-1(e)); and 3) amplitude mismatches exhibit distinct high/low-frequency patterns, most pronounced in the high-frequency band, highlighting the need for frequency-aware modeling (see Fig.
+1(c)); 2) average phase spectra differ minimally and primarily in low frequencies, indicating phase as shared information (see Fig.
+1(e)); and 3) amplitude mismatches exhibit distinct high- and low-frequency patterns, most pronounced in the high-frequency band, highlighting the need for frequency-aware modeling (see Fig.
 1(d)).
-Motivated by these insights, we propose Phase-Amplitude Decoupling (PAD), which disentangles phase components capturing modality-invariant structures from amplitude components encoding modality-specific details.
+Motivated by these insights, we propose Phase-Amplitude Decoupling (PAD), which decouples phase components capturing modality-shared structures from amplitude components encoding modality-complementary details.
 By explicitly aligning shared phase representations, PAD enhances complementary amplitude feature extraction and reduces cross-modal conflicts.
-PAD comprises two modules: Phase Spectrum Correction (PSC), which aligns and reinforces structural details via convolution-guided scaling; and Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive MLPs.
+PAD comprises two modules: Phase Spectrum Correction (PSC), which aligns and reinforces structural details via convolution-guided scaling; and Amplitude Spectrum Fusion (ASF), which dynamically integrates high- and low-frequency patterns using frequency-adaptive multilayer perceptrons (MLPs).
 Our contributions are fourfold:
-  • We systematically analyze registration datasets to establish frequency distributions of SAR-RGB common (phase) and unique (amplitude) patterns, offering interpretable physical priors for multi-source fusion.
-  • We introduce frequency-domain amplitude-phase decoupling for multi-modal fusion, orthogonally separating phase (modality-shared) and amplitude (modality-specific) features to resolve heterogeneity-induced conflicts.
+  • We systematically analyze registered datasets to establish the frequency distributions of SAR-RGB shared (phase) and complementary (amplitude) patterns, providing interpretable physical priors for multi-source fusion.
+  • We introduce frequency-domain amplitude-phase decoupling for multi-modal fusion, orthogonally separating phase (modality-shared) and amplitude (modality-complementary) features to resolve heterogeneity-induced conflicts.
   • PSC employs convolution-guided multiplicative scaling to align and reinforce critical structural details for robust fusion.
-  • ASF uses frequency-domain MLPs to integrate complementary high-frequency (local details) and low-frequency (global structure) components, leveraging SAR’s morphological sensitivity and RGB’s spectral richness for improved segmentation.
+  • ASF leverages frequency-adaptive MLPs to integrate complementary high-frequency (local details) and low-frequency (global structure) components, exploiting SAR’s morphological sensitivity and RGB’s spectral richness for improved segmentation.
 Figure 2: Flowchart of the PAD Framework.
-Registered RGB and SAR images, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are processed by asymmetric backbones to extract features $x_{\mathrm{RGB}}$ and $x_{\mathrm{SAR}}$ at each stage.
-These features are passed through $n$ PAD fusion modules, each comprising: (a) the SCF module, (c) the ASF module, and (d) the PSC module.
-(b) FD and (e) FR denote the spatial-to-frequency and frequency-to-spatial transformations, respectively.
-Each stage’s fused features are concatenated channel-wise (outer circle $C$) and fed into a shared decoder.
-Here, $\mathcal{A}$ and $\mathcal{P}$ represent amplitude and phase spectra.
+Registered RGB and SAR inputs, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are processed by asymmetric backbones to extract features $x_{\mathrm{RGB}}$ and $x_{\mathrm{SAR}}$ at each stage.
+These features pass through $n$ PAD fusion modules, each comprising: (a) SCF; (b) FD (spatial-to-frequency transform); (c) ASF; (d) PSC; and (e) FR (frequency-to-spatial transform).
+At each stage, the fused features are concatenated channel-wise (outer circle, $C$) and fed into a shared decoder.
+Here, $\mathcal{A}$ and $\mathcal{P}$ denote the amplitude and phase spectra, respectively.
 II Related Work
@@
 II-A1 Cross-Modal Transfer Learning
-Transfer learning [26] adapts knowledge from one modality to another, but the pronounced differences between SAR and RGB data often degrade learned representations, limiting multi-modal applicability.
+Transfer learning [27] adapts knowledge from one modality to another, but the pronounced differences between SAR and RGB often degrade learned representations, limiting their multi-modal applicability.
 II-A2 Hierarchical Feature Fusion
-Feature fusion has evolved through three stages: a) Basic Fusion Operators: Early methods used concatenation or weighted summation [27], but these are prone to modality dominance and redundancy.
-b) Attention-Guided Fusion: Spatial-channel attention mechanisms address these issues [28, 29].
-For example, MCANet’s cross-modal attention (MCAM) models inter-modality dependencies; MRFS [30] introduces progressive cycle attention for infrared-visible fusion; and CMX [31] employs bidirectional channel-spatial attention.
+Feature fusion has evolved through three stages: a) Basic Fusion Operators: Early methods employed concatenation or weighted summation [28], but these are prone to modality dominance and redundancy.
+b) Attention-Guided Fusion: Spatial-channel attention mechanisms alleviate these issues [29, 30].
+For example, MCANet’s cross-modal attention (MCAM) models inter-modality dependencies; MRFS [31] introduces progressive cycle attention for infrared-visible fusion; and CMX [32] applies bidirectional channel-spatial attention.
 However, these paradigms incur high computational costs.
-c) Complementarity-Aware Fusion: Recent approaches explicitly disentangle shared and modality-specific features.
+c) Complementarity-Aware Fusion: Recent approaches explicitly decouple shared and complementary features.
 ASANet [33] uses semantic focusing modules to emphasize complementary cues; CDDFuse [13] applies contrastive learning to separate private and shared representations; and CEN [34] leverages dynamic channel exchange with independent batch normalization.
@@
 Frequency representation learning has emerged as a powerful tool across various vision tasks due to its capacity for global modeling and structural decomposition.
-In this subsection, we review three relevant areas:
+We briefly review three relevant areas:
 II-B1 Spectral Decomposition for Structure-Style Separation
-Fourier-based decomposition isolates phase (structural semantics) and amplitude (appearance) for domain generalization and style transfer [34, 35].
-Early methods used feature statistics for style normalization [36, 37], while recent works employ Fourier transforms for clearer separation, enabling content-preserving augmentation by swapping amplitudes across domains [38, 39].
+Fourier-based decomposition isolates phase (structural semantics) and amplitude (appearance) for domain generalization and style transfer [35].
+Early methods relied on feature statistics for style normalization [36, 37], while recent works apply Fourier transforms for clearer separation, enabling content-preserving augmentation by swapping amplitudes across domains [38, 39].
 II-B2 Frequency-Driven Global Modeling
@@
 Fast Fourier Convolution (FFC) [40] and FFTformer [41] leverage spectral processing to replace or accelerate spatial attention.
-Spectral token mixing has also been effective in lightweight transformers [42, 43].
+Spectral token mixing has also proven effective in lightweight transformers [42, 43].
 II-B3 Frequency-Based Inductive Biases in Dense Prediction
-Spectral priors enhance dense prediction tasks.
+Spectral priors benefit dense prediction tasks [44, 45, 46, 47].
 DFF [48] uses frequency-domain dynamic filtering to amplify invariant low-frequency features, improving cross-domain generalization.
-FreqFusion [45] designs learnable high- or low-pass filters for semantic consistency and boundary refinement.
-FCENet [14] employs frequency-selective modules for cross-modal spectral correlations in RGB-NIR denoising.
-In summary, frequency-based representations offer a flexible, physically interpretable design space for vision tasks; yet, few frameworks target task-specific fusion in remote sensing.
-We address this gap in the next section.
+FreqFusion [49] designs learnable high- and low-pass filters for semantic consistency and boundary refinement.
+FCENet [15] employs frequency-selective modules to model cross-modal spectral correlations in RGB-NIR denoising.
+In summary, frequency-based representations offer a flexible and physically interpretable design space for vision tasks.
+Yet, few frameworks address task-specific fusion in remote sensing.
+We target this gap in the next section.
 II-C Frequency-Domain Fusion Methods
-Applying frequency representations to multi-modal fusion is recent.
+The application of frequency representations to multi-modal fusion is relatively recent.
 Shi et al.
-[15] introduce Frefusion, using spectral attention to merge modalities in the frequency domain.
-Frequency spectra naturally decompose signals into phase (modality-invariant structures) and amplitude (modality-specific details), offering a principled fusion framework.
-However, existing methods often transplant spatial operators, MLPs [16], convolutions [14], or attention [15], into the spectral domain, modulating magnitude while neglecting the distinct roles of phase and amplitude.
-Consequently, shared information may be preserved, but complementary cues are suppressed.
+[16] introduce Frefusion, which applies spectral attention to merge modalities in the frequency domain.
+Frequency spectra naturally decompose signals into phase (modality-shared structures) and amplitude (modality-complementary details), offering a principled basis for fusion.
+However, existing methods often transplant spatial operators, such as MLPs [17], convolutions [15], or attention [16], into the spectral domain, modulating magnitude while neglecting the distinct roles of phase and amplitude.
+As a result, shared information may be preserved, but complementary cues are frequently suppressed.
 In contrast, our PAD framework explicitly decouples phase and amplitude spectra.
-Through systematic spectral analysis (see Section III-A), we show that SAR and RGB share strong phase consistency while amplitude captures complementary patterns.
-PAD introduces (1) a Phase Spectrum Correction module for robust structural alignment and (2) an Amplitude Spectrum Fusion module with frequency-adaptive refinement.
+Through systematic spectral analysis (see §III-A), we show that SAR and RGB exhibit strong phase consistency, while amplitude captures complementary patterns.
+PAD introduces 1) a Phase Spectrum Correction (PSC) module, which performs robust structural alignment, and 2) an Amplitude Spectrum Fusion (ASF) module, which conducts frequency-adaptive refinement.
 This decomposition-correction-fusion pipeline distinguishes PAD from previous FFT-based approaches by providing a physically interpretable strategy for multi-modal fusion.
 The two most related works, FeINFN [17] and SFINet++ [50], also leverage amplitude and phase in the frequency domain for high-frequency detail enhancement and global context modeling.
-However, they lack a physically grounded interpretation of amplitude-phase roles and do not explicitly decouple shared versus modality-specific information.
+However, they lack a physically grounded interpretation of amplitude-phase roles and do not explicitly decouple shared versus complementary information.
 In contrast, PAD is the first framework to formulate and exploit amplitude-phase decomposition for principled alignment and complementary integration across modalities.
 III Proposed Method
-To address the fundamental conflict between modality-invariant geometry and modality-specific appearance in SAR-RGB fusion, we introduce Phase-Amplitude Decoupling (PAD), a frequency-aware multi-modal fusion framework.
-The overall architecture, shown in Fig.
-2, comprises three modules: the Spatial Complementary Fusion (SCF) module, which preserves fine-grained spatial priors; the Phase Spectrum Correction (PSC) module, which aligns and reinforces shared structural details; and the Amplitude Spectrum Fusion (ASF) module, which adaptively integrates complementary high- and low-frequency amplitude cues.
-Unlike previous methods that apply spectral operations uniformly, PAD explicitly decouples and recombines phase and amplitude with dedicated modules to enable physics-consistent fusion.
+To address the fundamental conflict between modality-shared geometry and modality-complementary appearance in SAR-RGB fusion, we introduce Phase-Amplitude Decoupling (PAD), a frequency-aware multi-modal fusion framework.
+The overall architecture in Fig.
+2 comprises three modules: the Spatial Complementary Fusion (SCF) module, which preserves fine-grained spatial priors; the Phase Spectrum Correction (PSC) module, which aligns and reinforces shared structural details; and the Amplitude Spectrum Fusion (ASF) module, which adaptively integrates complementary high- and low-frequency amplitude cues.
+Unlike previous methods that apply spectral operations uniformly, PAD explicitly decouples and then recombines phase and amplitude with dedicated modules for physics-consistent fusion.
 Input : SAR/RGB image pairs $\{I_{\mathrm{SAR}}^{(i)},I_{\mathrm{RGB}}^{(i)}\}_{i=1}^{N}$
@@
 1
+21ex
 1: Preprocessing
 //Convert RGB image to grayscale
-2 $I_{\mathrm{RGB}}\leftarrow 0.299R+0.587G+0.114B$
-//Retain SAR image as is
+$I_{\mathrm{RGB}}\leftarrow 0.299R+0.587G+0.114B$ //Retain SAR image as is
 3 $I_{\mathrm{SAR}}\leftarrow I_{\mathrm{SAR}}$
-4
 2: Spectral Decomposition
@@
   //Frequency domain conversion
-6       Compute $\mathcal{A}_{\mathrm{SAR}}^{(i)}$ and $\mathcal{P}_{\mathrm{SAR}}^{(i)}$ via Eqs. 11, 12 and 13
-7       Compute $\mathcal{A}_{\mathrm{RGB}}^{(i)}$ and $\mathcal{P}_{\mathrm{RGB}}^{(i)}$ similarly
+5  Compute $\mathcal{A}_{\mathrm{SAR}}^{(i)}$ and $\mathcal{P}_{\mathrm{SAR}}^{(i)}$ via 2D FFT and FFTshift; Compute $\mathcal{A}_{\mathrm{RGB}}^{(i)}$ and $\mathcal{P}_{\mathrm{RGB}}^{(i)}$ similarly
+ //Spatial difference computation
+  Compute $\Delta I^{(i)}\leftarrow\mathrm{RSD}^{(i)}$ via Eq. 1 //Amplitude difference computation
+  Compute $\Delta\mathcal{A}^{(i)}\leftarrow\mathrm{RAD}^{(i)}$ via Eq. 2 //Phase difference computation
+6  Compute $\Delta\mathcal{P}^{(i)}\leftarrow\mathrm{APPD}^{(i)}$ via Eq. 3
+7 end for
 8
-      //Spatial difference computation
-9       Compute $\Delta I^{(i)}\leftarrow\mathrm{RSD}^{(i)}$ via Eq. 1
-       //Amplitude difference computation
-10       Compute $\Delta\mathcal{A}^{(i)}\leftarrow\mathrm{RAD}^{(i)}$ via Eq. 2
-       //Phase difference computation
-11       Compute $\Delta\mathcal{P}^{(i)}\leftarrow\mathrm{APPD}^{(i)}$ via Eq. 3
-12
-13 end for
-14
 3: Dataset-level Aggregation
 //Mean spatial, amplitude, phase difference computation
-15 $\mathrm{RSD}\leftarrow\mathtt{Mean}(\{\Delta I^{(i)}\}_{i=1}^{N})$
-16 $\mathrm{RAD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{A}^{(i)}\}_{i=1}^{N})$
-17 $\mathrm{APPD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{P}^{(i)}\}_{i=1}^{N})$
+$\mathrm{RSD}\leftarrow\mathtt{Mean}(\{\Delta I^{(i)}\}_{i=1}^{N})$ $\mathrm{RAD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{A}^{(i)}\}_{i=1}^{N})$ $\mathrm{APPD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{P}^{(i)}\}_{i=1}^{N})$
 Algorithm 1 Spectral Difference Analysis.
 III-A Frequency-Domain Priors for Amplitude-Phase Decoupling
-To validate the physical significance of phase-amplitude decoupling and reveal the shared and complementary relations between RGB and SAR, we conducted a systematic spectral analysis on WHU-OPT-SAR dataset, the largest RGB-SAR segmentation benchmark.
-Firstly, we visualized the averaged difference spectra (see Fig.
-1(c)-(e)) using Algorithm 1 to compare modality differences in spatial, amplitude, and phase domains.
+To validate the physical significance of amplitude-phase decoupling and reveal the shared and complementary relations between RGB and SAR, we conduct a systematic spectral analysis on WHU-OPT-SAR, the largest RGB-SAR segmentation benchmark.
+First, we visualize averaged difference spectra (see Fig.
+1(c)-(e)) using Algorithm 1 to compare modality differences in the spatial, amplitude, and phase domains.
 This analysis shows that frequency-domain representations capture cross-modal sharing and complementarity more effectively than spatial representations.
@@
 $\mathrm{RAD}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|\mathcal{A}_{\mathrm{SAR}}^{(i)}-\mathcal{A}_{\mathrm{RGB}}^{(i)}\right|}{\mathcal{A}_{\mathrm{RGB}}^{(i)}+\epsilon},$ (2)
-$\mathrm{APPD}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\arg\exp j(\mathcal{P}_{\mathrm{SAR}}^{(i)}-\mathcal{P}_{\mathrm{RGB}}^{(i)})\right|,$ (3)
-where $I$, $\mathcal{A}$, and $\mathcal{P}$ denote pixel values, amplitude spectra, and phase spectra for SAR and RGB images; $\arg(\cdot)$ computes the complex angle; and $\epsilon$ prevents division by zero.
-By leveraging the periodicity of complex exponentials, APPD maps phase differences into $[-\pi,\pi)$, correcting wrap-around artifacts (e.g., $2\pi\to 0$).
-Figure 1(c)-(e) reveals three priors: 1) Spatial differences (RSD) exhibit anisotropic distributions.
-2) Amplitude differences (RAD) concentrate in high-frequency regions.
-3) Phase differences (APPD) remain minimal, with slight low-frequency clustering.
-These frequency-domain metrics demonstrate superior explainability over spatial RSD, confirming that spectral decomposition better captures cross-modal specificity.
+$\mathrm{APPD}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\arg\exp\left(\mathrm{i}(\mathcal{P}_{\mathrm{SAR}}^{(i)}-\mathcal{P}_{\mathrm{RGB}}^{(i)})\right)\right|,$ (3)
+where $I$ denotes pixel intensities; $\mathcal{A}$ and $\mathcal{P}$ are amplitude and phase spectra, respectively; $\arg(\cdot)$ returns the complex angle; $\mathrm{i}=\sqrt{-1}$; and $\epsilon>0$ prevents division by zero.
+By leveraging the periodicity of complex exponentials, Eq.
+3 maps phase differences to $\pi,\pi)$, correcting wrap-around artifacts (e.g., $2\pi\to 0$).
+Figure 1(c)-(e) reveals three priors: 1) spatial differences (RSD) exhibit anisotropic distributions; 2) amplitude differences (RAD) concentrate in high-frequency regions; and 3) phase differences (APPD) remain minimal, with slight low-frequency clustering.
+These frequency-domain metrics provide better explainability than spatial RSD, confirming that spectral decomposition more faithfully captures cross-modal characteristics.
 Physically, low-frequency phase disparities stem from geometric distortions (e.g., side-looking SAR vs.
-orthographic RGB), while high-frequency variations arise from SAR speckle noise.
-Amplitude disparities reflect modality-specific sensitivity to object scales: low frequencies encode large structures (farmlands, forests) and high frequencies capture fine details (roads).
-These analyses yield two key insights that drive our architecture design: 1) Phase Sharing: Consistently low APPD values indicate strong phase consistency.
-The PSC module thus performs pixel-level alignment while preserving this shared information.
-2) Amplitude Complementarity: RAD’s frequency-dependent divergence motivates ASF’s dynamic frequency-radius adjustment to handle amplitude variations.
-Secondly, we conducted sensitivity experiments on sampling rates (see Fig.
-1(f)-(g)) to evaluate the stability of our spectral metrics.
-By downsampling input images at various ratios and computing the mean and variance of RSD, RAD, and APPD across the dataset, we observed that RAD and APPD remain nearly constant, whereas RSD exhibits larger fluctuations.
-This invariance of RAD and APPD to sampling artifacts confirms that frequency-domain complementarity arises from intrinsic modality characteristics rather than from individual changes.
-TABLE I: Statistical Analysis of the Average Periodic Phase Difference Spectrum.
-Metric Low Frequency High Frequency Whole Image
+orthographic RGB), while high-frequency variations arise from SAR speckle.
+Amplitude disparities reflect modality-complementary sensitivity to object scales: low frequencies encode large structures (farmlands, forests), whereas high frequencies capture fine details (roads).
+These analyses yield two key insights that guide our design: Phase sharing.
+Consistently low APPD indicates strong phase consistency; PSC therefore performs pixel-level alignment while preserving this shared information.
+Amplitude complementarity.
+The frequency-dependent divergence of RAD motivates ASF’s dynamic frequency-radius adjustment to handle amplitude variations.
+Second, we conduct sensitivity experiments on sampling rates (see Fig.
+1(f)-(g)).
+By downsampling images at various ratios and computing the mean/variance of RSD, RAD, and APPD across the dataset, we observe that RAD and APPD remain nearly constant, whereas RSD fluctuates more.
+This invariance indicates that frequency-domain complementarity arises from intrinsic modality characteristics rather than sampling artifacts.
+TABLE I: Statistical Analysis of the Average Periodic Phase Difference (APPD) Spectrum.
+Metric
+Low frequency
+High frequency
+ALL
 Shapiro-Wilk statistic 1.0044 1.0050 1.0131
@@
 (a) APPD spectrum.
-(b)-(d) Frequency histograms with Gaussian fits for low-frequency (LF; radius = 0.5 $\times$ half-diagonal), high-frequency (HF), and all spectra (ALL).
-(e)-(f) Mean-variance trajectories across sampling rates.
-All spectra exhibit near-Gaussian invariance, with LF and ALL showing slight deviations due to structural noise in the low-frequency range, indicating that phase, particularly HF, can serve as a robust shared feature.
-Thirdly, to assess the statistical consistency of cross-modal phase spectra, we performed normality tests (Shapiro-Wilk) and calculated skewness and kurtosis for APPD values at each frequency band (see Table I).
+(b-d) Frequency histograms with Gaussian fits for low frequency (LF; radius = $0.5\times$ half-diagonal), high frequency (HF), and all spectra (ALL).
+(e-f) Mean-variance trajectories across sampling rates.
+Across sampling rates, all spectra exhibit near-Gaussian behavior, while slight deviations in LF and ALL are attributable to structural noise at low frequencies, indicating that the phase, particularly HF, serves as a robust shared feature.
+Third, we assess statistical consistency of cross-modal phase spectra via Shapiro-Wilk tests and by computing skewness and kurtosis of APPD within each frequency band (see Table I).
 Frequency histograms and mean-variance trajectories (see Fig.
-3) demonstrate that APPD distributions closely follow a Gaussian profile, especially in high-frequency regions.
-Low-frequency components show minor deviations, likely due to structural noise, but overall the near-Gaussian behavior validates phase spectra as a reliable shared feature.
-In a word, phase is so similar between RGB and SAR that it is almost unaffected by individual differences.
-Figure 4: Visualization of Spatial and Frequency Spectrum Differences under Missing RGB Regions.
-(a)-(b) RGB images with missing regions; (c) variation in RSD, RAD, and APPD, showing RSD’s greater sensitivity to missing data; (d)-(f) difference spectra including missing regions; (g)-(i) difference spectra excluding missing regions.
-These results demonstrate that frequency-domain metrics (RAD, APPD) are more robust to missing data than spatial-domain RSD.
-Figure 5: Visualization of Spatial and Frequency Spectrum Differences under Cloud Obscuration.
-(a)-(c) Example observations with cloud masks derived from RGB, SAR, and NIR channels in WHU-OPT-SAR [47].
-(d) Comparison of RSD, RAD, and APPD stability between cloud-covered and cloud-free areas; arrows denote the value of decline.
-Finally, we evaluated phase invariance under two challenging scenarios: 1) partial RGB data removal (see Fig.
-4) and 2) cloud-obscured observations (see Fig.
+3) show that APPD distributions closely follow a Gaussian profile, especially in high-frequency regions; small low-frequency deviations likely result from structural noise.
+In summary, RGB and SAR phases are highly similar and largely unaffected by instance-level variations.
+Figure 4: Visualization of Spatial-Frequency Spectrum Differences under Missing RGB Regions.
+(a-b) Two RGB images have missing regions in the original dataset.
+(c) Variation in RSD, RAD, and APPD, indicating RSD’s greater sensitivity to missing data.
+(d-f) Difference spectra computed including the missing regions.
+(g-i) Difference spectra computed excluding the missing regions.
+Overall, frequency-domain metrics (RAD and APPD) are more robust to missing data than the spatial-domain RSD.
+Figure 5: Visualization of Spatial-Frequency Spectrum Differences under Cloud Obscuration.
+(a-c) Example observations with cloud masks derived from the RGB, SAR, and NIR channels in WHU-OPT-SAR [51], respectively.
+(d) Comparison of RSD, RAD, and APPD between cloud-covered and cloud-free areas; arrows indicate the value of the decrease.
+Finally, we evaluate phase invariance under two challenging scenarios: 1) partial RGB removal (see Fig.
+4), and 2) cloud occlusion (see Fig.
 5).
-In the former, RSD increased dramatically when RGB pixels were missing (compare Fig.
-4(d) and (g)), while APPD remained stable (compare Fig.
+In the former, RSD increases markedly when RGB pixels are missing (see Fig.
+4(d) and (g)), whereas APPD remains stable (see Fig.
 4(f) and (i)).
-In the latter, APPD variance ($\Delta\approx 10^{-0.037}\%$) was negligible compared to that of RSD ($\Delta\approx 10^{5.81}\%$) and RAD ($\Delta\approx 10^{0.58}\%$) as shown in Fig.
+In the latter, the variance of APPD ($\Delta\approx 10^{-0.037}$) is negligible compared with that of RSD ($\Delta\approx 10^{5.81}$) and RAD ($\Delta\approx 10^{0.58}$) in Fig.
 5(d).
-These results confirm that phase information, particularly at high frequencies, retains robust consistency across modalities, underscoring its role as a shared feature to mitigate inter-modal conflicts and enhance complementary fusion.
+These results confirm that phase, particularly at high frequencies, retains robust cross-modal consistency, underscoring its role as a shared feature to mitigate inter-modal conflicts and enhance modality-complementary fusion.
 III-B Overall Framework
 As illustrated in Fig.
-2, the proposed PAD framework integrates SAR and RGB modalities via a dual-branch encoder and a shared decoder, with its primary innovation in cross-modal feature fusion through phase-amplitude decoupling.
-The encoder employs asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$, each with independent parameters to preserve modality-specific attributes.
-Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave scattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images.
-Both backbones produce hierarchical feature maps at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$ for $i\in\{0,1,2,3\}$:
-$x_{RGB}^{i}=\mathcal{B}_{\mathrm{RGB}}^{(i)}\left(I_{\mathrm{RGB}}^{i}\right),\quad x_{SAR}^{i}=\mathcal{B}_{\mathrm{SAR}}^{(i)}\left(I_{\mathrm{SAR}}^{i}\right).$ (4)
-Although frequency-domain fusion is more robust to modality differences, it can discard rich local interactions in the spatial domain.
-Therefore, we first introduce a lightweight SCF module to retain fine-grained spatial priors.
-The SCF module (see Eq.
-5) processes the bimodal features by computing pixel-wise fusion weights via spatial attention gating.
-This operation preliminarily integrates complementary features in the spatial domain and reduces feature dimensionality, constraining the solution space for subsequent frequency-domain operations:
-$X^{i}=\mathrm{SCF}\left(x_{RGB}^{i},x_{SAR}^{i}\right),$ (5a)
-$\mathcal{S}=\mathrm{Sigmoid}\left(\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1},x_{2}\right)\right)\right),$ (5b)
+2, PAD integrates SAR and RGB via a dual-branch encoder and a shared decoder, with its primary innovation lying in cross-modal feature fusion through phase-amplitude decoupling.
+The encoder employs asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$, each with independent parameters to preserve modality-complementary attributes.
+Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave backscattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images.
+Both backbones produce hierarchical features at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$ for $i\in\{0,1,2,3\}$:
+$x_{\mathrm{RGB}}^{i}=\mathcal{B}_{\mathrm{RGB}}^{(i)}\left(I_{\mathrm{RGB}}\right),\quad x_{\mathrm{SAR}}^{i}=\mathcal{B}_{\mathrm{SAR}}^{(i)}\left(I_{\mathrm{SAR}}\right).$ (4)
+Although frequency-domain fusion is robust to modality differences, it may discard rich local interactions.
+To retain fine-grained spatial priors, we introduce a lightweight SCF module.
+SCF computes pixel-wise fusion weights via spatial attention gating and reduces feature dimensionality to constrain the solution space for subsequent spectral operations:
+$X^{i}$ $=\mathrm{SCF}\left(x_{\mathrm{RGB}}^{i},x_{\mathrm{SAR}}^{i}\right),$ (5a)
+$\mathcal{S}$ $=\sigma\left(\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1},x_{2}\right)\right)\right),$ (5b)
 $x_{1}^{\prime}$ $=\mathcal{S}\odot x_{1},\quad x_{2}^{\prime}=\left(1-\mathcal{S}\right)\odot x_{2},$ (5c)
-$x=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1}^{\prime},x_{2}^{\prime}\right)\right).$ (5d)
-The fused feature $X^{i}$ is then transformed via FD into separate amplitude $\mathcal{A}^{i}$ and phase $\mathcal{P}^{i}$ spectra:
-$\mathcal{A}^{i},\mathcal{P}^{i}=\mathrm{FD}\left(X^{i}\right).$ (6)
-The amplitude spectrum $\mathcal{A}^{i}$ is enhanced by the ASF module, which uses a trainable frequency radius $r$ to distinguish and reinforce high-frequency details while preserving low-frequency structures:
-$\mathcal{A^{\prime}}^{i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right),$ (7)
-where the PSC module applies multiplicative scaling to correct the phase:
-$\mathcal{P^{\prime}}^{i}=\mathrm{PSC}\left(\mathcal{P}^{i}\right).$ (8)
-Finally, FR reconstructs the fused features:
-$\bm{F^{\prime}}^{i}=\mathrm{FR}\left(\mathcal{A^{\prime}}^{i},\mathcal{P^{\prime}}^{i}\right).$ (9)
-The shared decoder $\mathcal{D}$ then aggregates the multi-stage fused features to generate high-precision land-cover segmentation maps:
-$\bm{\hat{Y}}=\mathcal{D}\left(\bm{F^{\prime}}^{0},\bm{F^{\prime}}^{1},\bm{F^{\prime}}^{2},\bm{F^{\prime}}^{3}\right).$ (10)
-III-C Frequency Decoupling & Frequency Recoupling
-To achieve physically disentangled representations in the spectral domain, we propose a bidirectional transformation framework comprising FD and its inverse, FR.
-First, the input image $X\in\mathbb{R}^{H\times W}$ is transformed from the spatial domain to the frequency domain using a 2D real-input fast Fourier transform (rFFT) [48], yielding the spectrum $\mathcal{F}\in\mathbb{C}^{H\times(\lfloor W/2\rfloor+1)}$:
-$\mathcal{F}\left(u,v\right)=\sum_{y=0}^{H-1}\sum_{x=0}^{W-1}X\left(y,x\right)\exp\left(-i2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
-where $(u,v)$ are frequency-domain coordinates.
-Exploiting the conjugate symmetry of real-valued signals, we retain only the non-redundant region ($0\leq u\leq\lfloor W/2\rfloor$, $0\leq v<H$).
-Next, an FFT shift relocates low-frequency components to the spectrum center:
-$\mathcal{F}_{\mathrm{shift}}\left(u,v\right)=\mathcal{F}\left(u\oplus\frac{W}{2},v\oplus\frac{H}{2}\right),$ (12)
-where $\oplus$ denotes modulo addition. Finally, we derive the amplitude and phase spectra as
-$\mathcal{A}=|\mathcal{F}|,\quad\mathcal{P}=\angle\mathcal{F},$ (13)
-where $\mathcal{A},\mathcal{P}\in\mathbb{R}^{H\times(\lfloor W/2\rfloor+1)}$ capturing radiometric and structural information, respectively.
-This decoupling allows the model to focus on modality-specific representations: SAR phase carries scattering patterns sensitive to physical surface characteristics, whereas RGB amplitude captures fine-grained textures and semantic content.
-By separating them, the network avoids early-stage interference between heterogeneous modalities, enhancing feature discriminability before integration.
-III-D Phase Spectrum Correction Module
-While previous works have assumed perfect phase alignment [17], our spectral analysis (see Fig.
-1(e)) reveals that low-frequency geometric inconsistencies persist, necessitating a learnable, spatially aware correction mechanism.
+$x$ $=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1}^{\prime},x_{2}^{\prime}\right)\right),$ (5d)
+where $(x_{1},x_{2})=(x_{\mathrm{RGB}}^{i},x_{\mathrm{SAR}}^{i})$; $\sigma(\cdot)$ is the Sigmoid; $\mathrm{Conv}_{1\times 1}$ denotes a $1\times 1$ convolution; $\mathrm{Concat}(\cdot,\cdot)$ concatenates along channels; and $\odot$ denotes the Hadamard (element-wise) product.
+The fused feature $X^{i}$ is then transformed to separate amplitude $\mathcal{A}^{i}$ and phase $\mathcal{P}^{i}$ spectra:
+$\mathcal{A}^{i},\mathcal{P}^{i}=\mathrm{FD}\left(X^{i}\right),$ (6)
+where $\mathrm{FD}(\cdot)$ denotes the frequency decoupling operator (defined in §III-C).
+ASF enhances the amplitude spectrum using a trainable frequency radius $r$ to distinguish and reinforce high-frequency details while preserving low-frequency structures:
+$\mathcal{A}^{\prime i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right).$ (7)
+PSC applies multiplicative scaling to correct the phase:
+$\mathcal{P}^{\prime i}=\mathrm{PSC}\left(\mathcal{P}^{i}\right).$ (8)
+Finally, inverse reconstruction produces fused features:
+$\bm{F}^{\prime i}=\mathrm{FR}\left(\mathcal{A}^{\prime i},\mathcal{P}^{\prime i}\right),$ (9)
+where $\mathrm{FR}(\cdot,\cdot)$ is the inverse frequency recoupling operator (defined in §III-C). A shared decoder $\mathcal{D}$ aggregates multi-stage fused features to generate the land-cover map:
+$\widehat{\bm{Y}}=\mathcal{D}\left(\bm{F}^{\prime 0},\bm{F}^{\prime 1},\bm{F}^{\prime 2},\bm{F}^{\prime 3}\right).$ (10)
+III-C Frequency Decoupling and Recoupling
+To obtain physically decoupled spectral representations, we employ a bidirectional transformation comprising frequency decoupling (FD) and inverse recoupling (FR).
+Given $X\in\mathbb{R}^{H\times W}$, a 2D real-input fast Fourier transform (rFFT) [52] yields:
+$\mathcal{F}(u,v)=\sum_{y=0}^{H-1}\sum_{x=0}^{W-1}X(y,x)\exp\left(-\mathrm{i}2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
+where $(u,v)$ are frequency coordinates. Exploiting conjugate symmetry for real inputs, we retain only the non-redundant region and center low-frequency components via an FFT shift:
+$\mathcal{F}_{\mathrm{shift}}(u,v)=\mathcal{F}\left(u\oplus\frac{W}{2},v\oplus\frac{H}{2}\right),$ (12)
+where $\oplus$ is modulo addition. Amplitude and phase are:
+$\mathcal{A}=\left|\mathcal{F}_{\mathrm{shift}}\right|,\quad\mathcal{P}=\angle\mathcal{F}_{\mathrm{shift}},$ (13)
+with $\mathcal{A},\mathcal{P}\in\mathbb{R}^{H\times(\lfloor W/2\rfloor+1)}$. FR applies the inverse operations to reconstruct spatial features from $(\mathcal{A},\mathcal{P})$.
+III-D Phase Spectrum Correction (PSC) Module
+While prior works often assume perfect phase alignment [18], our spectral analysis (see Fig.
+1(e)) shows that low-frequency geometric inconsistencies persist, necessitating a learnable, spatially aware correction mechanism.
 Inspired by channel attention, we propose a lightweight Phase Spectrum Correction (PSC) module tailored for frequency-domain phase alignment.
-As illustrated in Fig.
+As shown in Fig.
 1(e), minor misalignments in the low-frequency phase spectra between SAR and RGB, attributable to geometric distortions (e.g., side-looking SAR vs.
-RGB orthographic projection), can compromise spatial consistency and boundary precision.
-Traditional rigid registration methods are inadequate for such non-rigid distortions.
-Thus, the PSC module performs adaptive geometric correction using learnable multiplicative masks.
-Unlike the global channel scaling in traditional attention mechanisms, the proposed PSC estimates a spatially varying modulation mask via a stack of $d$ consecutive $1\times 1$ convolutions, enabling localized, fine-grained phase adjustment.
-Given an input phase-spectrum tensor $P\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$, we first compress the channel dimension via $d$ $1\times 1$ convolution (reducing $C$ to $C/r$, $r\in\{4,8\}$) followed by leaky linear rectification function (Leaky ReLU) [49].
-A second $1\times 1$ convolution produces a single-channel Phase Correction Mask (PCM), constrained to $(0,1)$ by a Sigmoid:
-$\mathrm{PCM}=\sigma\left(\mathrm{Conv}_{d}^{1\times 1}\left(\delta\left(\cdots\delta\left(\mathrm{Conv}_{1}^{1\times 1}\left(P\right)\right)\cdots\right)\right)\right),$ (14)
-where $d$ denotes the number of stacked convolutions and $\delta(\cdot)$ is Leaky ReLU. Pixel-wise multiplicative scaling is then applied:
-$P^{\prime}=P\odot\left(1+\mathrm{PCM}\right),$ (15)
-where $\odot$ is the Hadamard product. This formulation enables localized, fine-grained phase adjustment while preserving phase periodicity and incurs minimal computational overhead.
-III-E Amplitude Spectrum Fusion Module
-Previous frequency-domain methods often apply fixed high- or low-pass filters or spatially uniform spectral attention, failing to adapt to content-specific frequency distributions.
+orthographic RGB), can compromise spatial consistency and boundary precision.
+Traditional rigid registration is inadequate for such non-rigid distortions.
+PSC performs adaptive geometric correction using learnable multiplicative masks.
+Unlike global channel scaling in traditional attention, PSC estimates a spatially varying modulation mask via a stack of $d$ consecutive $1{\times}1$ convolutions, enabling localized, fine-grained phase adjustment.
+Given an input phase-spectrum tensor $P\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$, we first compress channels with $d$ successive $1{\times}1$ convolutions (reducing $C$ to $C/r$, $r\in\{4,8\}$), each followed by a Leaky ReLU [53].
+A final $1{\times}1$ convolution produces a single-channel Phase Correction Mask (PCM), constrained to $(0,1)$ by a Sigmoid:
+$\mathrm{PCM}=\sigma\left(\mathrm{Conv}^{1\times 1}_{d}\left(\delta\left(\cdots\delta\left(\mathrm{Conv}^{1\times 1}_{1}\left(P\right)\right)\cdots\right)\right)\right),$ (14)
+where $\delta(\cdot)$ is Leaky ReLU and $\mathrm{Conv}^{1\times 1}_{k}$ denotes the $k$-th $1{\times}1$ convolution in the stack. Pixel-wise multiplicative scaling is then applied:
+$P^{\prime}=P\odot\left(1+\mathrm{PCM}\right).$ (15)
+This formulation enables localized, fine-grained phase adjustment while preserving phase periodicity and incurring minimal computational overhead.
+III-E Amplitude Spectrum Fusion (ASF) Module
+Previous frequency-domain methods often employ fixed high-/low-pass filters or spatially uniform spectral attention, which cannot adapt to content-specific frequency distributions.
 In contrast, our analysis in Fig.
-1(d) shows that the spectral discrepancy between SAR and RGB modalities is primarily concentrated in the medium-to-high frequency components of the amplitude spectra.
-To exploit this prior while maintaining gradient stability, we design the ASF module to employ a trainable frequency radius for dynamic separation and enhancement of complementary frequency bands via three sequential operations.
-First, we compute a normalized frequency distance metric for each spectral component $i$:
-$\mathrm{Dist}(i)=\frac{\sqrt{\left(X_{i}-X_{\mathrm{center}}\right)^{2}+\left(Y_{i}-Y_{\mathrm{center}}\right)^{2}}}{\sqrt{X_{\mathrm{max}}^{2}+Y_{\mathrm{max}}^{2}}},$ (16)
-where quantifies component positions relative to the spectrum center $(X_{\mathrm{center}},Y_{\mathrm{center}})$.
-A trainable boundary radius $r$, constrained via a Sigmoid activation to yield $R=\sigma(r)\in[0,1]$, dynamically partitions the spectrum into low- and high-frequency regions.
-This constraint matches the normalized spectral dimensions and prevents $R$ from diverging to unreasonable values during training, which would otherwise lead to unstable partitioning and degraded performance.
-To maintain differentiability, we employ a soft partition:
-$\mathrm{Mask}_{\mathrm{high-freq}}=\sigma\left(\left(\mathrm{Dist}-R\right)\cdot\tau\right),$ (17)
-where temperature $\tau=10$. High-frequency components are enhanced via a position-wise MLP:
-$A_{\mathrm{high-freq}}=\mathrm{MLP}\left(A\odot\mathrm{Mask}_{\mathrm{high-freq}}\right),$ (18)
-then concatenated with the original amplitude $A$ and fused by a $1\times 1$ convolution and residual connection:
-$A^{\prime}=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(A,A_{\mathrm{high-freq}}\right)\right),\quad A\leftarrow A+A^{\prime}.$ (19)
-This module preserves low-frequency structures while adaptively reinforcing high-frequency details that capture complementary semantic information.
+1(d) shows that RGB-SAR spectral discrepancies are concentrated in medium-to-high amplitude frequencies.
+To exploit this prior while maintaining gradient stability, ASF uses a trainable frequency radius to dynamically separate and enhance complementary bands via three steps.
+First, for each spectral component indexed by $i$, we compute a normalized radial distance:
+$\mathrm{Dist}(i)=\frac{\sqrt{\left(X_{i}-X_{\mathrm{c}}\right)^{2}+\left(Y_{i}-Y_{\mathrm{c}}\right)^{2}}}{\sqrt{X_{\max}^{2}+Y_{\max}^{2}}},$ (16)
+where $(X_{\mathrm{c}},Y_{\mathrm{c}})$ is the spectrum center and $(X_{\max},Y_{\max})$ are the maximal coordinate indices along each axis, so $\mathrm{Dist}(i)\in[0,1]$ quantifies each component’s position relative to the center.
+A trainable boundary radius $r$ is mapped by a Sigmoid to $R=\sigma(r)\in[0,1]$, partitioning the spectrum into low-/high-frequency regions and preventing divergence during training.
+To maintain differentiability, we use a soft mask:
+$\mathrm{Mask}_{\mathrm{hi}}=\sigma\left(\left(\mathrm{Dist}-R\right)\cdot\tau\right),$ (17)
+where $\tau=10$ is a temperature hyperparameter. High-frequency components are enhanced via a position-wise MLP:
+$A_{\mathrm{hi}}=\mathrm{MLP}\left(A\odot\mathrm{Mask}_{\mathrm{hi}}\right),$ (18)
+and then concatenated with the original amplitude $A$ and fused by a $1{\times}1$ convolution with a residual update:
+$A^{\prime}=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(A,A_{\mathrm{hi}}\right)\right),\quad A\leftarrow A+A^{\prime},$ (19)
+where $A\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$ is the input amplitude spectrum, and $\mathrm{MLP}(\cdot)$ is a channel-wise two-layer perceptron with nonlinearity.
+This preserves low-frequency structure while adaptively reinforcing high-frequency details carrying modality-complementary semantics.
 The learnable radius $R$ self-adjusts to input characteristics, eliminating manual threshold tuning.
 III-F Loss Function
-Our optimization framework integrates three objectives to jointly promote semantic segmentation accuracy, multi-scale feature learning, and amplitude-phase coherence.
+We integrate three objectives to jointly promote semantic accuracy, multi-scale representation learning, and amplitude-phase coherence.
 III-F1 Primary Semantic Supervision
-The primary segmentation loss is defined as the standard cross-entropy loss:
 $\mathcal{L}_{\mathrm{seg}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}_{i,c},$ (20)
-where $\Omega$ denotes the set of foreground pixels (excluding background label 255), $y_{i,c}\in\{0,1\}$ indicates the ground-truth class, and $\hat{p}_{i,c}$ is the predicted probability.
+where $\Omega$ is the set of labeled pixels (excluding the ignore label 255), $y_{i,c}\in\{0,1\}$ is the one-hot ground truth for class $c$, and $\hat{p}_{i,c}$ is the predicted probability.
 III-F2 Auxiliary Semantic Supervision
-An auxiliary segmentation head is added at the second-to-last stage ($t=2$), also using cross-entropy loss:
-$\mathcal{L}_{\mathrm{aux}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}^{(a)}_{i,c}.$ (21)
+An auxiliary head at the penultimate stage ($t=2$) uses the same objective:
+$\mathcal{L}_{\mathrm{aux}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}^{(a)}_{i,c}$, (21)
+where $\hat{p}^{(a)}_{i,c}$ denotes auxiliary predictions.
 III-F3 Phase-Amplitude Decoupling Constraint
-To resolve the coupling between amplitude and phase spectra and enforce consistency between fused and original amplitudes, we introduce an amplitude consistency loss that regularizes $\mathcal{A}^{\prime}_{t}$ to remain faithful to its original decoupled state $\mathcal{A}_{t}$ across feature scales:
-$\mathcal{L}_{\mathrm{amp}}=\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}\alpha_{t}\left\|\mathcal{A}^{\prime}_{t}-\mathcal{A}_{t}\right\|_{F}^{2},$ (22)
-where $\mathcal{T}=\{0,1,2,3\}$ indexes the feature scales, $\alpha_{t}=0.5+0.5\frac{t}{3}$ imposes depth-progressive weighting, and $\|\cdot\|_{F}$ is the Frobenius norm.
+To regularize the fused amplitude towards its decoupled counterpart across scales, we define:
+$\mathcal{L}_{\mathrm{amp}}=\frac{1}{\left|\mathcal{T}\right|}\sum_{t\in\mathcal{T}}\alpha_{t}\left\|\mathcal{A}^{\prime}_{t}-\mathcal{A}_{t}\right\|_{F}^{2},$ (22)
+where $\mathcal{T}=\{0,1,2,3\}$ indexes feature scales; $\alpha_{t}=0.5+0.5t/3$ applies depth-progressive weighting; $\|\cdot\|_{F}$ is the Frobenius norm; and $\mathcal{A}_{t},\mathcal{A}^{\prime}_{t}$ are the original and fused amplitude spectra at scale $t$.
 III-F4 Composite Objective
-The overall loss is
 $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{seg}}+\lambda_{1}\mathcal{L}_{\mathrm{aux}}+\lambda_{2}\mathcal{L}_{\mathrm{amp}},$ (23)
-with $\lambda_{1}=0.4$ balancing auxiliary supervision and $\lambda_{2}=0.1$ controlling the amplitude regularization strength.
-This composite loss promotes both geometric invariance, by stabilizing phase-sensitive features through $\mathcal{L}_{\mathrm{amp}}$, and semantic invariance via hierarchical amplitude refinement.
+where $\lambda_{1}=0.4$ balances auxiliary supervision and $\lambda_{2}=0.1$ controls amplitude regularization.
+This composite loss promotes geometric invariance by stabilizing phase-sensitive features via $\mathcal{L}_{\mathrm{amp}}$, and semantic robustness via hierarchical amplitude refinement.
 IV Experimental Results
-IV-A Dataset Description
-WHU-OPT-SAR [28] is a multi-modal dataset for land use classification in Hubei Province, China.
-It contains 100 co-registered RGB, NIR, and SAR image triples with corresponding land type labels.
-RGB images are acquired by GF-1, and SAR images by GF-3; both are resampled to 5 m resolution and sized 5556 $\times$ 3704 pixels.
-Seven primary land use categories are covered: farmland, urban areas, villages, water bodies, forests, roads, and others.
-We crop each image into non-overlapping 512 $\times$ 512 patches, yielding 7,000 patches, of which 80% (5,600) are used for training and 20% (1,400) for testing.
-DDHR-SK is a subset of the DDHR multi-modal dataset for land cover classification in Pohang City, South Korea [29].
-RGB images from GF-2 and SAR images from GF-3 are paired, and RGB images are processed with a fog filter in GIMP to simulate adverse conditions.
-The dataset comprises 6,173 cloud-affected RGB-SAR pairs, resampled to 1 m resolution and resized to 256 $\times$ 256 pixels.
-It includes five land cover types: buildings, farmland, greenery, water, and roads, with 3,087 training and 3,086 validation images in the public split.
-IV-B Training Setting
-All models were trained using the MMSegmentation framework [50] on a single NVIDIA H800 GPU (comparisons used an NVIDIA A100).
-We used a batch size of 4 for 80 k iterations, validating every 8 k iterations.
-Optimization was performed with AdamW [51], an initial learning rate of $1\times 10^{-4}$, and a cosine annealing schedule (T_max = total iterations).
-A fixed seed (42) ensured reproducibility.
-We report mean IoU (mIoU), Kappa coefficient, overall accuracy (OA), and F1 score, selecting the iteration with the best mIoU.
-To mitigate class-imbalance bias, we compute the Kappa coefficient per sample and report its mean:
-$\kappa_{i}=\frac{P_{o}^{(i)}-P_{e}^{(i)}}{1-P_{e}^{(i)}},\quad\mathrm{mKappa}=\frac{1}{M}\sum_{i=1}^{M}\kappa_{i},$ (24)
-where $P_{o}^{(i)}$ and $P_{e}^{(i)}$ are the observed and expected agreements for sample $i$.
-IV-C Experimental Results on Datasets
-To assess the effectiveness of PAD, we replicated the original configurations of 17 comparison models, nine state-of-the-art multi-modal segmentation approaches and eight leading unimodal methods, on both WHU-OPT-SAR and DDHR-SK.
+IV-A Experimental Setup
+IV-A1 Dataset Description
+WHU-OPT-SAR [29] is a multi-modal land-use dataset covering Hubei Province, China.
+It contains 100 co-registered RGB-SAR-NIR triples labeled with seven categories: farmland, urban, village, water, forest, roads, and others.
+RGB (GF-1) and SAR (GF-3) images are resampled to 5 m resolution and sized $5556\times 3704$ pixels.
+Each image is cropped into non-overlapping $512\times 512$ patches, yielding 7,000 samples split 80%/20% for training/testing.
+DDHR-SK [30] is a fog-degraded subset of DDHR from Pohang, South Korea, pairing GF-2 RGB and GF-3 SAR images.
+It contains 6,173 cloud-affected RGB-SAR pairs, resampled to 1 m and resized to $256\times 256$ pixels, with five categories: buildings, farmland, greenery, water, and roads.
+The public split provides 3,087 training and 3,086 validation samples.
+IV-A2 Training Settings
+All models are trained using the MMSegmentation framework [54] on a single NVIDIA H800 GPU (comparisons use an NVIDIA A800).
+We use a batch size of 4 for 80k iterations, validating every 8k iterations.
+Optimization is performed with AdamW, an initial learning rate of $1\times 10^{-4}$, and a cosine-annealing schedule (T_max = total iterations).
+A fixed seed (42) ensures reproducibility.
+IV-A3 Evaluation Metrics
+We report mean Intersection-over-Union (mIoU), overall accuracy (OA), F1 score, and Cohen’s Kappa (mKappa), selecting the checkpoint with the highest mIoU.
+To mitigate class-imbalance bias, Kappa is computed for each test sample and then averaged across all samples to obtain mKappa.
+IV-B Main Quantitative Results
+To assess PAD’s effectiveness, we replicate the original configurations of 20 comparison models, 12 state-of-the-art multi-modal segmentation approaches and 8 leading unimodal methods, on both WHU-OPT-SAR and DDHR-SK.
 TABLE II: Quantitative Results (%) on WHU-OPT-SAR.
-We report per-category IoU for seven land cover classes (Farmland, City, Village, Water, Forest, Road, Others), along with OA, mKappa, mF1, and mIoU.
-Bold and underlined values indicate the best and second-best performance, respectively.
-Method Venue IoU per Category OA mKappa mF1 mIoU
+Per-class IoU is reported for seven land-cover categories (Farmland, City, Village, Water, Forest, Road, and Others), along with OA, mKappa, mF1, and mIoU.
+The best and second-best results are highlighted in bold and underlined, respectively.
+Method Venue Per-class IoU OA mKappa mF1 mIoU
 Farmland City Village Water Forest Road Others
-RGB DANet [18] CVPR’19 67.51 55.22 46.71 60.85 83.13 38.89 20.43 83.19 49.74 67.40 53.25
+RGB
+DANet [19] CVPR’19 67.51 55.22 46.71 60.85 83.13 38.89 20.43 83.19 49.74 67.40 53.25
+CGNet [21] TIP’20 61.33 45.96 37.98 49.33 80.98 1.92 0.19 79.26 37.42 50.54 39.67
 OCRNet [20] ECCV’20 67.68 56.48 47.47 61.22 82.98 37.20 17.97 83.23 49.11 66.95 53.00
-CGNet [20] TIP’20 61.33 45.96 37.98 49.33 80.98 1.92 0.19 79.26 37.42 50.54 39.67
+SETR [23] CVPR’21 65.25 55.07 42.97 57.80 82.44 31.91 18.36 81.94 46.19 64.73 50.54
 SegFormer [22] NeurIPS’21 68.18 56.91 46.82 60.85 83.19 35.24 18.73 83.37 49.46 66.79 52.84
-SETR [22] CVPR’21 65.25 55.07 42.97 57.80 82.44 31.91 18.36 81.94 46.19 64.73 50.54
 Mask2Former [24] CVPR’22 67.09 56.64 43.97 60.09 81.47 36.24 21.03 82.69 47.69 66.65 52.36
@@
 TransUNet [26] MIA’24 67.62 56.28 46.08 60.66 83.25 38.76 19.32 83.18 49.89 67.20 53.14
-SAR DANet [18] CVPR’19 60.42 54.12 33.05 54.45 80.57 24.89 15.11 79.44 35.28 60.16 46.09
+SAR
+DANet [19] CVPR’19 60.42 54.12 33.05 54.45 80.57 24.89 15.11 79.44 35.28 60.16 46.09
+CGNet [21] TIP’20 56.72 46.81 26.32 46.30 77.53 0.27 0.00 76.39 24.64 47.00 36.28
 OCRNet [20] ECCV’20 59.81 52.48 31.21 54.03 79.57 22.49 13.92 78.94 32.62 58.74 44.79
-CGNet [20] TIP’20 56.72 46.81 26.32 46.30 77.53 0.27 0.00 76.39 24.64 47.00 36.28
+SETR [23] CVPR’21 60.35 52.29 31.41 53.56 80.15 18.66 15.18 79.12 34.25 58.33 44.51
 SegFormer [22] NeurIPS’21 60.17 54.71 32.30 54.07 79.79 21.02 15.13 79.31 33.60 59.24 45.31
-SETR [22] CVPR’21 60.35 52.29 31.41 53.56 80.15 18.66 15.18 79.12 34.25 58.33 44.51
 Mask2Former [24] CVPR’22 60.39 52.86 35.69 53.85 78.36 29.16 17.20 79.26 32.76 61.35 46.79
@@
 TransUNet [26] MIA’24 61.93 54.85 33.42 56.26 80.63 27.56 14.96 80.14 37.45 61.13 47.09
-SAR & RGB MCANet [28] JAG’21 64.21 45.49 42.93 56.20 81.81 25.93 8.47 80.80 45.47 59.94 46.44
-CEN [33] TPAMI’22 66.46 56.28 45.46 58.09 82.34 32.51 17.82 82.32 47.47 65.36 51.28
-TokenFusion [52] CVPR’22 67.87 57.26 46.96 61.66 83.31 36.34 18.76 83.38 49.87 67.09 53.16
-DDHRNet [29] JAG’22 59.12 45.67 30.77 47.96 79.59 0.47 0.03 78.00 27.18 48.36 37.66
-CDDFuse [12] CVPR’23 59.88 35.08 25.22 52.31 78.77 17.64 15.00 77.02 38.69 54.29 40.56
-CMX [31] TITS’23 69.24 58.20 48.45 64.06 83.83 39.99 21.92 84.18 52.76 69.01 55.10
-MRFS [30] CVPR’24 69.39 58.13 49.43 64.90 83.87 41.26 21.72 84.21 54.33 69.38 55.53
-FTransUNet [6] TGRS’24 64.87 53.84 43.71 55.63 81.53 0.00 0.00 81.25 43.53 52.97 42.80
-ASANet [32] ISPRS’24 69.18 56.95 48.58 64.06 83.79 41.48 21.74 84.15 52.62 69.05 55.11
+SAR & RGB
+MCANet [29] JAG’21 65.14 53.97 45.43 54.59 81.87 31.37 13.43 81.74 44.28 63.37 49.40
+DDHRNet [30] JAG’22 64.80 49.42 40.98 54.11 81.31 22.96 10.32 80.94 44.49 59.84 46.27
+CEN [34] TPAMI’22 67.60 55.73 46.46 60.65 83.28 34.22 15.94 83.17 49.37 65.79 51.98
+TokenFusion [55] CVPR’22 68.43 58.34 45.76 61.09 83.79 32.03 15.20 83.65 49.45 65.67 52.09
+CDDFuse [13] CVPR’23 60.07 48.64 23.96 50.67 79.47 15.16 11.43 78.02 51.36 54.54 41.34
+CMX [32] TITS’23 70.04 58.41 49.04 64.77 84.12 41.76 22.21 84.53 53.46 69.60 55.76
+FTransUNet [6] TGRS’24 68.22 56.28 46.76 62.35 83.46 35.35 17.23 83.50 50.94 66.61 52.81
+ASANet [33] ISPRS’24 69.71 58.09 49.05 63.52 84.05 41.50 20.57 84.29 52.83 69.04 55.21
+MRFS [31] CVPR’24 69.59 58.04 49.15 63.78 84.15 41.26 22.88 84.24 54.14 69.48 55.55
+ASMFNet [56] JSTARS’24 67.31 55.73 46.56 59.51 83.03 32.63 14.60 82.91 49.51 65.09 51.34
+MIEFNet [57] GRSL’24 68.72 57.16 48.04 63.23 83.64 39.35 18.41 83.87 50.58 67.89 54.08
+MFNet [58] TGRS’25 68.97 58.39 47.78 63.38 83.88 39.51 20.33 84.02 51.86 68.47 54.61
 PAD (Ours) 69.85 58.10 49.70 65.39 84.05 43.24 23.47 84.56 53.93 70.14 56.26
-IV-C1 Evaluations on WHU-OPT-SAR
+IV-B1 Evaluations on WHU-OPT-SAR
 As shown in Table II, PAD achieves new state-of-the-art results: 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1.
-This surpasses the best unimodal method (DANet-RGB [18]: 53.25% mIoU, 83.19% OA) and the multi-modal baseline MRFS [30] (55.53% mIoU, 84.21% OA).
-In fine-grained classes, PAD attains 43.24% IoU on roads and 49.70% on villages, improving over MRFS by 1.98% and 0.27%, respectively.
+This surpasses the best unimodal method (DANet-RGB [19]: 53.25% mIoU, 83.19% OA) and the multi-modal baseline CMX [32] (55.76% mIoU, 84.53% OA).
+For fine-grained classes, PAD attains 43.24% IoU on roads and 49.70% on villages, improving over MRFS by 1.48% and 0.66%, respectively.
 Figure 6: Visualization Results on WHU-OPT-SAR.
-Four representative scenarios: (a) building-dense area, (b) specular water bodies, (c) vegetated region, and (d) urban-rural transition.
-For each scenario, images are shown (left to right): RGB, SAR, pseudo-color ground truth, and prediction maps from six competing methods.
-We compare PAD against single-modal baselines (DANet-RGB, TransUNet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS).
-Regions of interest are highlighted with gray boxes (buildings/water) and circles (vegetation); see Section IV-C for details.
-Visual results in Fig.
-6 further illustrate PAD’s benefits.
-In (a), PAD recovers occluded road segments (cyan) in building-dense areas via SAR phase cues, whereas RGB-only models miss them.
-In (b), PAD distinguishes specular reflections from true water boundaries under cloud interference.
-In (c)-(d), PAD resolves vegetation spectral ambiguities and maintains consistency across urban-rural transitions.
-TABLE III: Quantitative Results (%) on DDHR-SK. We report per-category IoU for seven land cover classes (Buildings, Roads, Greenery, Water, Farmland), along with OA, mKappa, mF1, and mIoU.
-Method Venue IoU per Category OA mKappa mF1 mIoU
+Four representative scenarios: (a) building-dense area; (b) specular water bodies; (c) vegetated region; and (d) urban-rural transition.
+For each scenario, images are shown (left to right): RGB, SAR, pseudo-color ground truth, and predictions from six competing methods.
+We compare PAD against unimodal baselines (DANet-RGB and TransUNet-SAR) and multi-modal approaches (CDDFuse, CMX, and MRFS).
+Regions of interest are highlighted with gray boxes (buildings/water) and circles (vegetation); see §IV-B for details.
+Qualitative results in Fig.
+6 further highlight PAD’s benefits.
+In Fig.
+6(a), PAD recovers occluded road segments (cyan) in building-dense areas via SAR phase cues, whereas RGB-only models miss them.
+In Fig.
+6(b), PAD distinguishes specular reflections from true water boundaries under cloud interference.
+In Fig.
+6(c)-(d), PAD resolves vegetation spectral ambiguities and maintains consistency across urban-rural transitions.
+TABLE III: Quantitative Results (%) on DDHR-SK. Per-class IoU is reported for five land-cover categories (Buildings, Roads, Greenery, Water, and Farmland), along with OA, mKappa, mF1, and mIoU.
+Method Venue Per-class IoU OA mKappa mF1 mIoU
 Buildings Roads Greenery Water Farmland
-RGB DANet [18] CVPR’19 94.46 75.11 91.10 99.03 90.59 96.23 73.35 94.57 90.06
+RGB
+DANet [19] CVPR’19 94.46 75.11 91.10 99.03 90.59 96.23 73.35 94.57 90.06
+CGNet [21] TIP’20 80.19 27.57 72.77 90.47 69.97 86.16 33.90 78.76 68.19
 OCRNet [20] ECCV’20 90.18 60.49 84.42 97.87 82.84 93.17 58.96 90.26 83.16
-CGNet [20] TIP’20 80.19 27.57 72.77 90.47 69.97 86.16 33.90 78.76 68.19
+SETR [23] CVPR’21 94.17 74.10 91.43 99.18 90.88 96.22 73.79 94.49 89.95
 SegFormer [22] NeurIPS’21 92.44 67.11 88.76 98.82 88.18 95.00 66.73 92.71 87.06
-SETR [22] CVPR’21 94.17 74.10 91.43 99.18 90.88 96.22 73.79 94.49 89.95
 Mask2Former [24] CVPR’22 94.16 80.46 91.08 98.68 86.88 96.07 74.05 94.76 90.25
@@
 TransUNet [26] MIA’24 90.67 62.68 85.64 98.41 83.71 93.64 61.82 90.95 84.22
-SAR DANet [18] CVPR’19 91.20 58.97 85.93 98.13 82.31 93.57 59.35 90.28 83.31
+SAR
+DANet [19] CVPR’19 91.20 58.97 85.93 98.13 82.31 93.57 59.35 90.28 83.31
+CGNet [21] TIP’20 81.16 19.58 69.45 90.27 60.78 84.63 25.42 74.96 64.25
 OCRNet [20] ECCV’20 86.38 38.70 78.36 96.18 71.34 89.52 41.07 83.54 74.19
-CGNet [20] TIP’20 81.16 19.58 69.45 90.27 60.78 84.63 25.42 74.96 64.25
+SETR [23] CVPR’21 90.97 59.57 84.71 98.16 80.51 93.27 59.39 89.99 82.79
 SegFormer [22] NeurIPS’21 88.81 46.92 81.99 97.57 76.75 91.51 49.55 86.73 78.41
-SETR [22] CVPR’21 90.97 59.57 84.71 98.16 80.51 93.27 59.39 89.99 82.79
 Mask2Former [24] CVPR’22 91.34 58.52 86.40 93.29 80.08 93.20 62.69 89.49 81.92
@@
 TransUNet [26] MIA’24 86.56 38.45 77.56 96.53 71.30 89.49 42.21 83.44 74.08
-SAR & RGB MCANet [28] JAG’21 85.12 47.88 78.69 96.96 73.18 89.59 46.13 85.55 76.37
-CEN [33] TPAMI’22 91.78 58.73 87.70 98.33 85.73 94.26 60.19 90.93 84.46
-TokenFusion [52] CVPR’22 93.90 70.67 90.62 98.94 89.21 95.79 69.64 93.70 88.67
-DDHRNet [29] JAG’22 79.78 13.85 69.85 93.73 62.17 84.73 24.59 73.75 63.88
-CDDFuse [12] CVPR’23 96.59 83.02 94.59 99.11 93.30 97.61 80.65 96.46 93.32
-CMX [31] TITS’23 96.43 82.59 94.37 99.38 93.72 97.58 81.08 96.44 93.30
-MRFS [30] CVPR’24 96.48 83.68 94.48 99.39 93.84 97.65 82.24 96.60 93.57
-FTransUNet [6] TGRS’24 84.69 35.21 76.51 96.40 71.11 88.56 36.55 82.35 72.78
-ASANet [32] ISPRS’24 95.04 77.09 92.64 99.24 91.82 96.73 75.77 95.21 91.17
+SAR & RGB
+MCANet [29] JAG’21 83.90 38.06 75.40 96.04 69.99 88.07 41.04 82.54 72.68
+DDHRNet [30] JAG’22 82.76 29.45 73.78 95.20 67.40 87.21 35.24 79.81 69.72
+CEN [34] TPAMI’22 86.61 40.08 79.34 97.57 74.26 90.20 46.91 84.50 75.57
+TokenFusion [55] CVPR’22 92.28 65.60 88.05 98.60 85.74 94.63 64.71 92.10 86.05
+CMX [32] TITS’23 93.50 72.03 90.02 99.00 88.03 95.59 70.48 93.65 88.52
+CDDFuse [13] CVPR’23 82.19 35.87 69.95 93.94 54.50 85.06 66.26 78.55 67.29
+FTransUNet [6] TGRS’24 86.78 40.51 78.70 97.24 73.13 89.91 46.02 84.35 75.27
+ASANet [33] ISPRS’24 86.67 43.18 79.02 97.28 73.99 90.17 46.40 85.03 76.03
+MRFS [31] CVPR’24 94.63 76.20 91.30 99.11 90.22 96.32 74.23 94.72 90.29
+ASMFNet [56] JSTARS’24 86.24 37.52 79.08 97.45 73.38 89.92 45.24 83.77 74.74
+MIEFNet [57] GRSL’24 90.56 63.90 85.20 98.39 82.19 93.46 61.20 90.89 84.05
+MFNet [58] TGRS’25 89.79 56.65 83.61 98.10 80.32 92.67 57.54 89.23 81.70
 PAD (Ours) 97.09 86.12 95.38 99.47 94.89 98.04 84.38 97.16 94.59
-IV-C2 Evaluations on DDHR-SK
+IV-B2 Evaluations on DDHR-SK
 As shown in Table III, PAD exhibits superior performance under adverse (cloudy) conditions on DDHR-SK.
-It achieves 94.59% mIoU and 97.65% OA, outperforming leading unimodal methods (e.g., Mask2former-RGB: 90.25% mIoU, 94.76% mF1; DANet-SAR: 83.31% mIoU, 90.28% mF1) and the multi-modal baseline MRFS (93.57% mIoU, 82.24% mKappa).
-PAD also attains 82.24% mKappa and 96.60% mF1, with boundary-sensitive categories achieving 99.47% IoU on water and 86.12% IoU on roads, surpassing MRFS by 2.44% and 0.08%, respectively.
-Qualitative results in Fig.
-7 further demonstrate PAD’s ability to delineate water infiltration in irrigated fields and maintain continuous road networks beneath fog.
+It achieves 94.59% mIoU and 98.04% OA, outperforming leading unimodal methods (e.g., Mask2Former-RGB: 90.25% mIoU, 94.76% mF1; DANet-SAR: 83.31% mIoU, 90.28% mF1) and the multi-modal baseline MRFS (90.29% mIoU, 74.23% mKappa).
+PAD also attains 84.38% mKappa and 97.16% mF1.
+Boundary-sensitive categories achieve 86.12% IoU on roads, surpassing MRFS by 9.92%.
 Figure 7: Visualization Results on DDHR-SK.
-Four boundary-focused scenarios: (a) coastal demarcation, (b) water-cropland interface, (c) vegetation-building complex, and (d) road network topology.
-For each scenario, panels show the RGB input, SAR coherence map, pseudo-color ground truth, and prediction maps from six methods.
-We compare PAD against unimodal baselines (Mask2former-RGB, DANet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS).
-Critical boundary regions are highlighted with orange boxes; see Section IV-C for details.
-The visual results further demonstrate PAD’s capability for cross-modal boundary refinement.
+Four boundary-focused scenarios: (a) coastal demarcation; (b) water-cropland interface; (c) vegetation-building complex; and (d) road-network topology.
+For each scenario, panels (left to right) show the RGB input, SAR coherence map, pseudo-color ground truth, and predictions from six methods.
+We compare PAD against unimodal baselines (Mask2Former-RGB and DANet-SAR) and multi-modal approaches (CDDFuse, CMX, and MRFS).
+Critical boundary regions are highlighted with orange boxes; see §IV-B for details.
+Visual results demonstrate PAD’s ability for cross-modal boundary refinement.
 In Fig.
@@
 7(b), PAD leverages SAR features to delineate water-infiltration boundaries in irrigated fields, overcoming the over-segmentation common in optical sensors.
-Moreover, in Fig.
-7(c) and Fig.
-7(d), PAD achieves continuous road networks under vegetation occlusion and precise material boundaries in vegetation-building complexes, respectively.
-TABLE IV: Computational Efficiency Comparison.
-Model size (#Params), inference time per image, and throughput (FPS) for competitive methods.
-Lower #Params and inference time, and higher FPS, indicate better efficiency.
-Method #Params (M) Inference Time (ms) FPS (frames/s)
-CMX [31] 15.00 206.70 1,957.8
-MRFS [30] 10.30 695.85 1,342.2
-ASANet [32] 7.82 72.91 2,245.2
-PAD (Ours) 3.88 5.16 3,228.7
-IV-C3 Computational Efficiency Analysis
-We evaluate the computational efficiency of PAD in terms of model size (Fusion Modules Only), inference time per image, and frames per second (FPS), as summarized in Table IV.
-We compare against four high-mIoU methods, CMX [31], MRFS [30], ASANet [32], and CDDFuse [12] (excluded from detailed analysis due to its two-stage design and lower efficiency).
-All metrics are aggregated across fusion stages.
-As shown in Table IV, PAD uses only 3.882 M parameters, achieves the fastest inference time (5.16 ms per image), and delivers the highest FPS (3,228.7).
-These results demonstrate that PAD not only matches or exceeds accuracy of existing methods but also offers real-time efficiency suitable for operational deployment.
-IV-C4 Ablation Study
-We performed comprehensive ablation studies on WHU-OPT-SAR to validate the contributions of our core components: the Amplitude Spectrum Fusion (ASF) module, the Phase Spectrum Correction (PSC) module, and Multi-Stage Fusion (MSF).
-[…]
-ID Components IoU per Category OA mKappa mF1 mIoU
-MSF ASF PSC Farmland City Village Water Forest Road Others
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+In Fig.
+7(c)-(d), PAD produces continuous road networks under vegetation occlusion and precise material boundaries in vegetation-building complexes.
+IV-C Ablation and Efficiency Analysis
+IV-C1 Ablation Study
+We conduct comprehensive ablations on WHU-OPT-SAR to validate the contributions of our core components: the Amplitude Spectrum Fusion (ASF) module, and the Phase Spectrum Correction (PSC) module.
+TABLE IV: Ablation Results (%) on Main Modules for WHU-OPT-SAR.
+and
+indicate the inclusion or exclusion of ASF and PSC, respectively.
+Components Per-class IoU OA mKappa mF1 mIoU
+ASF PSC Farmland City Village Water Forest Road Others
+69.46 57.16 49.01 65.15 83.87 41.57 21.39 84.32 51.96 69.23 55.37
+69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
+69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
+69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
 Main Modules
-Eight variants (see Table V) were evaluated.
-The baseline (I), which replaces MSF with early fusion and substitutes PAD (ASF+PSC) with pixel-wise addition, achieves the lowest mIoU (54.98%).
-Enabling MSF alone (II) yields a +0.39% gain, ASF alone (III) reaches 55.86% mIoU, and PSC alone (IV) provides a smaller improvement.
-Combining MSF with ASF (V) or PSC (VI) yields similar gains (56.11-56.12% mIoU).
-Integrating all three modules achieves the best performance (56.26% mIoU), demonstrating their effective synergy.
-TABLE VI: Ablation Study on Fusion Modules (%) for WHU-OPT-SAR. We compare alternative amplitude-only, phase-only, and full fusion designs by integrating each into our framework.
-Fusion Module IoU per Category OA mKappa mF1 mIoU
+Four variants (see Table IV) are evaluated.
+The baseline, which replaces substitutes PAD (ASF+PSC) with pixel-wise addition, achieves the lowest mIoU (55.37%).
+Enabling ASF alone reaches 56.11% mIoU; PSC alone provides 56.12% mIoU.
+Integrating two modules achieves the best performance (56.26% mIoU), demonstrating effective synergy.
+TABLE V: Ablation on Fusion Modules (%) for WHU-OPT-SAR. Amplitude-only, phase-only, and combined (amplitude+phase) fusion designs are compared, each integrated into our framework.
+Fusion module Per-class IoU OA mKappa mF1 mIoU
 Farmland City Village Water Forest Road Others
-Amplitude FelNFN [16] 66.93 55.80 45.15 61.27 82.77 35.54 19.17 82.95 47.31 66.46 52.38
+Amplitude
+FelNFN [17] 66.93 55.80 45.15 61.27 82.77 35.54 19.17 82.95 47.31 66.46 52.38
 FCENet [15] 68.27 55.18 46.79 61.17 83.43 37.09 16.71 83.43 49.82 66.52 52.66
 FGSwin [59] 69.87 57.43 49.03 65.04 84.12 43.16 22.26 84.51 53.53 69.70 55.85
-Ours 69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
-Phase FelNFN [16] 60.66 53.25 31.99 54.14 80.64 19.84 15.02 79.42 33.32 58.89 45.08
+PAD (Ours) 69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
+Phase
+FelNFN [17] 60.66 53.25 31.99 54.14 80.64 19.84 15.02 79.42 33.32 58.89 45.08
 Channel attention 69.58 58.66 49.78 64.28 83.96 43.43 22.65 84.39 53.54 69.93 56.05
-Ours 69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
-Whole FelNFN [16] 66.43 51.97 44.36 58.48 82.80 32.33 12.64 82.33 47.35 63.63 49.86
-FCENet++ [14] 66.87 53.01 46.41 60.30 82.64 30.21 14.51 82.78 43.83 64.33 50.56
+PAD (Ours) 69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
+Whole
+FelNFN [17] 66.43 51.97 44.36 58.48 82.80 32.33 12.64 82.33 47.35 63.63 49.86
+FCENet [15] 66.87 53.01 46.41 60.30 82.64 30.21 14.51 82.78 43.83 64.33 50.56
 FGSwin [59] 69.88 57.86 49.48 65.26 84.15 42.49 22.36 84.56 53.41 69.76 55.92
@@
 SFINet++ [50] 68.23 55.65 47.31 62.35 83.10 38.48 19.04 83.50 49.56 67.43 53.45
-Ours 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
+PAD (Ours) 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
 Fusion Modules
-To further validate ASF and PSC, we replaced each with five existing frequency-domain designs, FelNFN [16], FCENet [14], FGSwin [53], Frefusion [15], and SFINet++ [46], while keeping the rest of our framework unchanged (see Table VI).
-These methods represent common strategies such as fixed-kernel spectral convolution (FelNFN), frequency-spatial interaction via Hadamard products/matrix multiplication (FCENet), learned spectral filtering (FGSwin), attention-based cross-domain fusion (Frefusion), and amplitude-phase parallel processing (SFINet++).
-Our ASF outperforms other amplitude-specific methods (56.11% vs.
+To further validate ASF and PSC, we replace each with five existing frequency-domain designs, FeINFN [17], FCENet [15], FGSwin [59], Frefusion [16], and SFINet++ [50], while keeping the rest of our framework unchanged (see Table V).
+These methods represent strategies such as fixed-kernel spectral convolution (FeINFN), frequency-spatial interaction via Hadamard products/matrix multiplication (FCENet), learned spectral filtering (FGSwin), attention-based cross-domain fusion (Frefusion), and amplitude-phase parallel processing (SFINet++).
+Our ASF outperforms other amplitude-focused methods (56.11% vs.
 55.85% mIoU for FGSwin), and PSC consistently exceeds phase baselines (56.12% vs.
-56.05% for channel attention and 45.08% for FelNFN).
-At the full-fusion level, our integrated design still leads (56.26% mIoU), indicating both modules are effective in isolation and synergistic when combined.
-These results confirm that adaptive frequency modulation (ASF) and spatially guided phase correction (PSC) offer structurally unique and more effective alternatives to existing paradigms.
-TABLE VII: Ablation Study on Loss Function Weights (%) for WHU-OPT-SAR. We vary $\lambda_{1}$ with fixed $\lambda_{2}=0.1$ and $\lambda_{2}$ with fixed $\lambda_{1}=0.4$.
-Configuration IoU per Category OA mKappa mF1 mIoU
+56.05% for channel attention and 45.08% for FeINFN).
+At the full-fusion level, our integrated design still leads (56.26% mIoU), indicating that both modules are effective in isolation and synergistic when combined.
+These results confirm that adaptive frequency modulation (ASF) and spatially guided phase correction (PSC) offer structurally distinct and more effective alternatives to existing paradigms.
+TABLE VI: Ablation on Loss-Function Weights (%) for WHU-OPT-SAR. We vary $\lambda_{1}$ (with $\lambda_{2}$ fixed at 0.1) and $\lambda_{2}$ (with $\lambda_{1}$ fixed at 0.4).
+Configuration Per-class IoU OA mKappa mF1 mIoU
 Farmland City Village Water Forest Road Others
@@
 Loss Function Weights
-We conducted weight ablations (see Table VII) to assess the trade-off between the amplitude consistency loss $\mathcal{L}_{\mathrm{amp}}$ and segmentation loss $\mathcal{L}_{\mathrm{seg}}$.
+We conduct weight ablations (see Table VI) to assess the trade-off between the amplitude consistency loss $\mathcal{L}_{\mathrm{amp}}$ and the segmentation loss $\mathcal{L}_{\mathrm{seg}}$.
 The best mIoU (56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$.
-Increasing $\lambda_{1}$ above 0.4 degrades performance, especially on under-represented classes, while too small $\lambda_{2}$ significantly reduces accuracy, confirming segmentation supervision must remain dominant.
-TABLE VIII: Ablation Study on Initial Frequency Radius (%) for WHU-OPT-SAR. We vary the trainable radius parameter in the ASF module.
-Radius IoU per Category OA mKappa mF1 mIoU
+Increasing $\lambda_{1}$ above 0.4 degrades performance, especially on under-represented classes, whereas too small $\lambda_{2}$ significantly reduces accuracy, confirming that segmentation supervision must remain dominant.
+TABLE VII: Ablation on Initial Frequency Radius (%) for WHU-OPT-SAR. The trainable radius parameter in the ASF module is varied.
+Radius Per-class IoU OA mKappa mF1 mIoU
 Farmland City Village Water Forest Road Others
-$0.01$ 70.08 57.44 49.31 66.29 84.18 43.08 22.25 84.66 53.92 69.88 56.09
-$0.1$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
-$1$ 69.79 58.19 49.71 64.81 84.15 43.44 23.03 84.52 53.88 70.04 56.16
-$10$ 69.71 57.71 49.69 65.63 84.18 43.21 23.56 84.57 53.93 70.12 56.24
-$100$ 69.90 57.23 49.57 65.23 84.13 43.34 22.94 84.53 53.88 69.93 56.05
+0.01 70.08 57.44 49.31 66.29 84.18 43.08 22.25 84.66 53.92 69.88 56.09
+0.1 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
+1 69.79 58.19 49.71 64.81 84.15 43.44 23.03 84.52 53.88 70.04 56.16
+10 69.71 57.71 49.69 65.63 84.18 43.21 23.56 84.57 53.93 70.12 56.24
+100 69.90 57.23 49.57 65.23 84.13 43.34 22.94 84.53 53.88 69.93 56.05
 Initial Frequency Radius
-We varied the initial frequency radius $r$, which controls the preserved low-frequency proportion in ASF (see Table VIII).
+We vary the initial frequency radius $r$, which controls the preserved low-frequency proportion in ASF (see Table VII).
 The optimal balance is at $r=0.1$ (56.26% mIoU).
 Values that overly suppress or over-smooth low-frequency content degrade performance, although the method remains robust for $r\in[0.1,10]$.
-IV-D Visualization at Distribution Level
-To further analyze feature distributions, we extracted multi-scale features from the RGB and SAR encoders and from the fusion modules.
-After upsampling to the original resolution, we concatenated features channel-wise.
-To balance computational efficiency with discriminative power, we applied a two-stage dimensionality reduction: a) random spatial sampling, and b) principal component analysis (PCA), retaining 50 components (over 89% variance).
-The final visualization was generated using t-SNE [54] (perplexity = 30, 1,000 iterations).
+TABLE VIII: Computational Efficiency Comparison.
+Model size (#Params), per-image inference time, and throughput (FPS) for competing methods.
+Fewer parameters and shorter inference times, along with higher FPS, indicate better efficiency.
+Method #Params (M) Inference time (ms) Throughput (FPS)
+CMX [32] 15.00 206.70 1,957.8
+MRFS [31] 10.30 695.85 1,342.2
+ASANet [33] 7.82 72.91 2,245.2
+PAD (Ours) 3.88 5.16 3,228.7
+IV-C2 Computational Efficiency
+We evaluate computational efficiency in terms of model size (fusion modules only), inference time per image, and frames per second (FPS), summarized in Table VIII.
+We compare against three high-mIoU methods, CMX [32], MRFS [31], and ASANet [33].
+All metrics are aggregated across fusion stages with the same baseline model.
+As shown in Table VIII, PAD uses only 3.882 M parameters, achieves the fastest inference time (5.16 ms per image), and delivers the highest FPS (3,228.7).
+These results indicate that PAD not only matches or exceeds the accuracy of existing methods but also offers real-time efficiency suitable for operational deployment.
+IV-D Further Analysis and Visualization
+IV-D1 Visualization at the Distribution Level
+To further analyze feature distributions, we extract multi-scale features from the RGB and SAR encoders and from the fusion modules.
+After upsampling to the original resolution, we concatenate features channel-wise.
+To balance computational efficiency with discriminative power, we apply a two-stage dimensionality reduction: a) random spatial sampling and b) principal component analysis (PCA), retaining 50 components (over 89% variance).
+The final visualization is generated using t-SNE [60] (perplexity = 30, 1,000 iterations).
 Refer to caption Refer to caption Refer to caption
@@
 Figure 8: t-SNE Visualization of Cross-Modal Feature Distribution Evolution.
-(a) RGB encoder features show spectral continuity but blurred class boundaries.
-(b) SAR encoder features capture geometric details but are affected by speckle noise.
-(c) PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separation.
+(a) RGB-encoder features show spectral continuity but exhibit blurred class boundaries; (b) SAR-encoder features capture geometric details but are affected by speckle noise; (c) PAD-fused features demonstrate enhanced intra-class compactness and improved inter-class separability.
 As shown in Fig.
 8, PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability.
-Quantitatively, mean intra-class standard variance is reduced by 20.39% for roads, 13.16% for water, 6.46% for forest, and 4.43% for farmland, compared to the best unimodal baselines.
-Moreover, Euclidean distances between critical class pairs increase significantly: City/Village by 75.75% (66.56 $to$ 116.95) and Farmland/Forest by 13.29% (107.64 $to$ 121.89), indicating that PAD captures cross-modal discriminative patterns effectively.
-TABLE IX: Statistical Comparison Between Models w/ and w/o ASF.
+Quantitatively, mean intra-class standard deviation is reduced by 20.39% for roads, 13.16% for water, 6.46% for forest, and 4.43% for farmland compared with the best unimodal baselines.
+Moreover, Euclidean distances between critical class pairs increase significantly: City/Village by 75.75% ($66.56\rightarrow 116.95$) and Farmland/Forest by 13.29% ($107.64\rightarrow 121.89$), indicating that PAD captures cross-modal discriminative patterns effectively.
+TABLE IX: Performance Comparison (%) of Models w/ and w/o PSC under Missing RGB Regions and Cloud Obscuration on WHU-OPT-SAR.
+Scenario Method Per-class IoU OA Kappa mF1 mIoU
+Farmland City Village Water Forest Road Others
+Missing RGB w/o PSC 74.65 75.91 58.03 80.18 56.84 57.81 59.49 82.28 70.94 79.23 66.13
+w/ PSC 74.65 76.25 58.05 80.25 57.09 58.54 59.82 82.35 71.12 79.42 66.38
+Cloud Obscuration w/o PSC 75.33 75.47 57.42 76.31 84.30 53.40 44.94 86.90 63.06 79.23 66.74
+w/ PSC 75.37 75.73 57.56 76.47 84.32 53.68 45.30 86.95 63.52 79.37 66.92
+IV-D2 Robustness of PSC under Challenging Scenarios
+To further evaluate the robustness of PSC beyond normal distributions, we conducted targeted ablation experiments under extreme long-tail conditions, including missing RGB regions and cloud obscuration.
+Results are summarized in Table IX.
+Missing Regions
+When large portions of the RGB modality are absent, models without PSC accumulate residual phase errors, particularly in categories with fine spatial structures (e.g., forest, road).
+With PSC, alignment remains stable and yields consistent improvements across macro-level metrics (e.g., +0.19 mF1, +0.25 mIoU).
+Cloud Obscuration
+Under cloud occlusion, traditional co-registration often provides weak or misleading guidance.
+PSC mitigates this issue by preserving phase consistency, resulting in systematic improvements in category-wise IoU (e.g., city, village) as well as overall agreement scores.
+These results confirm that PSC not only aligns with physical and statistical expectations but also maintains robust performance under adverse long-tail scenarios, where purely physical models typically degrade.
+TABLE X: Statistical Comparison between Models w/ and w/o ASF.
 Metric w/ ASF w/o ASF
@@
 Mean Spearman correlation 0.0482 0.0488
-Paired t-test p-value 0.0016
+Paired $t$-test pp-value 0.0016
 Wilcoxon test $p$-value $6.4\times 10^{-7}$
 Cohen’s $d$ (effect size) -0.038
-IV-E Noise Sensitivity of ASF
-To assess whether ASF amplifies high-frequency speckle noise in SAR images, we conducted a correlation analysis using the Equivalent Number of Looks (ENL), a standard measure of speckle.
-For each image, we computed pixel-wise ENL maps and then measured the Pearson and Spearman correlations between ENL and classification accuracy.
-As shown in Table IX, the model with ASF exhibits nearly identical correlation coefficients to the baseline, differences in mean correlation are below 0.001, and Cohen’s $d=-0.038$ indicates a negligible effect size.
+IV-D3 Noise Sensitivity of ASF
+To assess whether ASF amplifies high-frequency speckle noise in SAR images, we perform a correlation analysis using the Equivalent Number of Looks (ENL), a standard measure of speckle.
+For each image, we compute pixel-wise ENL maps and then measure Pearson and Spearman correlations between ENL and classification accuracy.
+As shown in Table X, the model with ASF exhibits nearly identical correlation coefficients to the baseline; differences in mean correlation are below 0.001, and Cohen’s d=0.038d=-0.038 indicates a negligible effect size.
 These results confirm that ASF does not increase noise sensitivity and remains robust in low-SNR regions.
-Overall, these results confirm that PAD significantly improves classification accuracy and prediction consistency by integrating multi-modal information, suppressing noise, and preserving discriminative spectral and geometric characteristics inherent in the data.
+Overall, these findings confirm that PAD significantly improves classification accuracy and prediction consistency by integrating multi-modal information, suppressing noise, and preserving discriminative spectral and geometric characteristics inherent in the data.
 V Conclusion
-In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land cover classification.
-Our frequency-domain analysis quantitatively demonstrates that phase components remain consistent across modalities, while amplitude spectra capture modality-specific details, providing the first physical prior for multi-source remote sensing fusion.
-PAD disentangles and aligns phase features via the Phase Spectrum Correction (PSC) module, and adaptively fuses amplitude components using the Amplitude Spectrum Fusion (ASF) module, which integrates SAR’s high-frequency details with RGB’s low-frequency context.
+In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land-cover classification.
+Our frequency-domain analysis quantitatively demonstrates that phase components remain consistent across modalities, whereas amplitude spectra capture modality-complementary details, providing the first physically grounded prior for multi-source remote-sensing fusion.
+PAD decouples and aligns phase features via the Phase Spectrum Correction (PSC) module and adaptively fuses amplitude components using the Amplitude Spectrum Fusion (ASF) module, effectively leveraging SAR’s high-frequency details together with RGB’s low-frequency context.
 Extensive experiments on two challenging datasets show that PAD outperforms existing methods in both accuracy and robustness.